### v2

# Continuum-Field-Theoretical Construction of Conserved Currents, Non-Invertible Symmetries, and Mixed Anomalies in (3+1)D Non-Abelian Topological Order

Abstract
A modern perspective views topological order as a symmetry-breaking phase of generalized symmetries, making the identification of such symmetries, especially in non-Abelian topological orders, a fundamental task. In this work, we investigate generalized symmetries, with particular emphasis on non-invertible ones, in three-dimensional non-Abelian topological orders hosting both particle- and loop-like excitations. We adopt a continuum topological field theory description, focusing on twisted $BF$ theories with gauge group $G=\prod_{i}\mathbb{Z}_{N_{i}}$ and an $a\wedge a\wedge b$ twisted term. This field theory supports Borromean-Rings braiding and realizes non-Abelian topological order, which for $G=(\mathbb{Z}_{2})^{3}$ admits a microscopic realization via the $\mathbb{D}_{4}$ Kitaev quantum double lattice model. We systematically identify all generalized symmetry operators by extracting conserved currents from the equations of motion. Two distinct classes of currents emerge: type-I currents, which generate invertible higher-form symmetries, and type-II currents, which give rise to non-invertible higher-form symmetries. The non-invertibility originates from projectors accompanying the symmetry operators, which restrict admissible gauge-field configurations. We further analyze the fusion rules of these symmetries, showing that invertible symmetries admit inverses, while non-invertible symmetries fuse through multiple channels. Finally, we study mixed anomalies among these generalized symmetries by simultaneously coupling multiple currents to proper types of background gauge fields and examining their gaugeability. We identify two types of mixed anomalies: one cancellable by topological field theories in one higher dimension, and another representing an intrinsic gauging obstruction encoded in the (3+1)(3+1)D continuum topological field theory. Our results provide a concrete field-theoretical construction of conserved currents and a unified characterization of invertible and non-invertible higher-form symmetries in three-dimensional non-Abelian topological orders, offering a natural framework for studying (2+1)(2+1)D systems via symmetry topological field theory and topological holography.
I Introduction
Symmetry has long been a central organizing principle in physics. Broadly speaking, a symmetry of a physical system is a transformation that leaves the underlying theory invariant when applied to the system as a whole. A familiar example is provided by the one-dimensional [^1] In this paper, “(n+1)(n+1)D” refers to $(n+1)$-dimensional spacetime with $n$-dimensional real space. We avoid using “nnD” or “(n)(n)D” unless otherwise specified. When referring specifically to spatial dimensions or to the dimensions of geometric objects, we use “nnd”, e.g., 33d ground state, 33d topological order, 22d square lattice, and 11d submanifold. (11d) transverse-field Ising model, which is invariant under a simultaneous flip of all spins. This global spin-flip operation acts on the entire system and leaves the Hamiltonian unchanged, and is therefore identified as a symmetry. Local operators may transform nontrivially under this operation and are said to carry symmetry charge. Since performing the global spin flip twice returns the system to itself, this symmetry has a $\mathbb{Z}_{2}$ group structure.
In recent years, the notion of symmetry has been substantially generalized from multiple perspectives across condensed matter, high-energy, and mathematical physics [1, 2, 3, 4, 5, 6, 7, 8]. Two complementary aspects are particularly important. First, symmetry operations need not act on the entire system; instead, they may be supported on submanifolds [1]. As a result, charged operators become extended rather than point-like, leading to the concept of higher-form symmetry. Within this framework, the $\mathbb{Z}_{2}$ symmetry of the 11d transverse-field Ising model discussed above is an example of a 0-form symmetry.
A canonical illustration of higher-form symmetry is provided by the 22d toric code model [9] on a square lattice, where symmetry operators can be constructed on closed loops rather than on the full lattice. These loop operators commute with the Hamiltonian and generate symmetries associated with one-dimensional submanifolds embedded in the two-dimensional system. Operators supported on one-dimensional paths that intersect these loops transform nontrivially under the symmetry and are therefore charged. Such symmetries are referred to as $1$-form symmetries, reflecting the co-dimensionality of their support in 22d real space. At long wavelengths, the supports of higher-form symmetries are manifold-like and can be continuously deformed without affecting their action. There exists another important generalization, namely subsystem symmetries, whose supports are rigid rather than deformable and are typically fixed to specific geometric patterns such as lines, fractals, or more exotic structures [10, 11, 12, 13]. In this work, however, we focus exclusively on higher-form symmetries and leave subsystem symmetries for future investigation.
Second, the algebraic structure organizing symmetry operators need not be a group [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. More generally, it may be described using the language of category theory, giving rise to non-invertible or categorical symmetries. A representative example arises in the 11d transverse-field Ising model at criticality, which admits a symmetry operation implementing Kramers-Wannier duality. This operation exchanges spin and domain-wall variables and commutes with the Hamiltonian, yet it does not admit an inverse in the conventional sense: acting with it projects out a subset of states. Such an operation therefore realizes a non-invertible symmetry. For non-invertible symmetries, composing two symmetry operations may not yield a single operator but rather a linear combination of symmetry operators governed by fusion rules. By contrast, for invertible symmetries the fusion rules reduce to the familiar group composition law.
These modern generalizations of symmetry provide powerful tools for the study of quantum many-body systems, particularly topologically ordered phases. Such phases, traditionally viewed as lying beyond the Landau-Ginzburg paradigm, can be reinterpreted as phases that spontaneously break generalized symmetries. For example, the 22d toric code model on a torus exhibits a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 11-form symmetry generated by operators supported on the two noncontractible cycles of the torus. Since these operators do not commute, no ground state can remain invariant under both simultaneously, leading inevitably to topological ground-state degeneracy. This phenomenon can be understood as spontaneous breaking of a higher-form symmetry. In this way, generalized symmetries offer a unifying framework for understanding both the structure and robustness of topologically ordered phases.
Building on the concept of generalized symmetry, a framework known as symmetry topological field theory (SymTFT), also referred to as the symmetry-topological-order correspondence, has been rapidly developed. The central idea is that a $(d+1)$-dimensional topological quantum field theory (TQFT), or equivalently a (d+1)(d+1)D topological order, encodes the symmetries of a (d1+1)(d-1+1)-dimensional quantum field theory. In this correspondence, topological operators in the bulk encode generalized symmetries of the TQFT and can terminate on the boundary in different ways, thereby characterizing distinct lower-dimensional systems. This framework has found broad applications, including the classification of gapped and gapless phases, the diagnosis of generalized symmetry anomalies, the analysis of renormalization-group flows, the study of mixed-state phases and phase transitions, and so on [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. To date, SymTFT has been particularly successful for (2+1)(2+1)D bulks and (1+1)(1+1)D boundaries. Extending this framework to (2+1)(2+1)D systems therefore naturally motivates a detailed understanding of generalized symmetries in (3+1)(3+1)D topologically ordered phases.
As effective field theories of topological order, TQFTs are expected to encode generalized symmetries in a natural and systematic manner. For instance, Wilson operators in Chern-Simons theory [36, 37, 38] act as generators of $1$-form symmetries. Analogous to the role of Chern-Simons theory in describing 22d topological orders, $BF$ theory [39, 40], together with its twisted variants [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], provides a unified field-theoretical framework for characterizing 33d topological orders and gauged symmetry-protected topological phases with particle and loop excitations. Geometrically, the canonical $BF$ term bdab\,da (with $b$ a $2$-form field and $a$ a $1$-form field) captures the characteristic particle-loop braiding in 33d topological order. [^2] Throughout this paper, we adopt a condensed notation for differential forms and omit the wedge symbol $\wedge$ when no confusion arises. For example, bdab\,da stands for $b\wedge da$, and similarly aadaaada denotes topological terms of $a\wedge a\wedge da$ type.
More intricate braiding processes, including multi-loop and Borromean rings braiding, are encoded by twisted terms such as aadaaada, aaaaaaaa, and aabaab [40, 44, 48, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. These braiding processes are subject to compatibility conditions imposed by gauge invariance [49]. Additional terms, such as the twisted $bb$ term and the $\theta$ term dadada\,da, encode emergent fermionic statistics [45, 50, 66, 67] and topological responses [41, 43, 51, 68, 69, 70, 71, 72], respectively. Further developments revealed symmetry fractionalization on loop excitations, leading to mixed three-loop braiding and the classification of symmetry-enriched topological phases in higher dimensions [73, 74, 75]. The fusion and shrinking rules of topological excitations in 33d topological orders have also been investigated within the $BF$-theory framework [76]. Motivated by this field-theoretical foundation, a diagrammatic representation of higher-dimensional topological orders, including 44d cases [77], was recently developed in Ref. [78]. Given this progress, a natural question arises: how can generalized symmetries of 33d topological orders be systematically identified directly from effective field theory?
In this work, we address this question by studying generalized symmetries in a class of (3+1)(3+1)D topological quantum field theories describing three-dimensional non-Abelian topological orders with nontrivial Borromean-Rings braiding. We focus on twisted $BF$ theories with gauge group $G=\prod_{i}\mathbb{Z}_{N_{i}}$ and an $a\wedge a\wedge b$ twisted term [48, 49, 76]. For the special case $G=(\mathbb{Z}_{2})^{3}$, the resulting non-Abelian topological order admits a microscopic realization in terms of the $\mathbb{D}_{4}$ Kitaev quantum double lattice model [79]. These field theories therefore provide a concrete continuum description in which higher-form generalized symmetries, including non-invertible ones, can be systematically analyzed. In quantum mechanics, an operator commuting with the Hamiltonian signals a symmetry, while in field theory the corresponding notion is a conserved current. Guided by this principle, we identify generalized symmetries directly from the effective field theory by extracting conserved currents from the equations of motion of the gauge fields.
We find that the equations of motion naturally separate into two distinct classes, leading to two types of conserved currents. Type-I currents are conserved identically and generate invertible higher-form symmetries. By contrast, type-II currents are conserved only upon imposing additional constraints on the gauge-field configurations. As a result, the associated symmetry operators necessarily involve projector structures that enforce these constraints. Since such projectors do not admit inverses, the corresponding higher-form symmetries are intrinsically non-invertible.
With the explicit symmetry operators constructed from these conserved currents, we analyze their fusion rules by examining the composition of symmetry operations. We show that invertible higher-form symmetries possess well-defined inverses and obey group-like fusion rules, whereas non-invertible symmetries fuse through multiple channels, reflecting the presence of projector-induced constraints. Finally, we investigate mixed anomalies among these generalized symmetries by simultaneously coupling multiple conserved currents to appropriate background gauge fields and attempting to gauge the corresponding symmetries. We identify two distinct types of mixed anomalies: one that can be canceled by embedding the theory into a topological field theory in one higher dimension, and another that represents an intrinsic gauging obstruction already encoded in the (3+1)(3+1)D continuum topological field theory.
This paper is organized as follows. In Sec. II, we review generalized symmetries, emphasizing the roles of conserved currents and topological operators. In Sec. III, we study generalized symmetries in the effective field theory of a class of 33d non-Abelian topological orders and construct the corresponding symmetry operators from conserved currents. Sec.p IV is devoted to the analysis of fusion rules. In Sec. V, we discuss anomalies of generalized symmetries. We conclude in Sec. VI with a summary and an outlook.
II Preliminaries of conserved currents and symmetry operators
In field theory, a symmetry is indicated by a conserved current, which in turn gives rise to a topological operator, namely the symmetry operator associated with the symmetry. From this perspective, a symmetry refers to the mathematical structure formed by these topological operators. In our discussion of generalized symmetries, we adopt the viewpoint that a higher-form symmetry is generated by topological operators originating from higher-form conserved currents. Moreover, this principle naturally extends to the construction of non-invertible symmetries, whose symmetry operators can be traced back to a special class of conserved currents, as will be discussed in Sec. III. Two representative examples of generalized symmetries in lattice systems are collected in Appendix A.
In the language of generalized symmetry, an ordinary global symmetry is referred to as a 0-form symmetry, and it is associated with a conserved vector current. As a concrete example, let us consider the global $U\left(1\right)$ symmetry corresponding to particle-number conservation in (3+1)(3+1)D. The spacetime current of particles satisfies the conservation law $\partial_{\mu}j_{\mu}=0$, where $j_{0}$ denotes the particle-number density and $j_{i}$ denotes the particle-current density. In the language of differential geometry, $j$ is a $1$-form, and the previous continuity equation can be written as $d\left(*j\right)=0$, where * denotes the Hodge dual. The total particle number of the 33d system is given by $Q=\int_{\Sigma_{3}}\rho\left(\vec{x},t\right)dx^{1}dx^{2}dx^{3}=\int_{\Sigma_{3}}*j$, where $\Sigma_{3}$ is the 33d space at a time slice. Let us consider another time slice $\Sigma_{3}^{\prime}$ such that $\Sigma_{3}^{\prime}-\Sigma_{3}=\partial M_{4}$ with $M_{4}$ a four-dimensional manifold. The increase of total particle number is given by $\int_{\Sigma_{3}^{\prime}}*j-\int_{\Sigma_{3}}*j=\int_{\partial M_{4}}*j$ which can be further written as $\int_{M_{4}}d\left(*j\right)$ by Stokes’ theorem and vanishes due to $d\left(*j\right)=0$. Therefore, the particle number $Q$ is independent of the choice of time slice and is a conserved quantity, implying the existence of a symmetry. This conserved particle number generates the global $U\left(1\right)$ symmetry, whose symmetry operator is
$U_{\alpha}\left(\Sigma_{3}\right)=\exp\left({\rm i}\alpha Q\right)=\exp\left({\rm i}\alpha\int_{\Sigma_{3}}*j\right)\,$, (1)
where the real parameter $\alpha$ takes value in $0,2\pi)$ as a compact phase angle. The compactness of $\alpha$ arises from the quantization of $Q$. The existence of this global symmetry can thus be traced back to the presence of a conserved current. The conservation law ensures that $U_{\alpha}$ is invariant under smooth deformations of $\Sigma_{3}$, hence the terminology topological operator of the 0-form symmetry (i.e., global symmetry) that act on the entire quantum system.
The concept of higher-form symmetry, e.g., $p$-form symmetry, generalizes the above structure by allowing the conserved current to be a $(p+1)$-form. Correspondingly, the associated symmetry operator is supported on a $\left(D-1-p\right)$-dimensional submanifold in (D1+1)(D-1+1)D spacetime. As an illustrative example of higher-form symmetries, we consider the (3+1)(3+1)D Maxwell’s electromagnetic theory with action $S=\frac{1}{4}\int d^{4}xF_{\mu\nu}F^{\mu\nu}$, where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ and $e$ is the gauge coupling constant; also see the similar discussion in, e.g., Refs [8, 1, 3, 6]. First, there exists a conserved $2$-form current satisfying $d\left(*J_{m}\right)=0$, with $J_{m}^{\mu\nu}=\frac{1}{4\pi}\epsilon^{\mu\nu\lambda\sigma}F_{\lambda\sigma}=\frac{1}{2\pi}\left(d\widetilde{A}\right)^{\mu\nu}$. Here $\widetilde{A}$ is defined by $d\widetilde{A}=*dA$. This continuity equation is equivalent to $dF=0$, namely the Bianchi identity. The conserved current $J_{m}$ gives rise to a “magnetic” $1$-form symmetry with symmetry operator
$U_{\alpha}^{\left(m\right)}\left(\Sigma_{2}\right)=\exp\left({\rm i}\alpha\int_{\Sigma_{2}}*J_{m}\right)=\exp\left({\rm i}\frac{\alpha}{2\pi}\int_{\Sigma_{2}}F\right).$ (2)
The integral $\frac{1}{2\pi}\int_{\Sigma_{2}}F$ counts the number of magnetic monopoles enclosed by the 22d submanifold $\Sigma_{2}$. Meanwhile, there exists another conserved $1$-form current, $J_{e}=F$, satisfying the continuity equation $d\left(*J_{e}\right)=d\left(*F\right)=0$ that holds wherever the electric charges are absent. The conserved current $J_{e}$ generates an “electric” $1$-form symmetry with symmetry operator
$U_{\alpha}^{\left(e\right)}\left(\Sigma_{2}\right)=\exp\left({\rm i}\alpha\int_{\Sigma_{2}}*J_{e}\right)=\exp\left({\rm i}\alpha\int_{\Sigma_{2}}*F\right).$ (3)
Another illustrative example is provided by the mutual Chern–Simons theory, which serves as the effective field theory of the 22d topological order,
$S=\int_{M_{3}}\frac{2}{2\pi}a^{1}da^{2}.$ (4)
The lattice realization of this field theory is the celebrated toric code model in 2d square lattice. Here $a^{1}$ and $a^{2}$ are $1$-form gauge fields. The equations of motion are $\frac{1}{\pi}da^{1}=0$ and $\frac{1}{\pi}da^{2}=0$. In analogy with the previous examples, we interpret these equations as continuity equations and introduce the conserved quantities $Q_{1}=\int_{\gamma}a^{1}$ and $Q_{2}=\int_{\gamma}a^{2}$. These quantities measure the number of topological excitations enclosed by the area bounded by a loop $\gamma$. They generate two minimal symmetry operators,
$U_{e}\left(\gamma\right)=\exp\left({\rm i}\int_{\gamma}a^{2}\right)\,,\,\,\,U_{m}\left(\gamma\right)=\exp\left({\rm i}\int_{\gamma}a^{1}\right)$ (5)
which satisfy the algebra $U_{e}\left(\gamma_{1}\right)U_{m}\left(\gamma_{2}\right)=\left(-1\right)^{{\rm Lk}\left(\gamma_{1},\gamma_{2}\right)}U_{m}\left(\gamma_{2}\right)U_{e}\left(\gamma_{1}\right)$. ${\rm Lk}\left(\gamma_{1},\gamma_{2}\right)$ denotes the linking number of the two loops. The equations of motion further imply that $\int_{\gamma}2a^{1}=0\mod 2\pi$ and $\int_{\gamma}2a^{2}=0\mod 2\pi$. Consequently, applying $U_{e}\left(\gamma\right)$ or $U_{m}\left(\gamma\right)$ twice yields the identity operation, and both symmetry operators have a $\mathbb{Z}_{2}$ group structure. In this sense, the field theory (4) exhibits a $\mathbb{Z}_{2,e}^{\left(1\right)}\times\mathbb{Z}_{2,m}^{\left(1\right)}$ 11-form symmetry implemented by $U_{e}\left(\gamma\right)$ and $U_{m}\left(\gamma\right)$, where the superscript $\left(1\right)$” indicates the $1$-form nature of the symmetry and the subscripts label the two symmetry generators.
III Field-theoretical construction of symmetry operators
In this section, we construct symmetry operators in effective field theories of 33d topological orders. We begin with the $\left(3+1\right)$D $\mathbb{Z}_{N}$ BFBF theory as a warm-up, which describes $\mathbb{Z}_{N}$ topological orders. We then turn to the main goal of this paper: identifying symmetry operators in the effective field theory of a class of 33d non-Abelian topological orders, namely Borromean-Rings (BR) topological orders. In particular, we show how conserved currents can be exploited to construct non-invertible symmetry operators.
A system with BR topological order [48, 49, 76, 79] exhibits nontrivial BR braiding statistics. A BR braiding process involves one particle and two loop excitations: the particle is moved around the two loops such that its spatial trajectory, together with the two loops, forms Borromean rings. Borromean rings are three closed curves in 33d space that are not pairwise linked, yet cannot be separated as a whole. Consequently, there is no net particle–loop braiding phase, since the particle trajectory is not linked with either loop individually. BR braiding is also distinct from multi-loop braiding because it necessarily involves a particle excitation. The corresponding TQFT is a $BF$ theory with an aabaab-type twisted term; see Eq. (14) below. Remarkably, BR topological order supports non-Abelian particle and loop excitations even when all excitations carry gauge charges and fluxes of an Abelian gauge group, making it an ideal playground for exploring generalized symmetries in $\left(3+1\right)$D systems.
In the following, we first derive the equations of motion (EoM) of gauge fields, from which we identify conserved currents. These conserved currents imply symmetries and allow us to construct the corresponding symmetry operators. We find two classes of conserved currents denoted as type-I and type-II respectively, as shown in Sec. III.3 and Sec. III.4. Type-I currents lead to group-like (i.e., invertible) symmetries. Type-II currents are conserved only under certain constraints such that corresponding conserved quantities are legitimate. Accordingly, the symmetry operators generated by these conserved quantities must be accompanied by projectors that enforce the constraints. Since projectors have no inverses, these symmetry operators are naturally non-invertible.
III.1 Higher-form symmetries in $\left(3+1\right)$D $\mathbb{Z}_{N}$ BFBF theory
The topological field theories describing the 22d and 33d toric code models discussed above are $\mathbb{Z}_{2}$ BFBF theories. It is straightforward to generalize the gauge group from $\mathbb{Z}_{2}$ to a general $\mathbb{Z}_{N}$. The $\left(3+1\right)$D $\mathbb{Z}_{N}$ BFBF theory is described by
$S=\int_{M_{4}}\frac{N}{2\pi}bda.$ (6)
Here $b$ is a $2$-form gauge field.
The equations of motion (EoMs) are
$\frac{N}{2\pi}da=0\quad\text{and}\quad\frac{N}{2\pi}db=0.$ (7)
We interpret them as continuity equations,
$d\left(*J_{e}\right)=0\quad\text{with}\quad J_{e}=-*a,$ (8)
$d\left(*J_{m}\right)=0\quad\text{with}\quad J_{m}=*b,$ (9)
where $J_{e}$ and $J_{m}$ are $3$-form and $2$-form conserved currents, respectively. Analogous to ordinary 0-form symmetries, these conserved currents give rise to conserved quantities
$Q_{e}\left(\gamma\right)=\int_{\gamma}*J_{e}=\int_{\gamma}a\,,\,\,\,Q_{m}\left(\sigma\right)=\int_{\sigma}*J_{m}=\int_{\sigma}b,$ (10)
where $\gamma$ and $\sigma$ are closed 11d and 22d submanifolds, respectively.
We now explain the physical meaning of these conserved quantities. Consider $Q_{e}\left(\gamma\right)=\int_{\gamma}a$ as an example. From the correspondence between gauge theory and topological order, the operator $e^{{\rm i}\int_{\gamma}a}$ is the Wilson operator associated with a particle excitation carrying one unit of gauge charge, and $\gamma$ can be interpreted as the worldline of this particle. Acting with $e^{{\rm i}\int_{\gamma}a}$ on a state amounts to transporting the particle along $\gamma$ in spacetime. This process probes the gauge flux carried by loop excitations piercing the surface bounded by $\gamma$.
From the perspective of conserved currents, since $a=a_{\mu}dx^{\mu}$, the current takes the form $J_{e}=-\frac{1}{3!}\epsilon^{\mu\nu\lambda\rho}a_{\mu}dx^{\nu}\wedge dx^{\lambda}\wedge dx^{\rho}$. For comparison, in a $\left(3+1\right)$D system with a 0-form $U\left(1\right)$ symmetry, the particle current $j=j_{\mu}dx^{\mu}$ is a $1$-form current, whose Hodge dual is $j=j_{0}dx^{1}\wedge dx^{2}\wedge dx^{3}-j_{1}dx^{0}\wedge dx^{2}\wedge dx^{3}+j_{2}dx^{0}\wedge dx^{1}\wedge dx^{3}-j_{3}dx^{0}\wedge dx^{1}\wedge dx^{2}$. Taking $M_{3}$ to be the entire 33d spatial manifold, $\int_{M_{3}}*j$ integrates the particle-number density $j_{0}$ over $M_{3}$, yielding the total particle number. By direct analogy, $\int_{\gamma}*J_{e}$ can be interpreted as a generalized “symmetry charge” associated with the submanifold $\gamma$.
Recall that $\int_{\gamma}*J_{e}$ counts the gauge fluxes encircled by $\gamma$; in $\left(3+1\right)$D, this quantity is given by the linking number between $\gamma$ and the worldsheets $\sigma$ of loop excitations. Equivalently, it can be computed by counting the intersections between $\gamma$ and a three-dimensional volume bounded by $\sigma$ in spacetime. In this sense, $J_{e}$ describes the flow of $2$-dimensional objects in $\left(3+1\right)$D spacetime, in close analogy with $j$ describing the flow of 0-dimensional particles. These 22d objects are precisely the worldsheets of loop excitations. The physical interpretation of $Q_{m}\left(\sigma\right)$ follows in a completely analogous manner.
Both $Q_{e}\left(\gamma\right)$ and $Q_{m}\left(\sigma\right)$ are invariant under smooth deformations of $\gamma$ and $\sigma$. To see this, consider a smooth deformation $\gamma_{1}\rightarrow\gamma_{2}$, which gives
$Q_{e}\left(\gamma_{1}\right)-Q_{e}\left(\gamma_{2}\right)=\int_{\partial\Sigma}*J_{e}=\int_{\Sigma}d\left(*J_{e}\right)$
with $\partial\Sigma=\gamma_{1}-\gamma_{2}$. By Stokes’ theorem and Eq. (8), we immediately obtain $Q_{e}\left(\gamma_{1}\right)=Q_{e}\left(\gamma_{2}\right)$. An identical argument applies to $Q_{m}\left(\sigma\right)$.
As in the case of ordinary global symmetries, we can define symmetry operators
$U_{q_{e}}\left(\gamma\right)=e^{{\rm i}q_{e}Q_{e}\left(\gamma\right)}=\exp\left({\rm i}q_{e}\int_{\gamma}a\right).$ (11)
The equations of motion imply that $\oint a\in\frac{2\pi}{N}\mathbb{Z}_{N}$, which in turn gives $NQ_{e}\left(\gamma\right)\in 2\pi\mathbb{Z}$, meaning that $q_{e}$ is equivalent to $q_{e}+N$. Furthermore, large gauge invariance under $\int_{\gamma}a\rightarrow\int_{\gamma}a+2\pi$ requires $q_{e}$ to be an integer. Consequently, $q_{e}$ takes values in $\mathbb{Z}/N\mathbb{Z}$. The symmetry operators satisfy
$U_{q_{e}}\left(\gamma\right)\times U_{q_{e}^{\prime}}\left(\gamma\right)=U_{q_{e}+q_{e}^{\prime}}\left(\gamma\right),\qquad U_{N}\left(\gamma\right)=1.$
Note that $U_{q_{e}}\left(\gamma\right)$ is supported on a 11d manifold $\gamma$ embedded in the 33d spatial manifold. In this sense, it implements a $\mathbb{Z}_{N}$ (31)(3-1)-form symmetry, i.e., a $2$-form symmetry, whose conserved current $J_{e}$ is a $3$-form.
Following the same logic, the equation of motion $\frac{N}{2\pi}db=0$ leads to another symmetry operator
$U_{q_{m}}\left(\sigma\right)=\exp\left({\rm i}q_{m}\int_{\sigma}b\right)$ (12)
with $q_{m}\in\mathbb{Z}/N\mathbb{Z}$. The operator $U_{q_{m}}\left(\sigma\right)$ implements a $\mathbb{Z}_{N}$ 11-form symmetry.
These two symmetry operators obey the commutation relation
$U_{q_{e}}\left(\gamma\right)U_{q_{m}}\left(\sigma\right)=e^{\frac{{\rm i}2\pi q_{e}q_{m}}{N}{\rm Lk}\left(\gamma,\sigma\right)}U_{q_{m}}\left(\sigma\right)U_{q_{e}}\left(\gamma\right),$ (13)
where ${\rm Lk}\left(\gamma,\sigma\right)$ is the linking number between $\gamma$ and $\sigma$. In this sense, $U_{q_{m}}\left(\sigma\right)$ carries charge under the symmetry generated by $U_{q_{e}}\left(\gamma\right)$, and vice versa.
This example illustrates that a $p$-form symmetry, much like a 0-form symmetry, originates from a conserved current. The essential difference is that the conserved quantity associated with a $p$-form symmetry is obtained by integrating the Hodge dual of the conserved current over a $\left(D-1-p\right)$-dimensional submanifold. Accordingly, the corresponding symmetry operator is a topological operator—protected by the conserved current—supported on a codimension-$p$ submanifold, i.e., a submanifold whose dimension is lower than that of the real space by $p$. Finally, we note that the $BF$ term in Eq. (6) can be generalized to a multi-component theory with a matrix-valued coefficient [41, 80]. In contrast to multi-component Chern–Simons theories, such a matrix can always be diagonalized via a combined general linear transformation.
III.2 Continuum topological field theory of Borromean-Rings topological order
The BR topological order [48, 49, 76, 79] can be effectively captured by a topological field theory with gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$, whose action takes the form
$S_{aab}=\int\sum_{i=1}^{3}\frac{N_{i}}{2\pi}b^{i}da^{i}+qa^{1}a^{2}b^{3},$ (14)
where $a^{i}$ and $b^{i}$ are $1$-form and $2$-form gauge fields, respectively. The first term consists of three conventional $BF$ terms, describing three decoupled $\mathbb{Z}_{N}$ ($N=N_{1},N_{2},N_{3}$) topological orders. The second term is a twisted term involving two $1$-form gauge fields ($a^{1},a^{2}$) and one $2$-form gauge field ($b^{3}$), which couples the three $\mathbb{Z}_{N}$ topological orders in a nontrivial manner. Its coefficient is $q=\frac{pN_{1}N_{2}N_{3}}{\left(2\pi\right)^{2}N_{123}}$, where $N_{123}$ denotes the greatest common divisor (GCD) of $N_{1},N_{2},N_{3}$, and $p\in\mathbb{Z}_{N_{123}}$ labels the level of the twisted term.
The three $BF$ terms impose flatness conditions on $a^{1},a^{2},b^{3}$, leading to the gauge transformations $a^{1,2}\rightarrow a^{1,2}+d\chi^{1,2}$ and $b^{3}\rightarrow b^{3}+dV^{3}$. In contrast, the twisted term $a^{1}a^{2}b^{3}$ induces nontrivial shift contributions to the gauge transformations of $b^{1},b^{2}$, and $a^{3}$. Explicitly, these gauge transformations take the form
$a^{3}\rightarrow$ $a^{3}+d\chi^{3}-\frac{2\pi q}{N_{3}}\left(\chi^{1}a^{2}+\frac{1}{2}\chi^{1}d\chi^{2}\right)$
$+\frac{2\pi q}{N_{3}}\left(\chi^{2}a^{1}+\frac{1}{2}\chi^{2}d\chi^{1}\right),$ (15)
$b^{1}\rightarrow$ $b^{1}+dV^{1}-\frac{2\pi q}{N_{1}}\left(\chi^{2}b^{3}-a^{2}V^{3}+\chi^{2}dV^{3}\right),$ (16)
$b^{2}\rightarrow$ $b^{2}+dV^{2}+\frac{2\pi q}{N_{2}}\left(\chi^{1}b^{3}-a^{1}V^{3}+\chi^{1}dV^{3}\right).$ (17)
To identify the conserved currents and the associated symmetry operators, we first derive the equations of motion. The EoMs are obtained by requiring the action to remain invariant (modulo $2\pi$) under infinitesimal variations of the gauge fields. For instance, under $b^{1}\rightarrow b^{1}+\delta b^{1}$, the condition $\delta S\left[\delta b^{1}\right]=0\mod{2\pi}$ implies $\frac{N_{1}}{2\pi}da^{1}=0$, which is the equation of motion for $a^{1}$. The full set of EoMs, together with the corresponding conserved currents and symmetry operators of $S_{aab}$, is summarized in Table 1. Among these symmetry operators, some are invertible whereas others are non-invertible. In particular, due to the presence of the twisted term, two distinct classes of EoMs and conserved currents emerge, which we refer to as type-I and type-II, respectively, and which will be discussed in detail in the following subsections.
Table 1: Equations of motion (EoMs), conserved currents, and symmetry operators of $S_{aab}=\int\sum_{i=1}^{3}\frac{N_{i}}{2\pi}b^{i}da^{i}+qa^{1}a^{2}b^{3}$, where $q=\frac{pN_{1}N_{2}N_{3}}{\left(2\pi\right)^{2}N_{123}}$ with $p\in\mathbb{Z}_{N_{123}}$. The first three symmetry operators are invertible and correspond to Wilson operators of Abelian particle and Abelian loop excitations. Specifically, $U_{e_{1}}\left(\gamma\right)$, $U_{e_{2}}\left(\gamma\right)$, and $U_{e_{3}}\left(\sigma\right)$ implement $\mathbb{Z}_{N_{1}}^{\left(2\right)}$, $\mathbb{Z}_{N_{2}}^{\left(2\right)}$, and $\mathbb{Z}_{N_{3}}^{\left(1\right)}$ symmetries, respectively, where the superscripts $\left(2\right)$ and $\left(1\right)$ denote 22-form and $1$-form symmetries. In contrast, the last three symmetry operators, $L^{1}\left(\sigma\right)$, $L^{2}\left(\sigma\right)$, and $L^{3}\left(\gamma\right)$, are non-invertible. They take the same functional form as Wilson operators associated with non-Abelian loop and non-Abelian particle excitations, up to overall normalization factors. The delta functions appearing in these symmetry operators enforce constraints on gauge-field configurations, under which the corresponding currents are conserved and define genuine symmetries (see Sec. III.4). The presence of these delta functions renders the symmetry operators non-invertible. Fusion of the symmetry operators listed here generally produces new symmetry operators; see Sec. IV for a detailed discussion.
EoM Currents Continuity Eq. Symmetry operators
$\frac{N_{1}}{2\pi}da^{1}=0$ $J_{e}^{1}=-*a^{1}$ $d\left(*J_{e}^{1}\right)=0$ $U_{e_{1}}\left(\gamma\right)=\exp\left({\rm i}e_{1}\int_{\gamma}a^{1}\right)$
$\frac{N_{2}}{2\pi}da^{2}=0$ $J_{e}^{2}=-*a^{2}$ $d\left(*J_{e}^{2}\right)=0$ $U_{e_{2}}\left(\gamma\right)=\exp\left({\rm i}e_{2}\int_{\gamma}a^{2}\right)$
$\frac{N_{3}}{2\pi}db^{3}=0$ $J_{e}^{3}=*b^{3}$ $d\left(*J_{e}^{3}\right)=0$ $U_{e_{3}}\left(\sigma\right)=\exp\left({\rm i}e_{3}\int_{\sigma}b^{3}\right)$
$\frac{N_{1}}{2\pi}db^{1}+qa^{2}b^{3}=0$ $J_{m}^{1}=*\left(b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}\right)$ $d\left(*J_{m}^{1}\right)=0$ $\begin{array}[]{cc}L^{1}\left(\sigma\right)=&\exp\left({\rm i}\int_{\sigma}b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}\right)\\ &\times\delta\left(\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}\right)\delta\left(\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}\right)\end{array}$
$\frac{N_{2}}{2\pi}db^{2}-qa^{1}b^{3}=0$ $J_{m}^{2}=*\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right)$ $d\left(*J_{m}^{2}\right)=0$ $\begin{array}[]{cc}L^{2}\left(\sigma\right)=&\exp\left({\rm i}\int_{\sigma}b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right)\\ &\times\delta\left(\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}\right)\delta\left(\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}\right)\end{array}$
$\frac{N_{3}}{2\pi}da^{3}+qa^{1}a^{2}=0$ $J_{m}^{3}=-*\left(a^{3}+\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\phi^{1}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi^{2}\right)$ $d\left(*J_{m}^{3}\right)=0$ $\begin{array}[]{cc}L^{3}\left(\gamma\right)=&\exp\left({\rm i}\int_{\gamma}a^{2}+\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\phi^{1}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi^{2}\right)\\ &\times\delta\left(\int_{\gamma}\frac{pN_{1}}{N_{123}}a^{1}\right)\delta\left(\int_{\gamma}\frac{pN_{2}}{N_{123}}a^{2}\right)\end{array}$
III.3 Type-I conserved currents and invertible symmetry operators
We begin by considering a local perturbation of the gauge field $b^{1}$. Under $b^{1}\rightarrow b^{1}+\delta b^{1}$, the variation of the action is $\delta S\left[\delta b^{1}\right]=\int_{M^{4}}\frac{N_{1}}{2\pi}\delta b^{1}da^{1}$. Requiring $\delta S\left[\delta b^{1}\right]=0$ for an arbitrary infinitesimal perturbation $\delta b^{1}$, we obtain the corresponding equation of motion
$\frac{N_{1}}{2\pi}da^{1}=0.$ (18)
This EoM implies flux quantization, $\oint_{\gamma}a^{1}=\frac{2\pi k_{1}}{N_{1}}$ with $k_{1}\in\mathbb{Z}_{N_{1}}$; namely, the flux of $a^{1}$ piercing the surface enclosed by $\gamma$ is quantized in integer multiples of $\frac{2\pi}{N_{1}}$. Since this EoM takes the form of a total derivative set to zero, it naturally resembles a continuity equation. Indeed, from Eq. (18) we can extract the continuity equation
$d\left(*J^{1}_{e}\right)=0\quad\text{with}\quad J^{1}_{e}=-*a^{1},$ (19)
where $J^{1}_{e}$ is a $3$-form conserved current. We refer to this class of EoMs and conserved currents as type-I.
Analogous to the case of an ordinary 0-form symmetry, we define a conserved quantity
$Q_{1,\gamma}=\int_{\gamma}*J^{1}_{e}=\int_{\gamma}a^{1},$ (20)
which depends only on the topological class of the closed curve $\gamma$. To see that $Q_{1,\gamma}$ is invariant under smooth deformations of $\gamma$, consider $Q_{1,\gamma}-Q_{1,\gamma^{\prime}}=\int_{\gamma-\gamma^{\prime}}*J^{1}_{e}=\int_{\Sigma}d\left(*J^{1}_{e}\right)=0$, where $\Sigma$ is the surface bounded by $\gamma-\gamma^{\prime}$ and we have used Stokes’ theorem. This invariance confirms the topological nature of $Q_{1,\gamma}$.
In direct analogy with ordinary global symmetries, we can define the associated symmetry operator
$U_{e_{1}}\left(\gamma\right)=\exp\left({\rm i}e_{1}\int_{\gamma}a^{1}\right).$ (21)
The object charged under this symmetry is the closed worldsheet of the loop excitation described by $b^{1}$. Correspondingly, $Q_{1,\gamma}$ counts the number of intersections between $\gamma$ and the three-dimensional volume bounded by the loop worldsheet. We therefore identify $U_{e_{1}}\left(\gamma\right)$ as a Wilson operator. Imposing large gauge invariance together with the quantization of $\int_{\gamma}a^{1}$, we find that $e_{1}$ is an integer defined modulo $N_{1}$. As a result, the symmetry operator $U_{e_{1}}\left(\gamma\right)$ with $e_{1}\in\left\{0,1,\cdots,N_{1}-1\right\}$ implements a $\mathbb{Z}_{N_{1}}$ 22-form symmetry.
Following the same reasoning, from the equations of motion for the gauge fields $a^{2}$ and $b^{3}$,
$\frac{N_{2}}{2\pi}da^{2}=0,\quad\frac{N_{3}}{2\pi}db^{3}=0,$ (22)
we identify the associated conserved currents
$J_{e}^{2}=-*a^{2},\quad J_{e}^{3}=*b^{3}.$ (23)
The corresponding symmetry operators are
$U_{e_{2}}\left(\gamma\right)=\exp\left({\rm i}e_{2}\int_{\gamma}a^{2}\right),$ (24)
$U_{e_{3}}\left(\sigma\right)=\exp\left({\rm i}e_{3}\int_{\sigma}b^{3}\right),$ (25)
where $e_{2}\in\left\{1,2,\cdots,N_{2}-1\right\}$ and $e_{3}\in\left\{1,2,\cdots,N_{3}-1\right\}$.
We note that the symmetry operators $U_{e_{1}}\left(\gamma\right)$, $U_{e_{2}}\left(\gamma\right)$, and $U_{e_{3}}\left(\sigma\right)$ coincide with the Wilson operators of Abelian particle and loop excitations. The operator $U_{e_{2}}\left(\gamma\right)$ implements a $\mathbb{Z}_{N_{2}}$ 22-form symmetry, while $U_{e_{3}}\left(\sigma\right)$ implements a $\mathbb{Z}_{N_{3}}$ 11-form symmetry, in the sense that $U_{e_{i}}\times U_{e_{i}^{\prime}}=U_{e_{i}+e_{i}^{\prime}}$ and $U_{N_{i}}=1$. Thus far, we have identified all invertible symmetry operators in the effective field theory description of the BR topological order.
III.4 Type-II conserved currents and non-invertible symmetry operators
Besides the type-I equations of motion discussed above, the theory also admits another class of EoMs. As an example, consider a local variation of $a^{1}$, $a^{1}\rightarrow a^{1}+\delta a^{1}$. The variation of the action is $\delta S\left[\delta a^{1}\right]=\int_{M^{4}}\frac{N_{1}}{2\pi}\delta a^{1}db^{1}+q\delta a^{1}a^{2}b^{3}$. Requiring $\delta S\left[\delta a^{1}\right]=0\mod{2\pi}$ yields
$\frac{N_{1}}{2\pi}db^{1}+qa^{2}b^{3}=0.$ (26)
In contrast to the EoMs (18) and (22), the EoM (26) is not simply a vanishing total derivative. We refer to EoMs of the form (26) as type-II, to distinguish them from the type-I EoMs in Eqs. (18) and (22).
A natural question is whether one can still extract a continuity equation from the type-II EoM (26). The answer is affirmative, but only under additional conditions. The key step is to rewrite the EoM into the form of a vanishing total derivative. To this end, we introduce two auxiliary fields $\phi^{2}$ and $\beta^{3}$ such that $d\phi^{2}=\frac{pN_{2}}{N_{123}}a^{2}$ and $d\beta^{3}=\frac{pN_{3}}{N_{123}}b^{3}$. We notice that $d\phi^{2}$ ($d\beta^{3}$) is an integral multiple of $a^{2}$ ($b^{3}$). Therefore, the fields $\phi^{2}$ and $\beta^{3}$ satisfy the Dirac quantization condition $\oint d\phi^{2}\in 2\pi\mathbb{Z}$ and $\oint d\beta^{3}\in 2\pi\mathbb{Z}$ when $\oint a^{2}$ and $\oint b^{3}$ take specific values. With $\phi^{2}$ and $\beta^{3}$, the EoM (26) can be rewritten as
$\frac{N_{1}}{2\pi}\left(db^{1}+\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}d\beta^{3}+\frac{1}{2}d\phi^{2}\frac{N_{3}}{2\pi}b^{3}\right)=0.$ (27)
Here, the two coefficients $\frac{1}{2}$ are not the result of the gauge invariance: the above equation still holds if one chooses two coefficient $c_{1}$ and $c_{2}$ with $c_{1}+c_{2}=1$. Instead, we split the term $\frac{2\pi qa^{2}b^{3}}{N_{1}}$ in a symmetric way in order to respect the fact that the twisted term in the action (14) is symmetric under exchanging $a^{1}$ and $a^{2}$, up to an overall anti-symmetry of the wedge product. In this way, the currents extracted from the EoMs have the same symmetric formalism, see Table 1. Using $\frac{N_{2}}{2\pi}da^{2}=0$ and $\frac{N_{3}}{2\pi}db^{3}=0$, we further obtain
$\frac{N_{1}}{2\pi}d\left(b^{1}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}\right)=0,$ (28)
from which we can extract the continuity equation
$d\left(*J_{m}^{1}\right)=0,\ J_{m}^{1}=*\left(b^{1}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}\right),$ (29)
where $J^{1}_{m}$ is a $2$-form current. Analogous to ordinary symmetries, this conservation law implies a conserved quantity
$Q_{m}^{1}=\int_{\sigma}*J_{m}^{1}=\int_{\sigma}b^{1}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3},$ (30)
where $\sigma$ is a closed 22d submanifold. The value of $Q_{m}^{1}$ is $\frac{2\pi k}{N_{1}}$ with $k\in\left\{0,1,\cdots,N_{1}-1\right\}$.
One might then expect that $Q_{m}^{1}$ generates a symmetry via the unitary operator
$\mathcal{U}_{1}=e^{{\rm i}Q_{m}^{1}}=\exp\left({\rm i}\int_{\sigma}b^{1}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}\right).$ (31)
However, an important subtlety arises: while the EoM (26) holds identically, the continuity equation (29)—and thus the conservation of $J^{1}_{m}$—is valid only when $d\phi^{2}=\frac{pN_{2}}{N_{123}}a^{2}$ and $d\beta^{3}=\frac{pN_{3}}{N_{123}}b^{3}$ are satisfied. Therefore, for $Q_{m}^{1}$ to be conserved, these relations must hold at least on the submanifold $\sigma$. Equivalently, these conditions can be expressed as $\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}=0$ and $\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}=0$, because on the closed manifold $\sigma$ one has $\int_{\sigma}d\beta^{3}=0$, and $\int_{\gamma}d\phi^{2}=0$ for any closed curve $\gamma\subset\sigma$. In other words, although $\mathcal{U}_{1}$ is a unitary operator originating from $Q_{m}^{1}$, it defines a symmetry operation only when $\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}=0$ and $\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}=0$ are satisfied. Accordingly, we define the symmetry operator as
$L^{1}\left(\sigma\right)=\mathcal{U}_{1}\delta\left(\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}\right)\delta\left(\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}\right).$ (32)
Here the delta functions are defined as $\delta\left(x\right)=1$ if and only if $x=0\mod 2\pi$, and $\delta\left(x\right)=0$ otherwise. These delta functions act as projectors onto the subspace characterized by $\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}=0$ and $\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}=0$.
When the symmetry operator $L^{1}\left(\sigma\right)$ is inserted into a correlation function, it first projects onto the subspace of gauge-field configurations satisfying the constraints encoded by the delta functions. Within this constrained subspace, $\mathcal{U}_{1}$ implements a unitary operation. Because $L^{1}\left(\sigma\right)$ contains projectors, it has no inverse: there is no operator $O$ such that $OL^{1}\left(\sigma\right)$ equals the identity. In this sense, $L^{1}\left(\sigma\right)$ is a non-invertible symmetry operator.
We further note that $L^{1}\left(\sigma\right)$ closely resembles the Wilson operator describing a loop excitation carrying one unit of $\mathbb{Z}_{N_{1}}$ gauge flux [76], up to an overall normalization factor. With this Wilson-operator picture in mind, the symmetry operator admits the following interpretation. Inserting such an operator creates a pair of loop excitations, transports one of them along $\sigma$, and finally annihilates the pair. During this process, the operator measures how many $\mathbb{Z}_{N_{1}}$ gauge charges (carried by particle excitations) are enclosed in the volume bounded by $\sigma$. This is analogous to the 22d toric code, where moving an $m$ excitation along a closed path measures the number of enclosed $e$ excitations. The enclosed number of $\mathbb{Z}_{N_{1}}$ gauge charges, $n_{e_{1}}$, is reflected by the phase factor $\exp\left({\rm i}\frac{2\pi}{N_{1}}\times n_{e_{1}}\right)$. Since the loop carries one unit of gauge flux, the measurement outcome is defined modulo $N_{1}$. The “symmetry” here is reflected in the fact that $\sigma$ can be smoothly deformed without changing $Q_{m}^{1}$, but only within the subspace where $\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}=0$ and $\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}=0$. Outside this subspace, the symmetry need not exist. In this way, we establish a direct connection between a non-Abelian loop excitation and a non-invertible symmetry.
Following this line of reasoning, we may ask what symmetry operator is associated with a loop excitation carrying multiple units, say $m_{1}$ units, of $\mathbb{Z}_{N_{1}}$ gauge flux. When the flux is not minimal, the operator may fail to distinguish certain gauge-charge numbers. For instance, in a 22d $\mathbb{Z}_{N=6}$ toric code model, moving a $2m$ particle along a closed path yields a phase factor $\exp\left({\rm i}\frac{2\pi}{6}\times 2\times n_{e}\right)$, where $n_{e}$ is the number of enclosed $e$ excitations. This implies that $n_{e}=1$ is indistinguishable from $n_{e}=4$, i.e., $n_{e}$ is identified with $n_{e}+3$. When a loop carrying $m_{1}$ units of gauge flux is transported along $\sigma$, the observable phase is quantized in units of $\frac{2\pi m_{1}}{N_{1}}$ rather than $\frac{2\pi}{N_{1}}$. Correspondingly, the conserved quantity $Q_{m_{1}}^{1}$, interpreted as the “effective” gauge-charge number detected by the loop, is defined in units of $\frac{2\pi m_{1}}{N_{1}}$. This can be captured by a current $J_{m_{1}}^{1}$ satisfying
$d\left(*J_{m_{1}}^{1}\right)=m_{1}\cdot d\left(*J_{m}^{1}\right)=0,$ (33)
such that $Q_{m_{1}}^{1}=\int_{\sigma}*J_{m_{1}}^{1}=\frac{2\pi m_{1}}{N_{1}}\cdot k$ with $k\in\mathbb{Z}$. Such a conserved current $J_{m_{1}}^{1}$ exists if we introduce two auxiliary fields $\phi_{m_{1}}^{2}$ and $\beta_{m_{1}}^{3}$ with $d\phi_{m_{1}}^{2}=\frac{m_{1}pN_{2}}{N_{123}}a^{2}$ and $d\beta_{m_{1}}^{3}=\frac{m_{1}pN_{3}}{N_{123}}b^{3}$. The conserved current is
$J_{m_{1}}^{1}=m_{1}b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta_{m_{1}}^{3}+\frac{1}{2}\phi_{m_{1}}^{2}\frac{N_{3}}{2\pi}b^{3}.$ (34)
We may view $J_{m_{1}}^{1}$ as arising from combining $m_{1}$ copies of $J_{m}^{1}$. The corresponding symmetry operator is
$L_{m_{1}}^{1}\left(\sigma\right)$ $=\mathcal{U}_{m_{1}}\delta\left(\int_{\sigma}\frac{m_{1}pN_{3}}{N_{123}}b^{3}\right)$
$\times\delta\left(\int_{\gamma\subset\sigma}\frac{m_{1}pN_{2}}{N_{123}}a^{2}\right),$ (35)
with
$\mathcal{U}_{m_{1}}=e^{{\rm i}Q^{1}_{m_{1}}}=e^{{\rm i}\left(\int_{\sigma}m_{1}b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta_{m_{1}}^{3}+\frac{1}{2}\phi_{m_{1}}^{2}\frac{N_{3}}{2\pi}b^{3}\right)}.$ (36)
The delta functions in $L^{1}_{m_{1}}\left(\sigma\right)$ project onto a larger subspace than those in $L^{1}\left(\sigma\right)$, in the sense that more gauge-field configurations are allowed. Thus, $\mathcal{U}_{m_{1}}$ implements a unitary action on a larger subspace. In the limiting case, $\mathcal{U}_{N_{1}}=1$ is a symmetry operation on the full Hilbert space (here the Hilbert space is spanned by configurations of the gauge fields). Again, $L_{m_{1}}^{1}\left(\sigma\right)$ is non-invertible because it is accompanied by two projectors. We also note that $L_{m_{1}}^{1}\left(\sigma\right)$ has the same mathematical form as the Wilson operator of a loop excitation carrying $m_{1}$ units of $\mathbb{Z}_{N_{1}}$ gauge flux [76], up to an overall normalization factor. Since $L_{m_{1}}^{1}\left(\sigma\right)$ acts on states in the Hilbert space, any overall coefficient can be absorbed into the normalization of the state.
In the same manner, we can obtain the EoM of gauge field $b^{2}$,
$\frac{N_{2}}{2\pi}db^{2}-qa^{1}b^{3}=0.$ (37)
From it we can extract a type-II conserved current, i.e., a continuity equation,
$d\left(*J_{m}^{2}\right)=0,\ J_{m}^{2}=*\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right),$ (38)
where the auxiliary fields $\phi^{1}$ and $\beta^{3}$ satisfy $d\phi^{1}=\frac{pN_{1}}{N_{123}}a^{1}$ and $d\beta^{3}=\frac{pN_{3}}{N_{123}}b^{3}$. The conserved quantity is $Q_{m}^{2}=\int_{\sigma}*J_{m}^{2}$, which leads to the symmetry operator
$L^{2}\left(\sigma\right)$ $=\mathcal{U}_{2}\delta\left(\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}\right)\delta\left(\int_{\gamma\subset\sigma}\frac{pN_{1}}{N_{123}}a^{1}\right),$ (39)
with $\mathcal{U}_{2}=\exp\left[{\rm i}\left(\int_{\sigma}b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right)\right].$
By combining $m_{2}$ copies of the current $J_{m}^{2}$, we obtain another type-II conserved current
$J_{m_{2}}^{2}=*\left(m_{2}b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta_{m_{2}}^{3}-\frac{1}{2}\phi_{m_{2}}^{1}\frac{N_{3}}{2\pi}b^{3}\right),$ (40)
provided that $d\phi_{m_{2}}^{1}=\frac{m_{2}pN_{1}}{N_{123}}a^{1}$ and $d\beta_{m_{2}}^{3}=\frac{m_{2}pN_{3}}{N_{123}}b^{3}$. The corresponding symmetry operator is
$L_{m_{2}}^{2}\left(\sigma\right)=\mathcal{U}_{m_{2}}\delta\left(\int_{\sigma}\frac{m_{2}pN_{3}}{N_{123}}b^{3}\right)\delta\left(\int_{\gamma\subset\sigma}\frac{m_{2}pN_{2}}{N_{123}}a^{2}\right),$ (41)
with
$\mathcal{U}_{m_{2}}=$ $e^{{\rm i}\int_{\sigma}*J_{m_{2}}^{2}}=e^{{\rm i}\left(\int_{\sigma}m_{2}b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta_{m_{2}}^{3}-\frac{1}{2}\phi_{m_{2}}^{1}\frac{N_{3}}{2\pi}b^{3}\right)}.$ (42)
$L_{m_{2}}^{2}\left(\sigma\right)$ is a non-invertible $1$-form symmetry operator, and it corresponds to the non-Abelian loop excitation carrying $m_{2}$ units of $\mathbb{Z}_{N_{2}}$ gauge flux.
The above discussion also applies to the EoM of gauge field $a^{3}$,
$\frac{N_{3}}{2\pi}da^{3}+qa^{1}a^{2}=0.$ (43)
We can similarly extract a type-II conserved current given by
$d\left(*J_{m}^{3}\right)=0,\ J_{m}^{3}=-*\left(a^{3}+\frac{1}{2}\phi^{1}\frac{N_{2}}{2\pi}a^{2}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi^{2}\right),$ (44)
when $d\phi^{1}=\frac{pN_{1}}{N_{123}}a^{1}$ and $d\phi^{2}=\frac{pN_{2}}{N_{123}}a^{2}$ are satisfied. The corresponding symmetry operator is
$L^{3}\left(\gamma\right)=\mathcal{U}_{3}\delta\left(\int_{\gamma}\frac{pN_{1}}{N_{123}}a^{1}\right)\delta\left(\int_{\gamma}\frac{pN_{2}}{N_{123}}a^{2}\right),$ (45)
where $\mathcal{U}_{3}=\exp\left[{\rm i}\left(\int_{\gamma}a^{3}+\frac{1}{2}\phi^{1}\frac{N_{2}}{2\pi}a^{2}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi^{2}\right)\right]$.
Combining several copies of $J_{m}^{3}$ yields a new type-II conserved current
$J_{m_{3}}^{3}=-*\left(m_{3}a^{3}+\frac{1}{2}\phi_{m_{3}}^{1}\frac{N_{2}}{2\pi}a^{2}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi_{m_{3}}^{2}\right),$ (46)
where the auxiliary fields $\phi_{m_{3}}^{1}$ and $\phi_{m_{3}}^{2}$ satisfy $d\phi_{m_{3}}^{1}=\frac{m_{3}pN_{1}}{N_{123}}a^{1}$ and $d\phi_{m_{3}}^{2}=\frac{m_{3}pN_{2}}{N_{123}}a^{2}$. The corresponding symmetry operator is
$L_{m_{3}}^{3}\left(\gamma\right)=\mathcal{U}_{m_{3}}\delta\left(\int_{\gamma}\frac{m_{3}pN_{1}}{N_{123}}a^{1}\right)\delta\left(\int_{\gamma}\frac{m_{3}pN_{2}}{N_{123}}a^{2}\right),$ (47)
with
$\mathcal{U}_{m_{3}}=$ $e^{{\rm i}\int_{\gamma}*J_{m_{3}}^{3}}=e^{{\rm i}\left(\int_{\gamma}m_{3}a^{2}+\frac{1}{2}\phi_{m_{3}}^{1}\frac{N_{2}}{2\pi}a^{2}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi_{m_{3}}^{2}\right)}.$ (48)
$L_{m_{3}}^{3}\left(\gamma\right)$ is a non-invertible $2$-form symmetry operator and corresponds to the non-Abelian particle excitation carrying $m_{3}$ units of $\mathbb{Z}_{N_{3}}$ gauge charge.
Before ending this section, we remark that inserting $L^{1}_{m_{1}}\left(\sigma\right)$ into a correlation function is not simply equivalent to inserting $m_{1}$ copies of $L^{1}\left(\sigma\right)$. This distinction will become clear in the next section, where we discuss fusion rules of symmetry operators. For completeness, the EoMs, conserved currents, corresponding continuity equations, and symmetry operators are summarized in Table 1.
IV Fusion rules of symmetry operators
In this section, we investigate the fusion rules of the symmetry operators identified in Sec. III. Fusion rules characterize the outcome of consecutively applying two symmetry operators supported on the same submanifold, and they encode the underlying algebraic structure of a generalized symmetry. From a field-theoretical perspective, fusion is realized by inserting symmetry operators sequentially into correlation functions.
For invertible symmetries, fusion rules are group-like: each symmetry operator is labeled by a group element, and their fusion follows the group multiplication law. By contrast, for non-invertible symmetries, fusion rules are intrinsically non-group-like. In particular, the fusion of two non-invertible symmetry operators can produce a linear combination of several symmetry operators, rather than a single one.
To analyze fusion rules in a systematic manner, it is crucial to first identify the complete set of symmetry operators in the theory. In Sec. IV.1, we show that additional symmetry operators can be generated from new conserved currents, which themselves arise from linear combinations of existing conserved currents. This construction elucidates the general structure of symmetry operators relevant for fusion.
In Sec. IV.2, we specialize to the simplest nontrivial case by choosing the gauge group in the action (14) to be $G=(\mathbb{Z}_{2})^{3}$. This choice allows us to explicitly demonstrate representative examples of fusion rules and to highlight characteristic features of non-invertible symmetries in a concrete and transparent setting. Guided by these examples, we present a general expression for fusion rules in Sec. IV.3, together with a practical prescription for computing fusion outcomes in a generic setup. Finally, in Sec. IV.4, we provide additional examples of fusion rules for different choices of gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$, illustrating how the fusion structure depends on the underlying gauge data.
IV.1 Generating new conserved currents by linear combinations
Since each conserved current corresponds to a symmetry operator, it is natural to combine multiple conserved currents to construct new ones, which in turn correspond to new symmetry operators. From the preceding discussion, we obtain the following “minimal” continuity equations:
$d\left(*J_{e}^{1}\right)=da^{1}=0,\ d\left(*J_{e}^{2}\right)=da^{2}=0,\ d\left(*J_{e}^{3}\right)=db^{3}=0,$ (49)
$d\left(*J_{m}^{1}\right)=db^{1}+\frac{pN_{2}N_{3}}{2\pi N_{123}}a^{2}b^{3}=0,$ (50)
$d\left(*J_{m}^{2}\right)=db^{2}-\frac{pN_{1}N_{3}}{2\pi N_{123}}a^{1}b^{3}=0,$ (51)
$d\left(*J_{m}^{3}\right)=da^{3}+\frac{pN_{1}N_{2}}{2\pi N_{123}}a^{1}a^{2}=0.$ (52)
New continuity equations can be obtained by taking linear combinations of these minimal ones. This viewpoint provides useful intuition for understanding the fusion of symmetry operators. In this way, we can systematically exhaust all symmetry operators allowed by the theory.
We first consider combinations of the type-I currents. It is straightforward to verify that
$e_{1}\cdot d\left(*J_{e}^{1}\right)=e_{1}\cdot da^{1}=d\left(e_{1}a^{1}\right)=0$ (53)
is also a valid continuity equation, where $e_{1}$ is an integer-valued coefficient. The parameter $e_{1}$ is defined modulo $N_{1}$, since $N_{1}*J_{e}^{1}$ is trivial. This leads to a new conserved current $J_{e_{1}}$ given by
$J_{e_{1}}=e_{1}a^{1}$ (54)
and a corresponding conserved quantity $Q_{e_{1}}=\int_{\gamma}*J_{e_{1}}=\int_{\gamma}e_{1}a^{1}$. As a result, we obtain a symmetry operator
$U_{e_{1}}\left(\gamma\right)=\exp\left({\rm i}\int_{\gamma}e_{1}a^{1}\right).$ (55)
We immediately recognize that $U_{e_{1}}\left(\gamma\right)$ represents the worldline operator of a particle excitation carrying $e_{1}$ units of $\mathbb{Z}_{N_{1}}$ gauge charge. Operationally, $U_{e_{1}}\left(\gamma\right)$ measures the number of $\mathbb{Z}_{N_{1}}$ gauge fluxes encircled by $\gamma$, denoted by $n_{m_{1}}$, which manifests itself as a phase factor $e^{i\theta_{e_{1}}}=\exp\left(\frac{2\pi e_{1}}{N_{1}}\times n_{m_{1}}\right)$. The associated symmetry implies that this phase remains invariant under smooth deformations of $\gamma$, namely, deformations that do not cross loops carrying $\mathbb{Z}_{N_{1}}$ gauge flux.
We can also combine different types of currents. A general combination of the type-I currents takes the form
$*J_{e_{1},e_{2},e_{3}}=$ $\delta^{\perp}\left(\gamma_{p}\right)\wedge e_{1}*J_{e}^{1}+\delta^{\perp}\left(\gamma_{p}\right)\wedge e_{2}*J_{e}^{2}$
$+e_{3}*J_{e}^{3},$ (56)
and the corresponding symmetry operator is given by
$U_{e_{1}e_{2}e_{3}}\left(\sigma,\gamma_{p}\right)=\exp\left({\rm i}\int_{\gamma_{p}}e_{1}a^{1}+{\rm i}\int_{\gamma_{p}}e_{2}a^{2}+{\rm i}\int_{\sigma}e_{3}b^{3}\right).$ (57)
Since $J_{e}^{1}$ and $J_{e}^{2}$ are $1$-forms, they must be promoted to $2$-forms, for instance in the form $\delta^{\perp}\left(\gamma_{p}\right)\wedge*J_{e}^{1}$, in order to be combined consistently with the $2$-form current $J_{e}^{3}$. Here, the $1$-form $\delta^{\perp}\left(\gamma_{p}\right)$ is a delta distribution supported on a closed curve $\gamma_{p}$, satisfying $\int_{\sigma}\delta^{\perp}\left(\gamma_{p}\right)\wedge(\cdots)=\int_{\gamma_{p}}(\cdots)$. Physically, $\gamma_{p}$ admits a natural interpretation: the operator $U_{e_{1},e_{2},e_{3}}$ represents the Wilson operator of a loop excitation decorated by particle excitations, where $\sigma$ is the worldsheet of the loop and $\gamma_{p}\subset\sigma$ is the worldline of the particle. It is straightforward to verify that $U_{e_{1},e_{2},e_{3}}\left(\sigma,\gamma_{p}\right)$ is an invertible symmetry operator.
It is more instructive to consider combinations involving type-II currents. As a representative example, starting from
$m_{1}\cdot d\left(*J_{m}^{1}\right)+m_{2}\cdot d\left(*J_{m}^{2}\right)=0,$ (58)
we obtain
$m_{1}\cdot db^{1}+\frac{m_{1}pN_{2}N_{3}}{2\pi N_{123}}a^{2}b^{3}+m_{2}\cdot db^{2}-\frac{m_{2}pN_{1}N_{3}}{2\pi N_{123}}a^{1}b^{3}=0.$ (59)
We therefore anticipate the existence of a new conserved current $J_{m_{1},m_{2}}$ satisfying $d\left(*J_{m_{1},m_{2}}\right)=m_{1}\cdot d\left(*J_{m}^{1}\right)+m_{2}\cdot d\left(*J_{m}^{2}\right)=0$. A naive choice, $J_{m_{1},m_{2}}=m_{1}*J_{m}^{1}+m_{2}*J_{m}^{2}$, indeed yields a conserved current, but it simply inherits all constraints required for the conservation of $J_{m}^{1}$ and $J_{m}^{2}$ individually. Remarkably, the presence of twisted terms allows the type-II currents to be nontrivially intertwined, making it possible to construct a new conserved current subject to weaker constraints. Indeed, one can explicitly verify that
$m_{1}\cdot d\left(*J_{m}^{1}\right)+m_{2}\cdot d\left(*J_{m}^{2}\right)=0=$ $\ d\left(m_{1}b^{1}\right)+d\left(m_{2}b^{2}\right)+\frac{1}{2}\frac{m_{1}pN_{2}}{N_{123}}a^{2}\frac{N_{3}}{2\pi}b^{3}+\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\frac{m_{1}pN_{3}}{N_{123}}b^{3}$
$-\frac{1}{2}\frac{m_{2}pN_{1}}{N_{123}}a^{1}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\frac{m_{2}pN_{3}}{N_{123}}b^{3}.$
Once we introduce several auxiliary fields satisfying $d\phi_{m_{1},m_{2}}^{1,2}=\frac{m_{1}pN_{2}}{N_{123}}a^{2}-\frac{m_{2}pN_{1}}{N_{123}}a^{1}$, $d\beta_{m_{1}}^{3}=\frac{m_{1}pN_{3}}{N_{123}}b^{3}$, and $d\beta_{m_{2}}^{3}=\frac{m_{2}pN_{3}}{N_{123}}b^{3}$, the above equation can be rewritten as
$0=$ $d\left(m_{1}b^{1}\right)+d\left(m_{2}b^{2}\right)+\frac{1}{2}d\phi_{m_{1},m_{2}}^{1,2}\frac{N_{3}}{2\pi}b^{3}$
$+\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}d\beta_{m_{1}}^{3}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}d\beta_{m_{2}}^{3}.$ (60)
We may therefore interpret this equation as $0=d\left(*J_{m_{1},m_{2}}\right)$ and identify a new conserved current $J_{m_{1},m_{2}}$ given by
$*J_{m_{1},m_{2}}=$ $m_{1}b^{1}+m_{2}b^{2}+\frac{1}{2}\phi_{m_{1},m_{2}}^{1,2}\frac{N_{3}}{2\pi}b^{3}$
$-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta_{m_{1}}^{3}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta_{m_{2}}^{3},$ (61)
provided that $d\phi_{m_{1},m_{2}}^{1,2}=\frac{m_{1}pN_{2}}{N_{123}}a^{2}-\frac{m_{2}pN_{1}}{N_{123}}a^{1}$, $d\beta_{m_{1}}^{3}=\frac{m_{1}pN_{3}}{N_{123}}b^{3}$, and $d\beta_{m_{2}}^{3}=\frac{m_{2}pN_{3}}{N_{123}}b^{3}$ are satisfied. For comparison, the conservation of $J^{1}_{m_{1}}$ requires $d\phi^{2}_{m_{1}}=\frac{m_{1}pN_{2}}{N_{123}}a^{2}$, while that of $J^{2}_{m_{2}}$ requires $d\phi^{1}_{m_{2}}=\frac{m_{2}pN_{1}}{N_{123}}a^{1}$; see Eqs. (34) and (40). When the currents $J^{1}_{m_{1}}$ and $J^{2}_{m_{2}}$ are combined, the condition for the resulting current to be conserved is relaxed to $d\phi_{m_{1},m_{2}}^{1,2}=\frac{m_{1}pN_{2}}{N_{123}}a^{2}-\frac{m_{2}pN_{1}}{N_{123}}a^{1}$.
The corresponding conserved quantity is $Q_{m_{1},m_{2}}=\int_{\sigma}*J_{m_{1},m_{2}}$, and the associated unitary operator is
[…]
$\left.\left.-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta_{m_{1}}^{3}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta_{m_{2}}^{3}\right)\right].$ (62)
Physically, $\mathcal{U}_{m_{1}m_{2}}$ corresponds to transporting a loop excitation along a 22d closed submanifold $\sigma$. The loop carries both $m_{1}$ units of $\mathbb{Z}_{N_{1}}$ gauge flux and $m_{2}$ units of $\mathbb{Z}_{N_{2}}$ gauge flux. This operation probes the numbers of $\mathbb{Z}_{N_{1}}$ and $\mathbb{Z}_{N_{2}}$ gauge charges, $n_{e_{1}}$ and $n_{e_{2}}$, resulting in the phase factor $\theta_{m_{1},m_{2}}=e^{{\rm i}\frac{2\pi m_{1}}{N_{1}}\times n_{e_{1}}+\frac{2\pi m_{2}}{N_{2}}\times n_{e_{2}}}$. The symmetry associated with this conserved current implies that $\theta_{m_{1},m_{2}}$ remains invariant under smooth deformations of $\sigma$. However, it is important to note that the current $J_{m_{1},m_{2}}$ is conserved only under specific conditions. Consequently, projectors must be introduced to enforce these conditions, and $\mathcal{U}_{m_{1}m_{2}}$ defines a symmetry operation only within the projected subspace. The resulting symmetry operator is
$L_{m_{1}m_{2}}\left(\sigma\right)$ $=\mathcal{U}_{m_{1}m_{2}}\delta\left(\int_{\gamma\subset\sigma}\frac{m_{1}pN_{2}}{N_{123}}a^{2}-\frac{m_{2}pN_{1}}{N_{123}}a^{1}\right)$
$\quad\times\delta\left(\frac{m_{1}pN_{3}}{N_{123}}\int_{\sigma}b^{3}\right)\delta\left(\frac{m_{2}pN_{3}}{N_{123}}\int_{\sigma}b^{3}\right),$ (63)
where the delta functions implement the projectors, $\delta\left(x\right)=1$ if $x=0\mod{2\pi}$ and $\delta\left(x\right)=0$ otherwise.
Another example of combining type-II currents involves $d\left(*J_{m}^{1}\right)=0$ and $d\left(*J_{m}^{3}\right)=0$. Since $J_{m}^{3}$ is a $2$-form, it must be promoted to a $3$-form as $\delta^{\perp}\left(\gamma_{p}\right)*J_{m}^{3}$ (with $\wedge$ omitted by convention). In this way, $\delta^{\perp}\left(\gamma_{p}\right)*J_{m}^{3}$ becomes a $3$-form supported only on the submanifold $\gamma_{p}$. We may then construct a new continuity equation starting from
$0=$ $m_{1}\cdot d\left(*J_{m}^{1}\right)+m_{3}\cdot d\left(\delta^{\perp}\left(\gamma_{p}\right)*J_{m}^{3}\right)$
$=$ $d\left(m_{1}b^{1}\right)+\frac{m_{1}pN_{2}N_{3}}{2\pi N_{123}}a^{2}b^{3}$
$-\delta^{\perp}\left(\gamma_{p}\right)d\left(m_{3}a^{3}\right)-\delta^{\perp}\left(\gamma_{p}\right)\frac{m_{3}pN_{1}N_{2}}{2\pi N_{123}}a^{1}a^{2}.$ (64)
We anticipate the existence of a new conserved current $J_{m_{1},m_{3}}$ satisfying $d\left(*J_{m_{1},m_{3}}\right)=m_{1}\cdot d\left(*J_{m}^{1}\right)+m_{3}\cdot d\left(\delta^{\perp}\left(\gamma_{p}\right)*J_{m}^{3}\right)=0$. Introducing auxiliary fields defined by $d\beta_{m_{1},m_{3}}^{1,3}=\frac{m_{1}pN_{3}}{N_{123}}b^{3}-\frac{m_{3}pN_{1}}{N_{123}}\delta^{\perp}\left(\gamma_{p}\right)a^{1}$, $d\phi_{m_{1}}^{2}=\frac{m_{1}pN_{2}}{N_{123}}a^{2}$, and $d\phi_{m_{3}}^{2}=\frac{m_{3}pN_{2}}{N_{123}}a^{2}$, we obtain
$0=$ $d\left(m_{1}b^{1}\right)+d\left(m_{3}\delta^{\perp}\left(\gamma_{p}\right)a^{3}\right)+\frac{1}{2}d\beta_{m_{1},m_{3}}^{1,3}\frac{N_{2}}{2\pi}a^{2}$
$+\frac{1}{2}d\phi_{m_{1}}^{2}\frac{N_{3}}{2\pi}b^{3}+\frac{1}{2}\delta^{\perp}\left(\gamma_{p}\right)d\phi_{m_{3}}^{2}\frac{N_{1}}{2\pi}a^{1}.$ (65)
We may therefore identify the conserved current as
$*J_{m_{1},m_{3}}=$ $m_{1}b^{1}+m_{3}\delta^{\perp}\left(\gamma_{p}\right)a^{3}+\frac{1}{2}\beta_{m_{1,}m_{3}}^{1,3}\frac{N_{2}}{2\pi}a^{2}$
$+\frac{1}{2}\phi_{m_{1}}^{2}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\delta^{\perp}\left(\gamma_{p}\right)\phi_{m_{3}}^{2}\frac{N_{1}}{2\pi}a^{1},$ (66)
provided that $d\beta_{m_{1},m_{3}}^{1,3}=\frac{m_{1}pN_{3}}{N_{123}}b^{3}-\frac{m_{3}pN_{1}}{N_{123}}\delta^{\perp}\left(\gamma_{p}\right)a^{1}$, $d\phi_{m_{1}}^{2}=\frac{m_{1}pN_{2}}{N_{123}}a^{2}$, and $d\phi_{m_{3}}^{2}=\frac{m_{3}pN_{2}}{N_{123}}a^{2}$. The associated conserved quantity is $Q_{m_{1},m_{3}}=\int_{\sigma}*J_{m_{1},m_{3}}$, and the corresponding unitary operator reads
$\mathcal{U}_{m_{1}m_{3}}=$ $\exp\left[{\rm i}\int_{\sigma}\left(m_{1}b^{1}+\frac{1}{2}\beta_{m_{1,}m_{3}}^{1,3}\frac{N_{2}}{2\pi}a^{2}+\frac{1}{2}\phi_{m_{1}}^{2}\frac{N_{3}}{2\pi}b^{3}\right)\right$.
$\left.+{\rm i}\int_{\gamma_{p}}\left(m_{3}a^{3}-\frac{1}{2}\phi_{m_{3}}^{2}\frac{N_{1}}{2\pi}a^{1}\right)\right],$ (67)
where we have used $\int_{\sigma}\delta^{\perp}\left(\gamma_{p}\right)(\cdots)=\int_{\gamma_{p}}(\cdots)$. This operator can be interpreted as the Wilson operator of a loop excitation carrying $\mathbb{Z}_{N_{1}}$ gauge flux, decorated by a particle excitation carrying $\mathbb{Z}_{N_{3}}$ gauge charge, up to an overall normalization factor. The role of $\delta^{\perp}\left(\gamma_{p}\right)$ is now transparent: since the particle is attached to the loop, its worldline $\gamma_{p}$ lies on the worldsheet $\sigma$ of the loop. Operationally, $\mathcal{U}_{m_{1}m_{3}}$ probes the numbers of $\mathbb{Z}_{N_{1}}$ gauge charges and $\mathbb{Z}_{N_{3}}$ gauge fluxes via a loop decorated by a particle. The appearance of $\delta^{\perp}\left(\gamma_{p}\right)$ ensures that the current $J_{m}^{3}$ is bound to $J_{m}^{1}$. As before, the current $J_{m_{1},m_{3}}$ is conserved only when specific conditions are satisfied. These conditions can be enforced by introducing projectors, leading to the symmetry operator
$L_{m_{1}m_{3}}\left(\sigma,\gamma_{p}\right)$ $=\mathcal{U}_{m_{1}m_{3}}\delta\left(\frac{m_{1}pN_{3}}{N_{123}}\int_{\sigma}b^{3}-\frac{m_{3}pN_{1}}{N_{123}}\int_{\gamma_{p}}a^{1}\right)$
$\quad\times\delta\left(\int_{\gamma\subset\sigma}\frac{m_{1}pN_{2}}{N_{123}}a^{2}\right)\delta\left(\int_{\gamma\subset\sigma}\frac{m_{3}pN_{2}}{N_{123}}a^{2}\right).$ (68)
Since $\gamma_{p}$ may be chosen as an arbitrary closed curve on $\sigma$, the integrals in the last two delta functions can equivalently be written as $\delta\left(\int_{\gamma_{p}}\cdots\right)\delta\left(\int_{\gamma_{p}}\cdots\right)$. Accordingly, the symmetry operator may also be expressed as
$L_{m_{1}m_{3}}\left(\sigma,\gamma_{p}\right)$ $=\mathcal{U}_{m_{1}m_{3}}\delta\left(\frac{m_{1}pN_{3}}{N_{123}}\int_{\sigma}b^{3}-\frac{m_{3}pN_{1}}{N_{123}}\int_{\gamma_{p}}a^{1}\right)$
$\quad\times\delta\left(\int_{\gamma_{p}}\frac{m_{1}pN_{2}}{N_{123}}a^{2}\right)\delta\left(\int_{\gamma_{p}}\frac{m_{3}pN_{2}}{N_{123}}a^{2}\right).$ (69)
When either $m_{1}$ or $m_{3}$ vanishes, the symmetry operator $L_{m_{1}m_{3}}\left(\sigma,\gamma_{p}\right)$ reduces to $L_{m_{3}}\left(\gamma_{p}\right)$ or $L_{m_{1}}\left(\sigma\right)$, respectively.
The above two examples illustrate how two conserved currents of type-II can be combined to produce a new one. More generally, we can combine the three distinct type-II currents to obtain a new conserved current $J_{m_{1},m_{2},m_{3}}$ such that
$d\left(*J_{m_{1},m_{2},m_{3}}\right)=0$ $=m_{1}\cdot d\left(*J_{m}^{1}\right)+m_{2}\cdot d\left(*J_{m}^{2}\right)$
$+m_{3}\cdot d\left(\delta^{\perp}\left(\gamma_{p}\right)*J_{m}^{3}\right).$ (70)
The right-hand side then yields
$0=$ $d\left(m_{1}b^{1}\right)+\frac{1}{2}\frac{m_{1}pN_{2}}{N_{123}}a^{2}\frac{N_{3}}{2\pi}b^{3}+\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\frac{m_{1}pN_{3}}{N_{123}}b^{3}$
$+$ $d\left(m_{2}b^{2}\right)-\frac{1}{2}\frac{m_{2}pN_{1}}{N_{123}}a^{1}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\frac{m_{2}pN_{3}}{N_{123}}b^{3}$
$+$ $d\left(\delta^{\perp}\left(\gamma_{p}\right)m_{3}a^{3}\right)-\delta^{\perp}\left(\gamma_{p}\right)\frac{1}{2}\frac{m_{3}pN_{1}}{N_{123}}a^{1}\frac{N_{2}}{2\pi}a^{2}-\delta^{\perp}\left(\gamma_{p}\right)\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\frac{m_{3}pN_{2}}{N_{123}}a^{2}.$ (71)
In order to extract a continuity equation, namely a vanishing total derivative, we introduce auxiliary fields satisfying $d\phi_{m_{1},m_{2}}^{2,1}=\frac{m_{1}pN_{2}}{N_{123}}a^{2}-\frac{m_{2}pN_{1}}{N_{123}}a^{1}$, $d\beta_{m_{1},m_{3}}^{3,1}=\frac{m_{1}pN_{3}}{N_{123}}b^{3}-\delta^{\perp}\left(\gamma_{p}\right)\frac{m_{3}pN_{1}}{N_{123}}a^{1}$, and $d\beta_{m_{2},m_{3}}^{3,2}=\frac{m_{2}pN_{3}}{N_{123}}b^{3}-\delta^{\perp}\left(\gamma_{p}\right)\frac{m_{3}pN_{2}}{N_{123}}a^{2}$. With these definitions, the above equation becomes
$0=$ $d\left(m_{1}b^{1}+m_{2}b^{2}+\delta^{\perp}\left(\gamma_{p}\right)m_{3}a^{3}\right)$
$+\frac{1}{2}d\phi_{m_{1},m_{2}}^{2,1}\frac{N_{3}}{2\pi}b^{3}+\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}d\beta_{m_{3},m_{1}}^{3,1}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}d\beta_{m_{2},m_{3}}^{3,2}$
$=$ $d\left(m_{1}b^{1}+m_{2}b^{2}+\delta^{\perp}\left(\gamma_{p}\right)m_{3}a^{3}\right)$
$+d\left(\frac{1}{2}\phi_{m_{1},m_{2}}^{2,1}\frac{N_{3}}{2\pi}b^{3}\right)-d\left(\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta_{m_{3},m_{1}}^{3,1}\right)+d\left(\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta_{m_{2},m_{3}}^{3,2}\right),$ (72)
from which we may identify
$*J_{m_{1},m_{2},m_{3}}=$ $m_{1}b^{1}+m_{2}b^{2}+\delta^{\perp}\left(\gamma_{p}\right)m_{3}a^{3}$
$+\frac{1}{2}\phi_{m_{1},m_{2}}^{2,1}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta_{m_{3},m_{1}}^{3,1}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta_{m_{2},m_{3}}^{3,2}.$ (73)
The unitary operator associated with the conserved quantity $Q_{m_{1},m_{2},m_{3}}=\int_{\sigma}*J_{m_{1},m_{2},m_{3}}$ is then given by
$\mathcal{U}_{m_{1}m_{2}m_{3}}\left(\sigma,\gamma_{p}\right)$
$=$ $\exp\left[{\rm i}\int_{\sigma}m_{1}b^{1}+{\rm i}\int_{\sigma}m_{2}b^{2}+{\rm i}\int_{\gamma_{p}}m_{3}a^{3}\right$.
$\left.+{\rm i}\int_{\sigma}\frac{1}{2}\left(\phi_{m_{1},m_{2}}^{2,1}\frac{N_{3}}{2\pi}b^{3}-\frac{N_{2}}{2\pi}a^{2}\beta_{m_{3},m_{1}}^{3,1}+\frac{N_{1}}{2\pi}a^{1}\beta_{m_{2},m_{3}}^{3,2}\right)\right].$ (74)
By further introducing projectors that enforce the conditions under which $d\left(*J_{m_{1},m_{2},m_{3}}\right)=0$, we obtain the symmetry operator
$L_{m_{1}m_{2}m_{3}}\left(\sigma,\gamma_{p}\right)=\mathcal{U}_{m_{1}m_{2}m_{3}}\mathbb{P}_{m_{1}m_{2}m_{3}},$ (75)
where the projectors are implemented by delta functions,
$\mathbb{P}_{m_{1}m_{2}m_{3}}$ $=\delta\left(\frac{m_{1}pN_{2}}{N_{123}}\int_{\gamma_{p}}a^{2}-\frac{m_{2}pN_{1}}{N_{123}}\int_{\gamma_{p}}a^{1}\right)$
$\times\delta\left(\frac{m_{1}pN_{3}}{N_{123}}\int_{\sigma}b^{3}-\frac{m_{3}pN_{1}}{N_{123}}\int_{\gamma_{p}}a^{1}\right)$
$\times\delta\left(\frac{m_{2}pN_{3}}{N_{123}}\int_{\sigma}b^{3}-\frac{m_{3}pN_{2}}{N_{123}}\int_{\gamma_{p}}a^{2}\right).$ (76)
Here, the first delta function was originally written as $\delta\left(\frac{m_{1}pN_{2}}{N_{123}}\int_{\gamma\subset\sigma}a^{2}-\frac{m_{2}pN_{1}}{N_{123}}\int_{\gamma\subset\sigma}a^{1}\right)$. Using the fact that $\gamma_{p}$ can be chosen as an arbitrary closed curve on $\sigma$, we replace $\int_{\gamma\subset\sigma}(\cdots)$ by $\int_{\gamma_{p}}(\cdots)$.
IV.2 A simple example of fusion rules of symmetry operators: $G=\left(\mathbb{Z}_{2}\right)^{3}$
From the previous discussion, we already know the general forms of the invertible and non-invertible symmetry operators, namely, $U_{e_{1}e_{2}e_{3}}\left(\sigma,\gamma_{p}\right)$ and $L_{m_{1}m_{2}m_{3}}\left(\sigma,\gamma_{p}\right)$ given by Eq. (57) and Eq. (75), respectively. Their fusion rules form an essential part of the algebraic data of the generalized symmetry in the theory. Here, “fusion” means inserting two symmetry operators at the same location (or, more precisely, on nearby supporting manifolds). The net effect, inside correlation functions, is equivalent to inserting another symmetry operator or a linear combination of symmetry operators, as dictated by the fusion rule.
We now present the simplest nontrivial example by choosing the gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$ as $G=\left(\mathbb{Z}_{2}\right)^{3}$. Consider first the fusion of two $U_{e_{1}=1}\left(\gamma_{p}\right)=\exp\left({\rm i}\int_{\gamma_{p}}a^{1}\right)$. Applying the operator twice gives
$\left\langle U_{e_{1}=1}\left(\gamma_{p}\right)\times U_{e_{1}=1}\left(\gamma_{p}\right)\mathcal{O}\right\rangle$ $=\left\langle\exp\left({\rm i}2\int_{\gamma_{p}}a^{1}\right)\mathcal{O}\right\rangle$
$=\left\langle U_{e_{1}=2}\left(\gamma_{p}\right)\mathcal{O}\right\rangle,$ (77)
where $\mathcal{O}$ is an arbitrary operator and $\left\langle\cdots\right\rangle$ denotes a correlation function. The fusion rule $U_{e_{1}=1}\left(\gamma_{p}\right)\times U_{e_{1}=1}\left(\gamma_{p}\right)=U_{e_{1}=2}\left(\gamma_{p}\right)$ should be understood as an equality inside correlation functions. To evaluate $\left\langle\exp\left({\rm i}2\int_{\gamma_{p}}a^{1}\right)\mathcal{O}\right\rangle$, we sum over all configurations of the gauge fields $\left[a^{i},b^{i}\right]$. When $\mathbb{Z}_{N_{1}}=\mathbb{Z}_{2}$, there are two topological sectors labeled by $w=\int_{\gamma_{p}}a^{1}\in\left\{0,\pi\right\}$, which cannot be related by gauge transformations. The correlation function is
$\left\langle e^{{\rm i}2\int_{\gamma_{p}}a^{1}}\mathcal{O}\right\rangle$
$=$ $\frac{1}{\mathcal{Z}}\sum_{w\in\left\{0,\pi\right\}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\gamma_{p}}a^{1}=w\right]e^{{\rm i}S_{aab}}e^{{\rm i}2\int_{\gamma_{p}}a^{1}}\mathcal{O},$ (78)
where $\mathcal{D}\left[a^{i},b^{i};\int_{\gamma_{p}}a^{1}=w\right]$ denotes configurations with $\int_{\gamma_{p}}a^{1}=w$. In either sector, $e^{{\rm i}2\int_{\gamma_{p}}a^{1}}=1$, and hence $e^{{\rm i}2\int_{\gamma_{p}}a^{1}}\mathcal{O}=\mathcal{O}$. Therefore, $\left\langle\exp\left({\rm i}2\int_{\gamma_{p}}a^{1}\right)\mathcal{O}\right\rangle=\left\langle\mathcal{O}\right\rangle$ for $N_{1}=2$. Equivalently, the fusion rule reduces to
$U_{e_{1}=1}\left(\gamma_{p}\right)\times U_{e_{1}=1}\left(\gamma_{p}\right)=1,$ (79)
where $1$ denotes the identity operator. We thus see that the $2$-form symmetry operator $U_{e_{1}=1}\left(\gamma_{p}\right)$ is invertible, with its inverse equal to itself when $\mathbb{Z}_{N_{1}}=\mathbb{Z}_{2}$. In this sense, the symmetry operators $U_{e_{1}}\left(\gamma_{p}\right)$ form a $\mathbb{Z}_{2}^{\left(2\right)}$ group, where the superscript denotes a $2$-form symmetry. Similar conclusions extend to an arbitrary cyclic group $\mathbb{Z}_{N_{1}}$, and likewise to $U_{e_{2}}\left(\gamma_{p}\right)$ and $U_{e_{3}}\left(\sigma\right)$. The fusion of $U_{e_{1}}\left(\gamma_{p}\right)$ with $U_{e_{2}}\left(\gamma_{p}\right)$ or $U_{e_{3}}\left(\sigma\right)$ is also straightforward:
$U_{e_{1}}\left(\gamma_{p}\right)\times U_{e_{2}}\left(\gamma_{p}\right)$ $=U_{e_{1},e_{2}}\left(\gamma_{p}\right)$
$=\exp\left({\rm i}\int_{\gamma_{p}}e_{1}a^{1}+{\rm i}\int_{\gamma_{p}}e_{2}a^{2}\right),$ (80)
$U_{e_{1}}\left(\gamma_{p}\right)\times U_{e_{3}}\left(\sigma\right)$ $=U_{e_{1},e_{3}}\left(\sigma,\gamma_{p}\right)$
$=\exp\left({\rm i}\int_{\gamma_{p}}e_{1}a^{1}+{\rm i}\int_{\sigma}e_{3}b^{3}\right).$ (81)
More generally, the fusion rules of $U_{e_{1}e_{2}e_{3}}\left(\sigma,\gamma_{p}\right)$ are group-like: each $U_{e_{1}e_{2}e_{3}}\left(\sigma,\gamma_{p}\right)$ can be labeled by an element of $\mathbb{Z}_{N_{1}}^{\left(2\right)}\times\mathbb{Z}_{N_{2}}^{\left(2\right)}\times\mathbb{Z}_{N_{3}}^{\left(1\right)}$, and fusion corresponds to group multiplication. This is consistent with the fact that $U_{e_{1}e_{2}e_{3}}\left(\sigma,\gamma_{p}\right)$ are also Wilson operators of Abelian topological excitations, whose fusion rules are likewise group-like. Moreover, for each $U_{e_{1}e_{2}e_{3}}\left(\sigma,\gamma_{p}\right)$, there exists an inverse $U_{e_{1}^{\prime}e_{2}^{\prime}e_{3}^{\prime}}\left(\sigma,\gamma_{p}\right)$ with $e_{i}+e_{i}^{\prime}=N_{i}$ for $i=1,2,3$ such that
$U_{e_{1}e_{2}e_{3}}\left(\sigma,\gamma_{p}\right)\times U_{e_{1}^{\prime}e_{2}^{\prime}e_{3}^{\prime}}\left(\sigma,\gamma_{p}\right)=1.$ (82)
Even in the simplest case of $G=\left(\mathbb{Z}_{2}\right)^{3}$, there exist non-invertible symmetry operators whose fusion is not group-like. Consider the fusion of two $L_{m_{1}=1}\left(\sigma\right)=\mathcal{U}_{m_{1}=1}\delta\left(\int_{\sigma}b^{3}\right)\delta\left(\int_{\gamma\subset\sigma}a^{2}\right)$. Inserting two copies into a correlation function yields
$\left\langle L_{m_{1}=1}\left(\sigma\right)\times L_{m_{1}=1}\left(\sigma\right)\mathcal{O}\right\rangle$
$=$ $\left\langle\mathcal{U}_{m_{1}=1}\mathcal{U}_{m_{1}=1}\delta\left(\int_{\sigma}b^{3}\right)\delta\left(\int_{\gamma\subset\sigma}a^{2}\right)\mathcal{O}\right\rangle.$ (83)
For $\mathbb{Z}_{N_{1}}=\mathbb{Z}_{2}$, $\mathcal{U}_{m_{1}=1}\mathcal{U}_{m_{1}=1}=1$ since $\mathcal{U}_{m_{1}=1}$ is unitary. Thus, the fusion of two $L_{m_{1}=1}\left(\sigma\right)$ leaves only the two delta functions, namely, the projectors. A delta function inside a correlation function selects a specific topological sector in the path-integral summation. For example, in $\left\langle\delta\left(\int_{\sigma}b^{3}\right)\mathcal{O}\right\rangle$, the values $\int_{\sigma}b^{3}\in\left\{0,\pi\right\}$ label two sectors. We can expand the delta function as $\delta\left(\int_{\sigma}b^{3}\right)=\frac{1}{2}\left(1+e^{{\rm i}\int_{\sigma}b^{3}}\right)$ because
$\left\langle\delta\left(\int_{\sigma}b^{3}\right)\mathcal{O}\right\rangle$
$=$ $\frac{1}{\mathcal{Z}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\sigma}b^{3}=0\right]e^{{\rm i}S_{aab}}\mathcal{O}$
$=$ $\frac{1}{\mathcal{Z}}\sum_{w\in\left\{0,\pi\right\}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\sigma}b^{3}=w\right]e^{{\rm i}S_{aab}}\frac{1}{2}\left(1+e^{{\rm i}\int_{\sigma}b^{3}}\right)\mathcal{O}.$ (84)
In this way, we obtain $\left\langle\delta\left(\int_{\sigma}b^{3}\right)\mathcal{O}\right\rangle=\left\langle\frac{1}{2}\left(1+e^{{\rm i}\int_{\sigma}b^{3}}\right)\mathcal{O}\right\rangle$. Similarly,
$\left\langle\delta\left(\int_{\sigma}b^{3}\right)\delta\left(\int_{\gamma\subset\sigma}a^{2}\right)\mathcal{O}\right\rangle$
$=$ $\left\langle\frac{1}{4}\left(1+e^{{\rm i}\int_{\gamma\subset\sigma}a^{2}}+e^{{\rm i}\int_{\sigma}b^{3}}+e^{{\rm i}\int_{\sigma}b^{3}+{\rm i}\int_{\gamma\subset\sigma}a^{2}}\right)\mathcal{O}\right\rangle,$ (85)
where the exponents are, in fact, invertible symmetry operators $U_{e_{1}e_{2}e_{3}}\left(\sigma,\gamma\subset\sigma\right)$. We therefore obtain the fusion rule
$L_{m_{1}=1}\times L_{m_{1}=1}=\frac{1}{4}\left(1+U_{010}+U_{001}+U_{011}\right),$ (86)
where we neglect the supporting manifolds of the symmetry operators for simplicity. This fusion rule shows that applying $L_{m_{1}=1}\left(\sigma\right)$ twice is equivalent, inside correlation functions, to inserting a linear combination of other symmetry operators, providing a concrete example of non-group-like fusion.
Next, we consider the fusion of $L_{m_{1}=1}\left(\sigma\right)$ and $L_{m_{2}=1}\left(\sigma\right)$. Inserting them in a correlation function, we get
$\left\langle L_{m_{1}=1}\left(\sigma\right)\times L_{m_{2}=1}\left(\sigma\right)\mathcal{O}\right\rangle$
$=$ $\left\langle\mathcal{U}_{m_{1}=1}\delta\left(\int_{\sigma}b^{3}\right)\delta\left(\int_{\gamma\subset\sigma}a^{2}\right)\right$.
$\left.\times\mathcal{U}_{m_{2}=1}\delta\left(\int_{\sigma}b^{3}\right)\delta\left(\int_{\gamma\subset\sigma}a^{1}\right)\mathcal{O}\right\rangle$ (87)
Since $\mathcal{U}$ operators are unitary operators, $\mathcal{U}_{m_{1}=1}\mathcal{U}_{m_{2}=1}=\mathcal{U}_{m_{1}=1,m_{2}=1}$, which is the unitary operator appearing in the symmetry operator $L_{m_{1}=1,m_{2}=1}\left(\sigma\right)$. This suggests that the fusion outcome is related to $L_{m_{1}=1,m_{2}=1}\left(\sigma\right)=\mathcal{U}_{m_{1}=1,m_{2}=1}\delta\left(\int_{\gamma\subset\sigma}a^{2}-a^{1}\right)\delta\left(\int_{\sigma}b^{3}\right)$. In the case of $G=\left(\mathbb{Z}_{2}\right)^{3}$, $\left\langle\delta\left(\int_{\gamma}a^{2}-a^{1}\right)\mathcal{O}\right\rangle=\left\langle\frac{1}{2}\left(1+e^{{\rm i}\int_{\gamma}\left(a^{2}+a^{1}\right)}\right)\mathcal{O}\right\rangle$, where we have inserted $e^{{\rm i}2\int_{\gamma}a^{1}}=1$ in the correlation function. We also find that $\left\langle\delta\left(\int_{\gamma}a^{2}\right)\delta\left(\int_{\gamma}a^{1}\right)\mathcal{O}\right\rangle=\left\langle\frac{1}{2}\left(1+e^{{\rm i}\int_{\gamma}a^{1}}\right)\delta\left(\int_{\gamma}a^{2}-a^{1}\right)\mathcal{O}\right\rangle$. Replacing the delta functions in $\left\langle L_{m_{1}=1}\left(\sigma\right)\times L_{m_{2}=1}\left(\sigma\right)\mathcal{O}\right\rangle$, we have
$\left\langle L_{m_{1}=1}\left(\sigma\right)\times L_{m_{2}=1}\left(\sigma\right)\mathcal{O}\right\rangle$
$=$ $\left\langle\frac{1}{2}\left(1+e^{{\rm i}\int_{\gamma\subset\sigma}a^{1}}\right)L_{m_{1}=1,m_{2}=1}\left(\sigma\right)\mathcal{O}\right\rangle,$ (88)
which yields the fusion rule
$L_{m_{1}=1}\times L_{m_{2}=1}=\frac{1}{2}L_{m_{1}=1,m_{2}=1}+\frac{1}{2}U_{100}L_{m_{1}=1,m_{2}=1}.$ (89)
Here, $U_{m_{1}=1}L_{m_{1}=1,m_{2}=1}$ denotes the symmetry operator obtained by fusing $U_{m_{1}=1}$ and $L_{m_{1}=1,m_{2}=1}$, whose expression is $e^{{\rm i}\int_{\gamma\subset\sigma}a^{1}+\ln\mathcal{U}_{m_{1}=1,m_{2}=1}}\delta\left(\int_{\gamma\subset\sigma}a^{2}-a^{1}\right)\delta\left(\int_{\sigma}b^{3}\right)$.
According to the correspondence between symmetry operators and topological excitations, the fusion rules of symmetry operators can be mapped to those of topological excitations [76]. One caveat is that in the fusion rules of topological excitations, the coefficients of different fusion channels are integers, which encode the quantum dimensions. Here, unlike Wilson operators of topological excitations, the symmetry operators do not come with such coefficients. This is because a symmetry operator acts on the Hilbert space, and an overall coefficient can be absorbed into a state vector. It would be interesting to understand how to encode the information about the quantum dimension directly at the level of symmetry operators.
IV.3 An overview of fusion rules of symmetry operators
In this section, we give an overview of the fusion rules of symmetry operators for a general gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$. Several concrete examples will be discussed in Sec. IV.4. As established in the previous sections, symmetry operators without projectors are invertible, in the sense that one can always find another operator whose insertion into a correlation function is equivalent to the identity. By contrast, symmetry operators accompanied by projectors (i.e., delta functions) are non-invertible. These projectors restrict the path integral to specific topological sectors (or subspaces), within which a conserved current and an associated symmetry can be defined. Since a projector does not admit an inverse, the corresponding symmetry operator is intrinsically non-invertible.
A general invertible symmetry operator is denoted by $U_{e_{1}e_{2}e_{3}}\left(\sigma,\gamma_{p}\right)$, see Eq. (57), or simply $U_{e_{1}e_{2}e_{3}}$ for brevity, where the integers $e_{i}$ take values from $1$ to $N_{i}-1$. As discussed earlier, $U_{e_{1}e_{2}e_{3}}$ can also be interpreted as the Wilson operator of an Abelian topological excitation. In this interpretation, the integers $e_{i}$ specify the numbers of $\mathbb{Z}_{N_{i}}$ gauge charges and fluxes carried by the excitation. Equivalently, $U_{e_{1}e_{2}e_{3}}$ can be viewed as the fusion product of $U_{e_{1}}\left(\gamma_{p}\right)$, $U_{e_{2}}\left(\gamma_{p}\right)$, and $U_{e_{3}}\left(\sigma\right)$. Here, $U_{e_{1}}\left(\gamma_{p}\right)$ and $U_{e_{2}}\left(\gamma_{p}\right)$ are $2$-form symmetry operators, while $U_{e_{3}}\left(\sigma\right)$ is a $1$-form symmetry operator. The fusion rules of invertible symmetry operators (with supporting manifolds suppressed) are group-like,
$U_{i_{1}j_{1}k_{1}}\times U_{i_{2}j_{2}k_{2}}=U_{\left(i_{1}+i_{2}\right)\left(j_{1}+j_{2}\right)\left(k_{1}+k_{2}\right)},$ (90)
where $i_{1}+i_{2}$ is understood modulo $N_{1}$ because $U_{N_{1}00}=1$; analogously, $j_{1}+j_{2}$ and $k_{1}+k_{2}$ are defined modulo $N_{2}$ and $N_{3}$, respectively.
A general non-invertible symmetry operator is given by $L_{m_{1}m_{2}m_{3}}\left(\sigma,\gamma_{p}\right)$, see Eq. (75), or simply $L_{m_{1}m_{2}m_{3}}$, where the integers $m_{i}$ range from 0 to $N_{i}-1$. Such an operator corresponds, up to an overall normalization factor, to the Wilson operator of a non-Abelian topological excitation. In this sense, the integers $m_{i}$ characterize the $\mathbb{Z}_{N_{i}}$ gauge fluxes and charges carried by the excitation. Each $L_{m_{1}m_{2}m_{3}}$ consists of a unitary operator $\mathcal{U}_{m_{1}m_{2}m_{3}}$ together with a projector $\mathbb{P}_{m_{1}m_{2}m_{3}}$ constructed from delta functions.
The fusion rules of the non-invertible symmetry operators $L_{m_{1}m_{2}m_{3}}$, for example $L_{ijk}\times L_{rst}$, can be obtained systematically as follows. One first treats the fusion of the unitary parts and the projectors separately. The fusion of the unitary operators is straightforward,
$\left\langle\mathcal{U}_{ijk}\times\mathcal{U}_{rst}\mathcal{O}\right\rangle=\left\langle\mathcal{U}_{\left(i+r\right)\left(j+s\right)\left(k+t\right)}\mathcal{O}\right\rangle,$ (91)
since they appear as multiplicative factors in the functional integral. This result reproduces precisely the unitary component of the symmetry operator $L_{\left(i+r\right)\left(j+s\right)\left(k+t\right)}$. By contrast, the fusion of the projectors yields the product $\mathbb{P}_{ijk}\mathbb{P}_{rst}$, which can be rewritten, inside correlation functions, as a linear combination of invertible symmetry operators, for example as in Eq. (84). Consequently, one can identify a set of invertible operators $U_{abc}$, possibly accompanied by coefficients $f_{abc}$, such that
$\left\langle\mathbb{P}_{ijk}\times\mathbb{P}_{rst}\mathcal{O}\right\rangle=\left\langle\sum_{a,b,c}f_{abc}U_{abc}\mathbb{P}_{\left(i+r\right)\left(j+s\right)\left(k+t\right)}\mathcal{O}\right\rangle.$ (92)
Combining the unitary and projector contributions, the general fusion rule takes the form
$L_{ijk}\times L_{rst}=\sum_{a,b,c}f_{abc}U_{abc}L_{\left(i+r\right)\left(j+s\right)\left(k+t\right)}.$ (93)
Therefore, the essential step in determining the fusion rules is the systematic manipulation of the projectors to extract the corresponding invertible symmetry operators $U_{abc}$ and their coefficients $f_{abc}$. For a general symmetry operator $L_{m_{1}m_{2}m_{3}}$, the explicit form of the associated projectors is given in Eq. (76).
IV.4 More examples of fusing (non-)invertible higher-form symmetry operators
In this section, we present further examples of fusion rules for different choices of gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$, illustrating in a concrete and systematic manner how fusion rules of symmetry operators can be computed in practice.
IV.4.1 $N_{1}=N_{2}=N_{3}=3$, $p=1$
We generalize the simplest case to $N_{1}=N_{2}=N_{3}=3$, for which $N_{123}=3$. The nontrivial values of $m_{i}$ and $p$ are $m_{i}=1,2$ and $p=1,2$. Here we focus first on the case $p=1$. In this setting, the delta functions appearing in the projectors reduce to $\delta\left(m_{1}\int_{\gamma_{p}}a^{2}-m_{2}\int_{\gamma_{p}}a^{1}\right)$, $\delta\left(m_{1}\int_{\sigma}b^{3}-m_{3}\int_{\gamma_{p}}a^{1}\right)$, and $\delta\left(m_{2}\int_{\sigma}b^{3}-m_{3}\int_{\gamma_{p}}a^{2}\right)$.
The symmetry operator $L_{100}\left(\sigma\right)$ takes the form $L_{100}\left(\sigma\right)=\mathcal{U}_{100}\delta\left(\int_{\sigma}b^{3}\right)\delta\left(\int_{\gamma_{p}}a^{2}\right)$. We first consider the fusion $L_{100}\left(\sigma\right)\times L_{100}\left(\sigma\right)$. The fusion of the unitary parts gives $\mathcal{U}_{100}\times\mathcal{U}_{100}=\mathcal{U}_{200}$, indicating that $L_{200}\left(\sigma\right)$ appears among the fusion channels. The fusion of the projectors is obtained by expanding the delta functions into sums of exponentials. Since the possible values of $\int_{\sigma}b^{3}$ belong to $\left\{0,\frac{2\pi}{N_{3}},\frac{4\pi}{N_{3}}\right\}$, we have $\delta\left(\int_{\sigma}b^{3}\right)=\frac{1}{3}\sum_{l_{3}=0}^{2}U_{00l_{3}}$ with $U_{00l_{3}}=e^{{\rm i}l_{3}\int_{\sigma}b^{3}}$. Therefore, $\delta\left(\int_{\sigma}b^{3}\right)\delta\left(\int_{\gamma_{p}}a^{2}\right)=\left(\frac{1}{3}\right)^{2}\sum_{l_{2},l_{3}=0}^{2}U_{0l_{2}l_{3}}$, where $U_{0l_{2}l_{3}}=e^{{\rm i}l_{2}\int_{\gamma_{p}}a^{2}+{\rm i}l_{3}\int_{\sigma}b^{3}}$ is an invertible symmetry operator. On the other hand, $L_{200}\left(\sigma\right)=\mathcal{U}_{200}\delta\left(2\int_{\sigma}b^{3}\right)\delta\left(2\int_{\gamma_{p}}a^{2}\right)$. Since $2$ is coprime with $N_{2}=N_{3}=3$, $L_{100}\left(\sigma\right)$ and $L_{200}\left(\sigma\right)$ share the same projectors. Matching the delta functions on both sides, we obtain
$L_{100}\left(\sigma\right)\times L_{100}\left(\sigma\right)=L_{200}\left(\sigma\right).$ (94)
Next, we consider the fusion $L_{100}\left(\sigma\right)\times L_{200}\left(\sigma\right)$. Fusing the unitary operators $\mathcal{U}_{100}$ and $\mathcal{U}_{200}$ yields the identity, while fusing two identical projectors simply reproduces the same projector. Consequently, the fusion rule is
$L_{100}\left(\sigma\right)\times L_{200}\left(\sigma\right)=\left(\frac{1}{3}\right)^{2}\sum_{l_{2},l_{3}=0}^{2}U_{0l_{2}l_{3}}\left(\sigma,\gamma_{p}\right).$ (95)
We now examine the fusion of $L_{100}\left(\sigma\right)$ and $L_{010}\left(\sigma\right)=\mathcal{U}_{010}\delta\left(\int_{\sigma}b^{3}\right)\delta\left(\int_{\gamma_{p}}a^{1}\right)$. The unitary parts fuse as $\mathcal{U}_{100}\times\mathcal{U}_{010}=\mathcal{U}_{110}$, which is precisely the unitary part of $L_{110}\left(\sigma\right)=\mathcal{U}_{110}\delta\left(\int_{\sigma}b^{3}\right)\delta\left(\int_{\gamma_{p}}a^{2}-\int_{\gamma_{p}}a^{1}\right)$. The fusion of the projectors gives $\delta\left(\int_{\sigma}b^{3}\right)\delta\left(\int_{\gamma_{p}}a^{2}\right)\delta\left(\int_{\gamma_{p}}a^{1}\right)$. To match this with the projector structure of $L_{110}\left(\sigma\right)$, we expand the delta functions as
$\delta\left(\int_{\gamma_{p}}a^{2}\right)\delta\left(\int_{\gamma_{p}}a^{1}\right)=\left(\frac{1}{3}\right)^{2}\sum_{l_{1},l_{2}=0}^{2}e^{{\rm i}\int_{\gamma_{p}}l_{1}a^{1}+l_{2}a^{2}},$ (96)
and
$\delta\left(\int_{\gamma_{p}}a^{2}-\int_{\gamma_{p}}a^{1}\right)=$ $\frac{1}{3}\sum_{c=0}^{2}e^{{\rm i}c\cdot\left(\int_{\gamma_{p}}a^{2}-a^{1}\right)}$
$=$ $\frac{1}{3}\left(e^{{\rm i}\int_{\gamma_{p}}a^{2}+2a^{1}}+e^{{\rm i}\int_{\gamma_{p}}2a^{2}+a^{1}}+1\right),$ (97)
where we have used $e^{{\rm i}3\int_{\gamma_{p}}a^{1}}=1$. From these expressions, one verifies that
$\delta\left(\int_{\gamma_{p}}a^{2}\right)\delta\left(\int_{\gamma_{p}}a^{1}\right)=\frac{1}{3}\sum_{l_{1}=0}^{2}U_{l_{1}00}\,\delta\left(\int_{\gamma_{p}}a^{2}-\int_{\gamma_{p}}a^{1}\right),$ (98)
so that the projectors are matched. The resulting fusion rule is
$L_{100}\left(\sigma\right)\times L_{010}\left(\sigma\right)=\frac{1}{3}\sum_{l_{1}=0}^{2}U_{l_{1}00}\left(\gamma_{p}\right)L_{110}\left(\sigma\right).$ (99)
As a more general example, consider the fusion $L_{120}\left(\sigma\right)\times L_{021}\left(\sigma\right)$. The corresponding symmetry operators are
$L_{120}\left(\sigma\right)$ $=\mathcal{U}_{120}\delta\left(\int_{\gamma_{p}}a^{2}-2\int_{\gamma_{p}}a^{1}\right)\delta\left(\int_{\sigma}b^{3}\right)\delta\left(2\int_{\sigma}b^{3}\right),$ (100)
and
$L_{021}\left(\sigma,\gamma_{p}\right)$ $=\mathcal{U}_{021}\delta\left(2\int_{\gamma_{p}}a^{1}\right)\delta\left(\int_{\gamma_{p}}a^{1}\right)\delta\left(2\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{2}\right).$ (101)
The fusion of the projectors yields $\mathbb{P}_{120}\times\mathbb{P}_{021}=\left(\frac{1}{3}\right)^{3}\sum_{l_{1},l_{2},l_{3}=0}^{2}U_{l_{1}l_{2}l_{3}}$, while the unitary parts fuse as $\mathcal{U}_{120}\times\mathcal{U}_{021}=\mathcal{U}_{111}$. Therefore, the fusion channels contain $L_{111}\left(\sigma,\gamma_{p}\right)=\mathcal{U}_{111}\mathbb{P}_{111}$, with projector
$\mathbb{P}_{111}=$ $\delta\left(\int_{\gamma_{p}}a^{2}-\int_{\gamma_{p}}a^{1}\right)\delta\left(\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{1}\right)\delta\left(\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{2}\right).$ (102)
Expanding $\mathbb{P}_{111}$ gives
$\mathbb{P}_{111}=\left(\frac{1}{3}\right)^{2}\sum_{i+j+k=0\mod 3}U_{ijk}.$ (103)
The constraints enforced by $\mathbb{P}_{111}$ imply $\int_{\gamma_{p}}a^{2}=\int_{\gamma_{p}}a^{1}=\int_{\sigma}b^{3}$. As a result, within correlation functions, $U_{ijk}$ and $U_{rst}$ are equivalent whenever $i+j+k=r+s+t=0\mod 3$. Consequently, all $U_{ijk}$ operators fall into three equivalence classes labeled by $i+j+k=0,1,2\mod 3$. One can verify that
$\sum_{l_{1},l_{2},l_{3}=0}^{2}U_{l_{1}l_{2}l_{3}}=\sum_{i+j+k\in 3\mathbb{Z}}U_{ijk}\times\left(1+U_{100}+U_{200}\right),$ (104)
where $\left(1+U_{100}+U_{200}\right)$ may equivalently be replaced by $\left(1+U_{010}+U_{020}\right)$ or $\left(1+U_{001}+U_{002}\right)$. We therefore arrive at the fusion rule
$L_{120}\times L_{021}=\frac{1}{3}\left(L_{111}+U_{100}L_{111}+U_{200}L_{111}\right).$ (105)
Here $U_{100}L_{111}$ and $U_{010}L_{111}$ should be regarded as equivalent symmetry operators, in the sense that their correlation functions with all other operators coincide. This equivalence follows from the fact that the projector in $L_{111}$ enforces $U_{100}$ and $U_{010}$ to take the same value.
In this subsection we have focused on the case $N_{1}=N_{2}=N_{3}=N=3$ and $p=1$. It is straightforward to extend these results to other values of $N$ and $p$. In Appendix B, we provide an explicit example for $N=6$ and $p=2$.
IV.4.2 $N_{1}=4$, $N_{2}=8$, $N_{3}=12$, $N_{123}=4$, $p=1$
In general, the gauge group of our effective field theory can be a product of three cyclic subgroups. In this example, we choose $N_{1}=4$, $N_{2}=8$, and $N_{3}=12$. We will consider two cases, $p=1$ and $p=2$, to investigate how the level of the twisted term influences the fusion channels. The comparison between these two cases is summarized in Table 2.
Table 2: Comparison of fusion rules between different levels of twisted term ($p$). The gauge group is $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}=\mathbb{Z}_{4}\times\mathbb{Z}_{8}\times\mathbb{Z}_{12}$. The greatest common divisor of $N_{1},N_{2},N_{3}$ is $N_{123}=4$. The nontrivial level of twisted term is $p\in\left\{1,2,3\right\}$. We can see that the fusion rules depend on the value of $p$.
Level of the twisted term Fusion rules
$p=1$ $L_{100}\times L_{100}=\frac{1}{4}\left(L_{200}+U_{020}L_{200}+U_{003}L_{200}+U_{023}L_{200}\right)$
$L_{100}\times L_{300}=\left(\frac{1}{4}\right)^{2}\sum_{l_{2},l_{3}=0}^{3}U_{0\left(2l_{2}\right)\left(2l_{3}\right)}$
$L_{123}\times L_{234}=\frac{1}{4}\sum_{l_{1}=0}^{3}U_{l_{1}00}L_{357}$
$p=2$ $L_{100}\times L_{100}=\frac{1}{4}\left(L_{200}+U_{040}L_{200}+U_{006}L_{200}+U_{046}L_{200}\right)$
$L_{100}\times L_{300}=\left(\frac{1}{2}\right)^{2}\sum_{l_{2},l_{3}=0}^{1}U_{0\left(4l_{2}\right)\left(6l_{3}\right)}$
$L_{123}\times L_{234}=\frac{1}{2}\left(L_{357}+U_{200}L_{357}\right)$
We consider several fusion rules in this setup: $L_{100}\times L_{100}$, $L_{100}\times L_{\left(N_{1}-1\right)00}$, and $L_{123}\times L_{234}$.
Fusing two $L_{100}=\mathcal{U}_{100}\delta\left(\frac{pN_{2}}{N_{123}}\int_{\gamma_{p}}a^{2}\right)\delta\left(\frac{pN_{3}}{N_{123}}\int_{\sigma}b^{3}\right)$ yields
$L_{100}\times L_{100}=\mathcal{U}_{200}\delta\left(\frac{pN_{2}}{N_{123}}\int_{\gamma_{p}}a^{2}\right)\delta\left(\frac{pN_{3}}{N_{123}}\int_{\sigma}b^{3}\right),$ (106)
where $\mathcal{U}_{200}$ is the unitary operator in $L_{200}=\mathcal{U}_{200}\delta\left(\frac{2pN_{2}}{N_{123}}\int_{\gamma_{p}}a^{2}\right)\delta\left(\frac{2pN_{3}}{N_{123}}\int_{\sigma}b^{3}\right)$. Plugging in the parameters, we have $\mathbb{P}_{100}\times\mathbb{P}_{100}=\delta\left(2\int_{\gamma_{p}}a^{2}\right)\delta\left(3\int_{\sigma}b^{3}\right)$. Since $\int_{\gamma_{p}}a^{2}=\frac{2\pi k}{8}$ with $k\in\left\{0,1,2,\cdots,7\right\}$, the possible values of $2\int_{\gamma_{p}}a^{2}$ are $\left\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\right\}$. Similarly, the possible values of $3\int_{\sigma}b^{3}$ are $\left\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\right\}$. The delta function $\delta\left(2\int_{\gamma_{p}}a^{2}\right)$ selects $2\int_{\gamma_{p}}a^{2}=0$ among the values. Therefore, we can write these projectors as a linear combination of invertible symmetry operators,
$\mathbb{P}_{100}\times\mathbb{P}_{100}=\frac{1}{4}\sum_{l_{2}=0}^{3}e^{{\rm i}l_{2}\cdot 2\int_{\gamma_{p}}a^{2}}\times\frac{1}{4}\sum_{l_{3}=0}^{3}e^{{\rm i}l_{3}\cdot 3\int_{\sigma}b^{3}}.$ (107)
The projector in $L_{200}$ is $\mathbb{P}_{200}=\delta\left(4\int_{\gamma_{p}}a^{2}\right)\delta\left(6\int_{\sigma}b^{3}\right)$, where $4\int_{\gamma_{p}}a^{2}$ takes value in $\left\{0,\pi\right\}$ and $6\int_{\sigma}b^{3}$ takes value in $\left\{0,\pi\right\}$. Thus, we have
$\mathbb{P}_{200}=\frac{1}{2}\sum_{k_{2}=0}^{1}e^{{\rm i}k_{2}\cdot 4\int_{\gamma_{p}}a^{2}}\times\frac{1}{2}\sum_{k_{3}=0}^{1}e^{{\rm i}k_{3}\cdot 6\int_{\sigma}b^{3}}.$ (108)
By comparing $\mathbb{P}_{100}\times\mathbb{P}_{100}$ and $\mathbb{P}_{200}$, we find
$\frac{1}{4}\sum_{l_{2}=0}^{3}e^{{\rm i}l_{2}\cdot 2\int_{\gamma_{p}}a^{2}}=\frac{1}{2}\sum_{k_{2}=0}^{1}e^{{\rm i}k_{2}\cdot 4\int_{\gamma_{p}}a^{2}}\times\frac{1}{2}\left(1+e^{{\rm i}2\int_{\gamma_{p}}a^{2}}\right)$
and
$\frac{1}{4}\sum_{l_{3}=0}^{3}e^{{\rm i}l_{3}\cdot 3\int_{\sigma}b^{3}}=\frac{1}{2}\sum_{k_{3}=0}^{1}e^{{\rm i}k_{3}\cdot 6\int_{\sigma}b^{3}}\times\frac{1}{2}\left(1+e^{{\rm i}3\int_{\sigma}b^{3}}\right).$
In other words,
$\mathbb{P}_{100}\times\mathbb{P}_{100}$
$=$ $\mathbb{P}_{200}\times\frac{1}{4}\left(1+e^{{\rm i}2\int_{\gamma_{p}}a^{2}}+e^{{\rm i}3\int_{\sigma}b^{3}}+e^{{\rm i}2\int_{\gamma_{p}}a^{2}+{\rm i}3\int_{\sigma}b^{3}}\right),$ (109)
and the fusion rule is
$L_{100}\times L_{100}$
$=$ $\frac{1}{4}\left(L_{200}+U_{020}L_{200}+U_{003}L_{200}+U_{023}L_{200}\right).$ (110)
Next, we consider $L_{100}\times L_{300}$, with $L_{300}=\mathcal{U}_{300}\delta\left(6\int_{\gamma_{p}}a^{2}\right)\delta\left(9\int_{\sigma}b^{3}\right)$. The possible values of $6\int_{\gamma_{p}}a^{2}$ are $\left\{0,\frac{3\pi}{2},3\pi,\frac{9\pi}{2}\right\}$, which are equivalent to $\left\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\right\}$. Thus, $\delta\left(6\int_{\gamma_{p}}a^{2}\right)=\frac{1}{4}\sum_{l_{2}=0}^{3}e^{{\rm i}l_{2}\cdot 6\int_{\gamma_{p}}a^{2}}$. As for $9\int_{\sigma}b^{3}$, it takes values in $\left\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\right\}$, the same as $3\int_{\sigma}b^{3}$. Therefore, we find
$\delta\left(6\int_{\gamma_{p}}a^{2}\right)=\delta\left(2\int_{\gamma_{p}}a^{2}\right),\ \delta\left(9\int_{\sigma}b^{3}\right)=\delta\left(3\int_{\sigma}b^{3}\right).$ (111)
Since $\mathcal{U}_{100}\times\mathcal{U}_{300}=1$, we obtain $L_{100}\times L_{300}=\delta\left(2\int_{\gamma_{p}}a^{2}\right)\delta\left(3\int_{\sigma}b^{3}\right)$. Expanding the delta functions, we have the fusion rule
$L_{100}\times L_{300}=\left(\frac{1}{4}\right)^{2}\sum_{l_{2},l_{3}=0}^{3}U_{0\left(2l_{2}\right)\left(3l_{3}\right)}.$ (112)
We now turn to a more involved fusion, $L_{123}\times L_{234}$, where $L_{123}=\mathcal{U}_{123}\mathbb{P}_{123}$ and $L_{234}=\mathcal{U}_{234}\mathbb{P}_{234}$. Substituting the parameters, we obtain the delta functions in $L_{123}$ and $L_{234}$, respectively,
$\mathbb{P}_{123}=$ $\delta\left(2\int_{\gamma_{p}}a^{2}-2\int_{\gamma_{p}}a^{1}\right)$
$\times\delta\left(3\int_{\sigma}b^{3}-3\int_{\gamma_{p}}a^{1}\right)\delta\left(6\int_{\sigma}b^{3}-6\int_{\gamma_{p}}a^{2}\right)$ (113)
$\mathbb{P}_{234}=$ $\delta\left(4\int_{\gamma_{p}}a^{2}-3\int_{\gamma_{p}}a^{1}\right)$
$\times\delta\left(6\int_{\sigma}b^{3}-4\int_{\gamma_{p}}a^{1}\right)\delta\left(9\int_{\sigma}b^{3}-8\int_{\gamma_{p}}a^{2}\right).$ (114)
We analyze these delta functions to obtain the fusion rule. We denote $\int_{\gamma_{p}}a^{1}=\frac{2\pi k_{1}}{4}$, $\int_{\gamma_{p}}a^{2}=\frac{2\pi k_{2}}{8}$, and $\int_{\sigma}b^{3}=\frac{2\pi k_{3}}{12}$, with $k_{1}\in\mathbb{Z}/4\mathbb{Z}$, $k_{2}\in\mathbb{Z}/8\mathbb{Z}$, and $k_{3}\in\mathbb{Z}/12\mathbb{Z}$. The delta functions in $L_{123}$ mean that the following equations must hold: $\frac{2\pi p}{N_{123}}\left(k_{2}-2k_{1}\right)=0\mod{2\pi}$, $\frac{2\pi p}{N_{123}}\left(k_{3}-3k_{1}\right)=0\mod{2\pi}$, and $\frac{2\pi p}{N_{123}}\left(2k_{3}-3k_{2}\right)=0\mod{2\pi}$. Likewise, the delta functions in $L_{234}$ enforce $\frac{2\pi p}{N_{123}}\left(2k_{2}-3k_{1}\right)=0\mod{2\pi}$, $\frac{2\pi p}{N_{123}}\left(2k_{3}-4k_{1}\right)=0\mod{2\pi}$, and $\frac{2\pi p}{N_{123}}\left(3k_{3}-4k_{2}\right)=0\mod{2\pi}$. Thus, $\mathbb{P}_{123}\times\mathbb{P}_{234}$ corresponds to the solutions of the above six equations. In our setup, $N_{123}=4$ and $p=1$, so we are solving integer equations defined modulo $N_{123}$, e.g., $k_{2}-2k_{1}=0\mod{N_{123}}$. The solution set is $k_{1}=0$, $k_{2}\in\left\{0,4\right\}$, and $k_{3}\in\left\{0,4,8\right\}$. In other words, the gauge field configurations should satisfy $\int_{\gamma_{p}}a^{1}=0$, $\int_{\gamma_{p}}a^{2}\in\left\{0,\pi\right\}$, and $\int_{\sigma}b^{3}\in\left\{0,\frac{2\pi}{3},\frac{4\pi}{3}\right\}$. Such configurations are exactly selected by $\delta\left(\int_{\gamma_{p}}a^{1}\right)$, $\delta\left(2\int_{\gamma_{p}}a^{2}\right)$, and $\delta\left(3\int_{\sigma}b^{3}\right)$, respectively. This means that
$\mathbb{P}_{123}\times\mathbb{P}_{234}$ $=\delta\left(\int_{\gamma_{p}}a^{1}\right)\delta\left(2\int_{\gamma_{p}}a^{2}\right)\delta\left(3\int_{\sigma}b^{3}\right)$
$=\left(\frac{1}{4}\right)^{3}\sum_{l_{1},l_{2},l_{3}=0}^{3}e^{{\rm i}\left(l_{1}\int_{\gamma_{p}}a^{1}+2l_{2}\int_{\gamma_{p}}a^{2}+3l_{3}\int_{\sigma}b^{3}\right)}$ (115)
According to Eq. (93), the fusion channels contain $L_{357}=\mathcal{U}_{357}\mathbb{P}_{357}$, where
$\mathbb{P}_{357}=$ $\delta\left(6\int_{\gamma_{p}}a^{2}-5\int_{\gamma_{p}}a^{1}\right)$
$\times\delta\left(9\int_{\sigma}b^{3}-7\int_{\gamma_{p}}a^{1}\right)\delta\left(15\int_{\sigma}b^{3}-14\int_{\gamma_{p}}a^{2}\right).$ (116)
The delta functions in $L_{357}$ enforce $\frac{2\pi p}{N_{123}}\left(3k_{2}-5k_{1}\right)=0$, $\frac{2\pi p}{N_{123}}\left(3k_{3}-7k_{1}\right)=0$, and $\frac{2\pi p}{N_{123}}\left(5k_{3}-7k_{2}\right)=0$, all of which are defined modulo $2\pi$. The independent equations are $k_{3}-k_{1}=0\mod 4$ and $k_{2}-3k_{1}=0\mod 4$. Using $4\int_{\gamma_{p}}a^{1}\in 2\pi\mathbb{Z}$, $8\int_{\gamma_{p}}a^{2}\in 2\pi\mathbb{Z}$, and $12\int_{\sigma}b^{3}\in 2\pi\mathbb{Z}$, we can simplified $\mathbb{P}_{357}$ as
$\mathbb{P}_{357}=$ $\delta\left(3\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{1}\right)\delta\left(2\int_{\gamma_{p}}a^{2}-3\int_{\gamma_{p}}a^{1}\right)$
$=$ $\left(\frac{1}{4}\right)^{2}\sum_{h_{1},h_{2}=0}^{3}e^{{\rm i}\left[3h_{1}\int_{\sigma}b^{3}-h_{1}\int_{\gamma_{p}}a^{1}+2h_{2}\int_{\gamma_{p}}a^{2}-3h_{2}\int_{\gamma_{p}}a^{1}\right]}.$ (117)
The second line uses the fact that $3\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{1}$ and $2\int_{\gamma_{p}}a^{2}-3\int_{\gamma_{p}}a^{1}$ both take values in $\left\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\right\}$. For convenience, we denote $\mathbb{P}_{357}=\left(\frac{1}{4}\right)^{2}A$ where
$A=\sum_{h_{1},h_{2}=0}^{3}e^{{\rm i}\left[3h_{1}\int_{\sigma}b^{3}+2h_{2}\int_{\gamma_{p}}a^{2}-\left(h_{1}+3h_{2}\right)\int_{\gamma_{p}}a^{1}\right]}.$
We expand the summation explicitly and find that
$A\times\sum_{l=0}^{3}e^{{\rm i}l\cdot\int_{\gamma_{p}}a^{1}}$ $=\sum_{l_{1},l_{2},l_{3}=0}^{3}e^{{\rm i}\left(l_{1}\int_{\gamma_{p}}a^{1}+2l_{2}\int_{\gamma_{p}}a^{2}+3l_{3}\int_{\sigma}b^{3}\right)}.$ (118)
Therefore, $\mathbb{P}_{123}\times\mathbb{P}_{234}=\mathbb{P}_{357}\times\frac{1}{4}\sum_{l=0}^{3}e^{{\rm i}l\cdot\int_{\gamma_{p}}a^{1}}$, and the fusion rule is
$L_{123}\times L_{234}=\frac{1}{4}\sum_{l_{1}=0}^{3}U_{l_{1}00}L_{357}.$ (119)
This example indicates that, for a general gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$, the greatest common divisor $N_{123}$ controls the fusion rules (in the next example, we will also see the effect of $p$ in the fusion rules).
IV.4.3 $N_{1}=4$, $N_{2}=8$, $N_{3}=12$, $N_{123}=4$, $p=2$
In this example, we set $p=2$. Notice that the parameter $p$ and $N_{123}$ together determine the expansions of delta functions. To see this explicitly, we again consider the symmetry operators $L_{100}$, $L_{\left(N_{1}-1\right)00}$, $L_{123}$, and $L_{234}$.
The fusion of two $L_{100}$’s is still
$L_{100}\times L_{100}=\mathcal{U}_{200}\delta\left(\frac{pN_{2}}{N_{123}}\int_{\gamma_{p}}a^{2}\right)\delta\left(\frac{pN_{3}}{N_{123}}\int_{\sigma}b^{3}\right).$ (120)
But now, with $p=2$ and $N_{123}=4$, these delta functions are $\mathbb{P}_{100}\times\mathbb{P}_{100}=\delta\left(4\int_{\gamma_{p}}a^{2}\right)\delta\left(6\int_{\sigma}b^{3}\right)$. These delta functions require that $\int_{\gamma_{p}}a^{2}\in\left\{0,\frac{4\pi}{8},\frac{8\pi}{8},\frac{12\pi}{8}\right\}$ and $\int_{\sigma}b^{3}\in\left\{0,\frac{4\pi}{12},\frac{8\pi}{12},\cdots,\frac{22\pi}{12}\right\}$, and $4\int_{\gamma_{p}}a^{2}\in\left\{0,\pi\right\}$ and $6\int_{\sigma}b^{3}\in\left\{0,\pi\right\}$. Therefore, we can express the delta functions as $\frac{1}{2}\sum_{l_{2}=0}^{1}e^{{\rm i}l_{2}\cdot 4\int_{\gamma_{p}}a^{2}}\times\frac{1}{2}\sum_{l_{3}=0}^{1}e^{{\rm i}l_{3}\cdot 6\int_{\sigma}b^{3}}$. On the other hand, for $L_{200}=\mathcal{U}_{200}\delta\left(\frac{2pN_{2}}{N_{123}}\int_{\gamma_{p}}a^{2}\right)\delta\left(\frac{2pN_{3}}{N_{123}}\int_{\sigma}b^{3}\right)$, the projectors are $\mathbb{P}_{200}=\delta\left(8\int_{\gamma_{p}}a^{2}\right)\delta\left(12\int_{\sigma}b^{3}\right)$, which impose no constraints on the values of $\int_{\gamma_{p}}a^{2}$ and $\int_{\sigma}b^{3}$. This is because $8\int_{\gamma_{p}}a^{2}$ and $12\int_{\sigma}b^{3}$ always equal $0\mod 2\pi$, and the delta functions always return $1$. Therefore,
$\mathbb{P}_{100}\times\mathbb{P}_{100}=\mathbb{P}_{200}\times\frac{1}{2}\sum_{l_{2}=0}^{1}e^{{\rm i}l_{2}\cdot 4\int_{\gamma_{p}}a^{2}}\times\frac{1}{2}\sum_{l_{3}=0}^{1}e^{{\rm i}l_{3}\cdot 6\int_{\sigma}b^{3}}$ (121)
and
$L_{100}\times L_{100}=\frac{1}{4}\left(L_{200}+U_{040}L_{200}+U_{006}L_{200}+U_{046}L_{200}\right).$ (122)
We can see the difference in the fusion rules when compared to Eq. (110) in the $p=1$ setup.
Next, we consider $L_{100}\times L_{300}$ where $L_{300}=\mathcal{U}_{300}\delta\left(\frac{3pN_{2}}{N_{123}}\int_{\gamma_{p}}a^{2}\right)\delta\left(\frac{3pN_{3}}{N_{123}}\int_{\sigma}b^{3}\right)$. As $p=2$, the projectors are $\mathbb{P}_{300}=\delta\left(12\int_{\gamma_{p}}a^{2}\right)\delta\left(18\int_{\sigma}b^{3}\right)$, which is actually $\mathbb{P}_{300}=\delta\left(4\int_{\gamma_{p}}a^{2}\right)\delta\left(6\int_{\sigma}b^{3}\right)$ since $8\int_{\gamma_{p}}a^{2}\in 2\pi\mathbb{Z}$ and $12\int_{\sigma}b^{3}\in 2\pi\mathbb{Z}$. When $p=2$, $L_{100}$ happens to share the same projectors with $L_{300}$. Therefore, when $p=2$,
$L_{100}\times L_{300}=\left(\frac{1}{2}\right)^{2}\sum_{l_{2},l_{3}=0}^{1}U_{0\left(4l_{2}\right)\left(6l_{3}\right)}.$ (123)
Again, we can see the difference when compared to Eq. (112) in the $p=1$ setup.
The third example in the $p=2$ setup is $L_{123}\times L_{234}$. The key to obtaining the fusion rule is the delta functions. Substituting $p=2$ into the expressions of $L_{123}$ and $L_{234}$, the delta functions are
$\mathbb{P}_{123}=$ $\delta\left(4\int_{\gamma_{p}}a^{2}-4\int_{\gamma_{p}}a^{1}\right)$
$\times\delta\left(6\int_{\sigma}b^{3}-6\int_{\gamma_{p}}a^{1}\right)\delta\left(12\int_{\sigma}b^{3}-12\int_{\gamma_{p}}a^{2}\right),$
$\mathbb{P}_{234}=$ $\delta\left(8\int_{\gamma_{p}}a^{2}-6\int_{\gamma_{p}}a^{1}\right)$
$\times\delta\left(12\int_{\sigma}b^{3}-8\int_{\gamma_{p}}a^{1}\right)\delta\left(18\int_{\sigma}b^{3}-16\int_{\gamma_{p}}a^{2}\right).$
Because $4\int_{\gamma_{p}}a^{1}\in 2\pi\mathbb{Z}$, $8\int_{\gamma_{p}}a^{2}\in 2\pi\mathbb{Z}$, and $12\int_{\sigma}b^{3}\in 2\pi\mathbb{Z}$, we can simplify these delta functions as
$\mathbb{P}_{123}=$ $\delta\left(4\int_{\gamma_{p}}a^{2}\right)\delta\left(6\int_{\sigma}b^{3}-2\int_{\gamma_{p}}a^{1}\right),$ (124)
$\mathbb{P}_{234}=$ $\delta\left(2\int_{\gamma_{p}}a^{1}\right)\delta\left(6\int_{\sigma}b^{3}\right).$ (125)
On the other hand, the delta functions in $L_{357}$ are
$\mathbb{P}_{357}=$ $\delta\left(12\int_{\gamma_{p}}a^{2}-10\int_{\gamma_{p}}a^{1}\right)$
$\times\delta\left(18\int_{\sigma}b^{3}-14\int_{\gamma_{p}}a^{1}\right)\delta\left(30\int_{\sigma}b^{3}-28\int_{\gamma_{p}}a^{2}\right)$ (126)
which can be simplified as
$\mathbb{P}_{357}=$ $\delta\left(4\int_{\gamma_{p}}a^{2}-2\int_{\gamma_{p}}a^{1}\right)\delta\left(6\int_{\sigma}b^{3}-2\int_{\gamma_{p}}a^{1}\right).$ (127)
In order to determine the fusion rule, we need to match $\mathbb{P}_{123}\times\mathbb{P}_{234}$ to $\mathbb{P}_{357}$. In $\mathbb{P}_{123}\times\mathbb{P}_{234}$, only $\delta\left(4\int_{\gamma_{p}}a^{2}\right)$, $\delta\left(2\int_{\gamma_{p}}a^{1}\right)$, and $\delta\left(6\int_{\sigma}b^{3}\right)$ are independent, since $\delta\left(6\int_{\sigma}b^{3}-2\int_{\gamma_{p}}a^{1}\right)$ automatically holds in the presence of $\delta\left(2\int_{\gamma_{p}}a^{1}\right)\delta\left(6\int_{\sigma}b^{3}\right)$. Therefore,
$\mathbb{P}_{123}\times\mathbb{P}_{234}=$ $\left(\frac{1}{2}\right)^{3}\sum_{l_{1},l_{2},l_{3}=0}^{1}e^{{\rm i}\left(l_{1}\cdot 2\int_{\gamma_{p}}a^{1}+l_{2}\cdot 4\int_{\gamma_{p}}a^{2}+l_{3}\cdot 6\int_{\sigma}b^{3}\right)}.$ (128)
We now turn to $\mathbb{P}_{357}$. Suppose $\int_{\gamma_{p}}a^{1}=\frac{2\pi k_{1}}{N_{1}}$, $\int_{\gamma_{p}}a^{2}=\frac{2\pi k_{2}}{N_{2}}$, and $\int_{\sigma}b^{3}=\frac{2\pi k_{3}}{N_{3}}$, where $k_{i}\in\left\{0,1,\cdots,N_{i}-1\right\}$. We have $4\int_{\gamma_{p}}a^{2}-2\int_{\gamma_{p}}a^{1}=\pi\left(k_{2}-k_{1}\right)$ and $6\int_{\sigma}b^{3}-2\int_{\gamma_{p}}a^{1}=\pi\left(k_{3}-k_{1}\right)$. By exhausting all combinations of $\left\{k_{1},k_{2},k_{3}\right\}$, we find that $4\int_{\gamma_{p}}a^{2}-2\int_{\gamma_{p}}a^{1}$ and $6\int_{\sigma}b^{3}-2\int_{\gamma_{p}}a^{1}$ both take values 0 or $\pi$ with equal probability. The effect of $\mathbb{P}_{357}$ is to select those combinations with $k_{2}-k_{1}=0\mod 2$ and $k_{3}-k_{1}=0\mod 2$. Expressing the delta functions in terms of invertible symmetry operators yields
$\mathbb{P}_{357}=$ $\frac{1}{2}\left(1+e^{{\rm i}4\int_{\gamma_{p}}a^{2}-{\rm i}2\int_{\gamma_{p}}a^{1}}\right)$
$\times\frac{1}{2}\left(1+e^{{\rm i}6\int_{\sigma}b^{3}-{\rm i}2\int_{\gamma_{p}}a^{1}}\right).$ (129)
By comparing $\mathbb{P}_{123}\times\mathbb{P}_{234}$ and $\mathbb{P}_{357}$, we find
$\mathbb{P}_{123}\times\mathbb{P}_{234}=\mathbb{P}_{357}\times\frac{1}{2}\left(1+e^{{\rm i}2\int_{\gamma_{p}}a^{1}}\right),$ (130)
which implies the fusion rule
$L_{123}\times L_{234}=\frac{1}{2}\left(L_{357}+U_{200}L_{357}\right).$ (131)
We summarize the fusion rules for different values of $p$ in Table 2. The fusion rules of invertible symmetry operators are group-like, while those of non-invertible symmetry operators depend on the fusion of the projectors (delta functions). For a general effective field theory $S_{aab}$ with gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$ and a level-$p$ twisted term, the greatest common divisor $N_{123}$ and $p$ determine how a projector can be expressed as a sum of invertible operators in the context of correlation functions, which in turn determines the fusion of the projectors.
V Anomalies of (non-)invertible higher-form symmetries
In this section, we analyze the anomalies of the generalized symmetries identified above. A symmetry is said to be anomalous if it cannot be gauged. At the field-theoretical level, gauging means coupling the symmetry current to a background gauge field and summing over all background-field configurations. After gauging, the original global symmetry is promoted to a gauge redundancy of the new theory.
For the $1$- and $2$-form symmetries studied in this paper—including both invertible and non-invertible ones—we find mixed anomalies, in the sense that certain pairs of symmetries cannot be gauged simultaneously. We diagnose a mixed anomaly by coupling the currents of two symmetries to their respective background fields and checking whether the resulting action remains invariant under background gauge transformations. For certain combinations, the total action fails to be gauge invariant, signaling a mixed anomaly. Furthermore, we find that in some cases the mixed anomaly can be canceled by a bulk term in one higher dimension, whereas in other cases it cannot.
Below, we first illustrate the mixed anomaly between two higher-form symmetries using the 33d toric code and its effective field theory. We then investigate the mixed anomalies among the generalized higher-form symmetries in the BR topological order described by $S_{aab}$ in Sec. III.
Table 3: Mixed anomalies of generalized symmetries in the effective field theory $S_{aab}$. $J_{e}^{1}=-*a^{1}$, $J_{e}^{2}=-*a^{2}$, and $J_{e}^{3}=*b^{3}$ are conserved currents that generate $2$-, 22-, and $1$-form invertible symmetries, respectively. $J_{m}^{1}=*\left(b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}\right)$, $J_{m}^{2}=*\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right)$, and $J_{m}^{3}=-*\left(a^{3}+\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\phi^{1}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi^{2}\right)$ are conserved currents that generate $1$-, 11-, and $2$-form non-invertible symmetries, where $d\phi^{1}=\frac{pN_{1}}{N_{123}}a^{1}$, $d\phi^{2}=\frac{pN_{2}}{N_{123}}a^{2}$, and $d\beta^{3}=\frac{pN_{3}}{N_{123}}b^{3}$. To gauge these generalized symmetries, we couple the currents to background fields: 22-form (33-form) background fields are denoted by $\mathcal{B}$ ($\mathcal{C}$). This table lists the outcomes of gauging different pairs of generalized symmetries. The mark $\checkmark$” indicates that there is no mixed anomaly for the corresponding pair. A higher-dimensional bulk term, e.g., $\int_{M_{5}}\frac{1}{2\pi N_{1}}\mathcal{B}_{m}^{1}\mathcal{C}_{e}^{1}$, indicates a mixed anomaly that can be canceled by anomaly inflow from a (4+1)(4+1)D bulk. The mark $\times$” indicates a mixed anomaly manifested by a lack of gauge invariance that cannot be canceled by adding any higher-dimensional bulk term.
Coupling terms $\int\frac{1}{2\pi}\mathcal{C}_{e}^{1}\wedge*J_{e}^{1}$ $\int\frac{1}{2\pi}\mathcal{C}_{e}^{2}\wedge*J_{e}^{2}$ $\int\frac{1}{2\pi}\mathcal{B}_{e}^{3}\wedge*J_{e}^{3}$ $\int\frac{1}{2\pi}\mathcal{B}_{m}^{1}\wedge*J_{m}^{1}$ $\int\frac{1}{2\pi}\mathcal{B}_{m}^{2}\wedge*J_{m}^{2}$ $\int\frac{1}{2\pi}\mathcal{C}_{m}^{3}\wedge*J_{m}^{3}$
$\int\frac{1}{2\pi}\mathcal{C}_{e}^{1}\wedge*J_{e}^{1}$ $\checkmark$ $\checkmark$ $\checkmark$ $\int_{M_{5}}\frac{1}{2\pi N_{1}}\mathcal{B}_{m}^{1}\mathcal{C}_{e}^{1} \checkmark \checkmark$
$\int\frac{1}{2\pi}\mathcal{C}_{e}^{2}\wedge*J_{e}^{2}$ $\checkmark$ $\checkmark$ $\checkmark$ $\int_{M_{5}}\frac{1}{2\pi N_{2}}\mathcal{B}_{m}^{2}\mathcal{C}_{e}^{2} \checkmark$
$\int\frac{1}{2\pi}\mathcal{B}_{e}^{3}\wedge*J_{e}^{3}$ $\checkmark$ $\checkmark$ $\checkmark$ $\int_{M_{5}}\frac{1}{2\pi N_{3}}\mathcal{B}_{e}^{3}\mathcal{C}_{m}^{3}$
$\int\frac{1}{2\pi}\mathcal{B}_{m}^{1}\wedge*J_{m}^{1}$ $\checkmark$ $\times$ $\times$
$\int\frac{1}{2\pi}\mathcal{B}_{m}^{2}\wedge*J_{m}^{2}$ $\checkmark$ $\times$
$\int\frac{1}{2\pi}\mathcal{C}_{m}^{3}\wedge*J_{m}^{3}$ $\checkmark$
V.1 Mixed anomalies in the 33d $\mathbb{Z}_{2}$ topological order (the toric code)
In this subsection, we use the 33d toric code as a benchmark example to show how its $1$-form and $2$-form symmetries exhibit a mixed anomaly and how the anomaly can be canceled by a bulk term. We start from its effective field theory $S_{\text{3dTC}}=\int\frac{2}{2\pi}bda$. As reviewed in Sec. III.1, the EoMs $\frac{2}{2\pi}da=0$ and $\frac{2}{2\pi}db=0$ can be viewed as continuity equations, with conserved currents $J_{e}=-*a$ and $J_{m}=*b$. These currents generate a $1$-form symmetry with symmetry operator $U_{m}\left(\sigma\right)=e^{{\rm i}\int_{\sigma}b}$ and a $2$-form symmetry with symmetry operator $U_{e}\left(\gamma\right)=e^{{\rm i}\int_{\gamma}a}$.
To gauge the $1$-form symmetry, we couple $J_{m}$ to a $2$-form background field $\mathcal{B}_{m}$ via $\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}\wedge*J_{m}=\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}\wedge b$. The total action (with $\wedge$ omitted) becomes
$S^{m}_{\text{gauged}}=\int_{M_{4}}\frac{2}{2\pi}bda+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}b.$ (132)
Under the background gauge transformation $\mathcal{B}_{m}\rightarrow\mathcal{B}_{m}+dV_{m}$, the action is invariant provided that $a\rightarrow a-\frac{1}{2}V_{m}$.
If we instead gauge the $2$-form symmetry, the coupling term is $\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}\wedge*J_{e}=\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}\wedge a$, where $\mathcal{C}_{e}$ is a $3$-form background field. The total action is
$S_{\text{gauged}}^{e}=\int_{M_{4}}\frac{2}{2\pi}bda+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}a.$ (133)
If $\mathcal{C}_{e}$ transforms as $\mathcal{C}_{e}\rightarrow\mathcal{C}_{e}+dT_{e}$, then $S_{\text{gauged}}^{e}$ is invariant under $b\rightarrow b+\frac{1}{2}T_{e}$.
We now ask what happens if we attempt to gauge the $1$-form and $2$-form symmetries simultaneously. Including both coupling terms, we obtain
$S_{\text{gauged}}^{m,e}=\int_{M_{4}}\frac{2}{2\pi}bda+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}b+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}a.$ (134)
Under the gauge transformations $\begin{cases}\mathcal{B}_{m}\rightarrow&\mathcal{B}_{m}+dV_{m}\\ a\rightarrow&a-\frac{1}{2}V_{m}\end{cases}$ and $\begin{cases}\mathcal{C}_{e}\rightarrow&\mathcal{C}_{e}+dT_{e}\\ b\rightarrow&b+\frac{1}{2}T_{e}\end{cases}$, the action acquires an additional term
$\Delta S_{\text{gauged}}^{m,e}=\int_{M_{4}}\frac{1}{4\pi}\mathcal{B}_{m}T_{e}-\frac{1}{4\pi}\mathcal{C}_{e}V_{m}-\frac{1}{4\pi}dT_{e}V_{m}.$ (135)
The violation of gauge invariance shows that the two symmetries cannot be gauged simultaneously; hence the 33d toric code exhibits a mixed anomaly between its $1$-form and $2$-form symmetries.
This mixed anomaly can be canceled by coupling the theory to a (4+1)(4+1)D bulk,
$S_{\text{bulk}}=-\int_{M_{5}}\frac{1}{4\pi}\mathcal{B}_{m}\mathcal{C}_{e}.$ (136)
Assuming that the background fields are locally flat, $d\mathcal{B}_{m}=0$ and $d\mathcal{C}_{e}=0$, one can verify that $S_{\text{gauged}}^{m,e}+S_{\text{bulk}}$ is gauge invariant under the above transformations.
V.2 Exploration of mixed anomalies of higher-form symmetries in Borromean-rings topological order
In Sec. III we identified the generalized symmetries of the effective field theory (14) and the corresponding conserved currents. Here we couple these currents to background fields to gauge the symmetries. For simplicity, we focus on the “minimal” conserved currents, since all other symmetry operators can be generated by combining them.
Among the minimal currents, $J_{e}^{1}=-*a^{1}$, $J_{e}^{2}=-*a^{2}$, and $J_{e}^{3}=*b^{3}$ generate $2$-, 22-, and $1$-form invertible symmetries, respectively; $J_{m}^{1}=*\left(b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}\right)$, $J_{m}^{2}=*\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right)$, and $J_{m}^{3}=-*\left(a^{3}+\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\phi^{1}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi^{2}\right)$ generate $1$-, 11-, and $2$-form non-invertible symmetries, where $d\phi^{1}=\frac{pN_{1}}{N_{123}}a^{1}$, $d\phi^{2}=\frac{pN_{2}}{N_{123}}a^{2}$, and $d\beta^{3}=\frac{pN_{3}}{N_{123}}b^{3}$. For $1$-form (22-form) symmetries, we couple them to $2$-form (33-form) background fields denoted by $\mathcal{B}$ ($\mathcal{C}$). As in the 33d toric code example, we first gauge one symmetry at a time and then attempt to gauge two symmetries simultaneously to diagnose possible mixed anomalies.
We begin by gauging a single higher-form symmetry. For example, to gauge the $2$-form symmetry generated by $J_{e}^{1}$, we couple it to a background field $\mathcal{C}_{e}^{1}$ via $\mathcal{C}_{e}^{1}\wedge*J_{e}^{1}=\mathcal{C}_{e}^{1}\wedge a^{1}$. The total action is
$S^{e_{1}}_{\text{gauged}}=\int_{M_{4}}\sum_{i=1}^{3}\frac{N_{i}}{2\pi}b^{i}da^{i}+qa^{1}a^{2}b^{3}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{1}a^{1},$ (137)
which is invariant under $\mathcal{C}_{e}^{1}\rightarrow\mathcal{C}_{e}^{1}+dT_{e}^{1}$ together with $b^{1}\rightarrow b^{1}+\frac{1}{N_{1}}T_{e}^{1}$. This preserved gauge invariance indicates no anomaly for this $2$-form symmetry at the level of the effective action.
To gauge the $1$-form symmetry generated by $J_{m}^{1}$, we add the coupling term $\mathcal{B}_{m}^{1}\wedge*J_{m}^{1}$ to $S_{aab}$,
$S_{\text{gauged}}^{m_{1}}=$ $\int_{M^{4}}\sum_{i=1}^{3}\frac{N_{i}}{2\pi}b^{i}da^{i}+qa^{1}a^{2}b^{3}$
$+\frac{1}{2\pi}\mathcal{B}_{m}^{1}\left(b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}\right).$ (138)
The action $S_{\text{gauged}}^{m_{1}}$ is invariant under $\mathcal{B}_{m}^{1}\rightarrow\mathcal{B}_{m}^{1}+dV_{m}^{1}$ together with $a^{1}\rightarrow a^{1}-\frac{1}{N_{1}}V_{m}^{1}$. Again, gauge invariance implies no anomaly for this $1$-form symmetry.
Similarly, one can gauge the remaining minimal currents. The total action with the corresponding coupling term is always invariant under suitable background gauge transformations. For later reference, we list the coupling terms and their associated gauge transformations:
$S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{2}\wedge*J_{e}^{2},\quad\begin{cases}\mathcal{C}_{e}^{2}\rightarrow&\mathcal{C}_{e}^{2}+dT_{e}^{2}\\ b^{2}\rightarrow&b^{2}+\frac{1}{N_{2}}T_{e}^{2}\end{cases};$ (139)
$S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}^{2}\wedge*J_{m}^{2},\quad\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases};$ (140)
$S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{e}^{3}\wedge*J_{e}^{3},\quad\begin{cases}\mathcal{B}_{e}^{3}\rightarrow&\mathcal{B}_{e}^{3}+dV_{e}^{3}\\ a^{3}\rightarrow&a^{3}-\frac{1}{N_{3}}V_{e}^{3}\end{cases};$ (141)
$S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{m}^{3}\wedge*J_{m}^{3},\quad\begin{cases}\mathcal{C}_{m}^{3}\rightarrow&\mathcal{C}_{m}^{3}+dT_{m}^{3}\\ b^{3}\rightarrow&b^{3}+\frac{1}{N_{3}}T_{m}^{3}\end{cases}.$ (142)
We now consider gauging two higher-form symmetries simultaneously by adding two coupling terms to $S_{aab}$. All combinations are summarized in Table 3. Some pairs can be gauged simultaneously and are free of mixed anomalies. In other cases, a mixed anomaly arises because the gauged action fails to remain gauge invariant. Some mixed anomalies can be canceled by coupling to a (4+1)(4+1)D bulk term, while others cannot.
Gauging two invertible symmetries is always free of mixed anomalies. For example, gauging the symmetries generated by $J_{e}^{1}$ and $J_{e}^{2}$ leads to
$S_{\text{gauged}}^{e_{1},e_{2}}=$ $S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{1}a^{1}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{2}a^{2},$ (143)
which is invariant under $\mathcal{C}_{e}^{1}\rightarrow\mathcal{C}_{e}^{1}+dT_{e}^{1}$, $\mathcal{C}_{e}^{2}\rightarrow\mathcal{C}_{e}^{2}+dT_{e}^{2}$ together with $b^{1}\rightarrow b^{1}+\frac{1}{N_{1}}T_{e}^{1}$ and $b^{2}\rightarrow b^{2}+\frac{1}{N_{2}}T_{e}^{2}$.
There is no mixed anomaly when gauging one invertible symmetry together with one non-invertible symmetry, provided that their currents do not involve gauge charges/fluxes of the same $\mathbb{Z}_{N_{i}}$ factor. For example, it is consistent to gauge the invertible symmetry generated by $J_{e}^{1}$ and the non-invertible symmetry generated by $J_{m}^{2}$. Indeed,
$S_{\text{gauged}}^{e_{1},m_{2}}=$ $S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{1}a^{1}$
$+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}^{2}\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right),$ (144)
remains gauge invariant under $\mathcal{C}_{e}^{1}\rightarrow\mathcal{C}_{e}^{1}+dT_{e}^{1}$ with $b^{1}\rightarrow b^{1}+\frac{1}{N_{1}}T_{e}^{1}$, and $\mathcal{B}_{m}^{2}\rightarrow\mathcal{B}_{m}^{2}+dV_{m}^{2}$ with $a^{2}\rightarrow a^{2}-\frac{1}{N_{2}}V_{m}^{2}$.
An invertible symmetry and a non-invertible symmetry can have a mixed anomaly if their currents involve gauge charges/fluxes of the same $\mathbb{Z}_{N_{i}}$ factor; such a mixed anomaly can nevertheless be canceled by a bulk term. For example, gauging the invertible symmetry generated by $J_{e}^{3}$ together with the non-invertible symmetry generated by $J_{m}^{3}$ is described by
$S_{\text{gauged}}^{e_{3},m_{3}}=$ $S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{e}^{3}b^{3}$
$+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{m}^{3}\left(a^{3}+\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\phi^{1}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi^{2}\right).$ (145)
Under the gauge transformations $\begin{cases}\mathcal{B}_{e}^{3}\rightarrow&\mathcal{B}_{e}^{3}+dV_{e}^{3}\\ a^{3}\rightarrow&a^{3}-\frac{1}{N_{3}}V_{e}^{3}\end{cases}$ and $\begin{cases}\mathcal{C}_{m}^{3}\rightarrow&\mathcal{C}_{m}^{3}+dT_{m}^{3}\\ b^{3}\rightarrow&b^{3}+\frac{1}{N_{3}}T_{m}^{3}\end{cases}$,
$\Delta S_{\text{gauged}}^{e_{3},m_{3}}=$ $\int_{M_{4}}\frac{1}{2\pi N_{3}}\mathcal{B}_{e}^{3}T_{m}^{3}-\frac{1}{2\pi N_{3}}\mathcal{C}_{m}^{3}V_{e}^{3}-\frac{1}{2\pi N_{3}}dT_{m}^{3}V_{e}^{3}.$ (146)
The non-vanishing $\Delta S_{\text{gauged}}^{e_{3},m_{3}}$ indicates a mixed anomaly. Equivalently, once we gauge the symmetry generated by $J_{e}^{3}$, gauge invariance requires $a^{3}\rightarrow a^{3}-\frac{1}{N_{3}}V_{e}^{3}$, which modifies the continuity equation $d\left(*J_{m}^{3}\right)=0$ to $d\left(*J_{m}^{3}\right)-\frac{1}{N_{3}}dV_{e}^{3}=0$. Thus $J_{m}^{3}$ is no longer conserved: gauging $J_{e}^{3}$ explicitly breaks the symmetry generated by $J_{m}^{3}$, and the two symmetries cannot be gauged simultaneously.
This mixed anomaly can be canceled by coupling to a (4+1)(4+1)D bulk with action
$S_{\text{bulk}}=-\int_{M_{5}}\frac{1}{2\pi N_{3}}\mathcal{B}_{e}^{3}\mathcal{C}_{m}^{3},$ (147)
so that the combined action is gauge invariant under the same transformations.
Finally, we encounter mixed anomalies between two non-invertible symmetries. As an illustrative example, consider gauging the non-invertible symmetries generated by $J_{m}^{1}$ and $J_{m}^{2}$ by adding the coupling terms
$S_{\text{gauging}}^{m_{1},m_{2}}=$ $S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\left(b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}\right)$
$+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}^{2}\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right).$ (148)
Assuming that the two symmetries can be gauged simultaneously would require $S_{\text{gauging}}^{m_{1},m_{2}}$ to be invariant under suitable background gauge transformations. In the presence of the first coupling term, the EoM of $b^{1}$ is modified to $\frac{N_{1}}{2\pi}da^{1}+\mathcal{B}_{m}^{1}=0$, which must be preserved under $\mathcal{B}_{m}^{1}\rightarrow\mathcal{B}_{m}^{1}+dV_{m}^{1}$. This forces $a^{1}\rightarrow a^{1}-\frac{1}{N_{1}}V_{m}^{1}$. However, $d\left(*J_{m}^{2}\right)=0$ is no longer valid under this transformation: gauging the symmetry generated by $J_{m}^{1}$ breaks the symmetry generated by $J_{m}^{2}$. Similarly, gauging the symmetry generated by $J_{m}^{2}$ requires $a^{2}\rightarrow a^{2}-\frac{1}{N_{2}}V_{m}^{2}$ under $\mathcal{B}_{m}^{2}\rightarrow\mathcal{B}_{m}^{2}+dV_{m}^{2}$, which in turn induces $J_{m}^{1}\rightarrow*J_{m}^{1}-\frac{1}{2}\frac{V_{m}^{2}}{2\pi}\beta^{3}$ and violates $d\left(*J_{m}^{1}\right)=0$. Therefore, these two symmetries cannot be gauged simultaneously: gauging one of them necessarily breaks the other. To our knowledge, we cannot find a higher-dimensional bulk term to cancel this mixed anomaly, which is explained in Appendix C.
VI Summary and Outlook
VI.1 Summary
In this paper, we investigate generalized symmetries in the topological field theory of (3+1)(3+1)D BR topological order. Our strategy is to construct symmetry operators from conserved currents: we extract continuity equations and conserved currents from the equations of motion of the gauge fields and then write down the associated symmetry operators. Geometrically, these generalized symmetries are classified as $1$-form or $2$-form, depending on the dimension of the supporting submanifold. Algebraically, the symmetries include both invertible and non-invertible ones. The invertible higher-form symmetries exhibit a $\mathbb{Z}_{N_{i}}$-type group structure: their symmetry operators obey group-like fusion rules that match the composition law of group elements, as summarized in Sec. IV.3. Equivalently, each such operator admits an inverse whose fusion yields the identity operator. By contrast, the non-invertible symmetry operators (again of either $1$-form or $2$-form type) consist of a unitary operator dressed by a projector. The projector enforces the constraints under which the corresponding current is conserved. Since a projector has no inverse, the resulting symmetry is intrinsically non-invertible. Correspondingly, the fusion of non-invertible symmetry operators can admit multiple fusion channels, as illustrated in Sec. IV.4.
We also analyze anomalies of these generalized symmetries by testing whether they can be gauged. Gauging a single symmetry amounts to coupling its current to a background field in a gauge-invariant way. However, when two symmetries are gauged simultaneously, a mixed anomaly may arise: the action with both coupling terms fails to remain gauge invariant. In some cases, such a mixed anomaly can be canceled by anomaly inflow from a higher-dimensional bulk, so that the combined bulk–boundary system is gauge invariant. In other cases, the mixed anomaly is intrinsic and persists because no bulk term can restore gauge invariance; see Table 3.
Overall, our results provide a field-theoretical characterization of invertible and non-invertible higher-form symmetries in 33d topological order. We expect that the current-based construction developed here can be generalized to other 33d topological orders, including those described by quantum doubles of finite groups. In parallel, recent progress has established diagrammatic representations for 33d and 44d topological orders from continuum field theories, yielding stringent consistency conditions for higher-dimensional topological orders [78]. It would be interesting to further clarify how the fusion rules of the symmetry operators constructed here are constrained by, and in turn inform, these diagrammatic consistency conditions.
VI.2 Outlook
Several future directions are motivated by the present study. Throughout this work, our analysis is primarily formulated at the level of continuum field theory. An important next step is to formulate and test generalized symmetries in concrete lattice models that realize three-dimensional topological orders. Natural candidates include three-dimensional quantum double models, Hamiltonian lattice realizations of (3+1)(3+1)D Dijkgraaf–Witten theories, and three-dimensional generalizations of string-net constructions [9, 64, 81, 82]. Recently, the authors of this work investigated microscopic constructions of excitations, fusion processes, and shrinking operations in non-Abelian topological orders using the three-dimensional quantum double model [79]. Building on this line of work, it would be valuable to explicitly construct invertible and non-invertible higher-form symmetry operators on three-dimensional lattices and to systematically study their properties—including anomalies and onsitability—directly at the microscopic level [83, 84, 85, 86]. Such constructions would provide a concrete bridge between continuum effective field theories and lattice Hamiltonians.
Another natural direction is to investigate how generalized symmetries in a three-dimensional topologically ordered phase constrain its boundary theories. Within the SymTFT framework, bulk generalized symmetries are expected to impose strong restrictions on admissible boundary degrees of freedom, anomalies, and symmetry realizations. A systematic analysis of boundary theories from this perspective may lead to a unified understanding of bulk–boundary correspondence in higher dimensions. In addition, it is intriguing to explore whether the $1$-form and $2$-form symmetries identified in this work organize into a nontrivial higher-group structure [87, 88, 89]. Clarifying such higher-group structures, both in continuum field theories and in lattice realizations, could further illuminate the internal symmetry organization of three-dimensional topological orders.
Beyond symmetry structure, a promising set of directions concerns the interplay between generalized symmetries and quantum-information properties. In particular, it would be interesting to study how generalized symmetries constrain the entanglement structure of quantum states. For example, one may investigate the circuit complexity required to prepare quantum states that differ by their generalized symmetry content, potentially revealing intrinsic notions of complexity protected by higher-form or non-invertible symmetries. Furthermore, extending the theory of anyon condensation to (3+1)(3+1)D systems remains an important open problem. Such a generalization is expected to be closely related to phase transitions between (3+1)(3+1)D topological phases characterized by distinct generalized symmetry structures.
Finally, it is important to explore how generalized symmetries constrain the behavior of quantum systems at finite temperature or under decoherence, where the system is in a mixed state rather than a pure state. Recent work has shown that three-dimensional systems below a nonzero temperature can still exhibit topological order due to anomalous $2$-form symmetries associated with emergent fermionic excitations [90]. Since $BF$ theories with a $b\wedge b$ twist describe a class of three-dimensional topological orders with fermionic excitations, they provide a natural arena to further investigate the role of anomalous generalized symmetries in realizing mixed-state topological order. Understanding these phenomena within a unified continuum field-theoretical framework may shed new light on the robustness of topological phases beyond the zero-temperature, closed-system setting.
Acknowledgements.
This work was in part supported by National Natural Science Foundation of China (NSFC) under Grants No. 12474149 and No. 12274250, Research Center for Magnetoelectric Physics of Guangdong Province under Grant No. 2024B0303390001, and Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices under Grant No. 2022B1212010008.
Appendix A Generalized symmetries in lattice models
In Sec. II we explain the concepts of generalized symmetries in the language of continuum field theory. As a comparison, in this appendix, we use two representative lattice model examples to illustrate the higher-form symmetry (the toric code models in 22d and 33d) and the non-invertible symmetry (the 11d critical Ising model).
A.1 Higher-form symmetries in 22d and 33d toric code models
The Hamiltonian of the toric code model on a square lattice is given by
$H_{\text{TC}}=-\sum_{v}A_{v}-\sum_{p}B_{p}=-\sum_{v}\prod_{\partial l\ni v}X_{l}-\sum_{p}\prod_{l\in\partial p}Z_{l},$ (149)
where qubits reside on the links of the lattice. The ground state satisfies $A_{v}=B_{p}=1$ for all vertices $v$ and plaquettes $p$. An excitation with $A_{v}=-1$ ($B_{p}=-1$) is referred to as an $e$ excitation (an $m$ excitation).
There exist two operators that commute with $H_{\text{TC}}$,
$W_{e}\left(L\right)=\prod_{l\in L}Z_{l},\qquad W_{m}\left(L^{*}\right)=\prod_{l\perp L^{*}}X_{l},$ (150)
where $L$ and $L^{*}$ are closed loops on the original lattice and its dual, respectively. Each operator generates a $\mathbb{Z}_{2}$ 11-form symmetry, denoted as $\mathbb{Z}_{2,e}^{(1)}$ and $\mathbb{Z}_{2,m}^{(1)}$, respectively.
We can identify operators charged under the $\mathbb{Z}_{2,e}^{(1)}$ and $\mathbb{Z}_{2,m}^{(1)}$ symmetry transformations,
$O_{x}=\prod_{l\perp c^{*}}X_{l},\qquad O_{z}=\prod_{l\in c}Z_{l},$ (151)
where $c^{*}$ ($c$) is a line on the dual (original) lattice. The lines $c^{*}$ and $c$ need not be closed. The actions of the symmetry operators on these charged operators are
$W_{e}\left(L\right)O_{x}W_{e}^{\dagger}\left(L\right)=\left(-1\right)^{{\rm Int}\left(L,c^{*}\right)}O_{x},$ (152)
$W_{m}\left(L^{*}\right)O_{z}W_{m}^{\dagger}\left(L^{*}\right)=\left(-1\right)^{{\rm Int}\left(L^{*},c\right)}O_{z},$ (153)
where ${\rm Int}\left(L,c^{*}\right)$ denotes the intersection number of $L$ and $c^{*}$. When $c^{*}$ ($c$) is an open line, the charged operator creates a pair of $e$ ($m$) excitations. Accordingly, applying $W_{e}\left(L\right)$ or $W_{m}\left(L^{*}\right)$ measures the number (modulo $2$) of $e$ or $m$ excitations enclosed by the area bounded by $L$ or $L^{*}$.
In addition, the two symmetry operators satisfy
$W_{e}\left(L\right)W_{m}\left(L^{*}\right)=\left(-1\right)^{{\rm Int}\left(L,L^{*}\right)}W_{m}\left(L^{*}\right)W_{e}\left(L\right).$ (154)
This relation implies that $W_{e}\left(L\right)$ is charged under the symmetry generated by $W_{m}\left(L^{*}\right)$, and vice versa.
The higher-form symmetries of the toric code model can also be described using its effective field theory,
$S=\int_{M_{3}}\frac{2}{2\pi}a^{1}da^{2},$ (155)
where $a^{1}$ and $a^{2}$ are $1$-form gauge fields. The equations of motion are $\frac{1}{\pi}da^{1}=0$ and $\frac{1}{\pi}da^{2}=0$. Interpreting these equations as continuity equations, we introduce the conserved quantities
$Q_{1}=\int_{\gamma}a^{1},\qquad Q_{2}=\int_{\gamma}a^{2},$ (156)
which measure the number of topological excitations enclosed by the area bounded by the closed curve $\gamma$. These conserved quantities generate the symmetry operators
$U_{e}\left(\gamma\right)=\exp\left({\rm i}\int_{\gamma}a^{2}\right),\qquad U_{m}\left(\gamma\right)=\exp\left({\rm i}\int_{\gamma}a^{1}\right),$ (157)
which satisfy
$U_{e}\left(\gamma_{1}\right)U_{m}\left(\gamma_{2}\right)=\left(-1\right)^{{\rm Lk}\left(\gamma_{1},\gamma_{2}\right)}U_{m}\left(\gamma_{2}\right)U_{e}\left(\gamma_{1}\right).$ (158)
These operators realize a $\mathbb{Z}_{2,e}^{(1)}\times\mathbb{Z}_{2,m}^{(1)}$ 11-form symmetry, consistent with the lattice description.
We now turn to the 33d toric code model to illustrate $1$-form and $2$-form symmetries. Consider a three-dimensional cubic lattice with qubits on the links. The Hamiltonian is
$H_{\text{3dTC}}=-\sum_{v}A_{v}-\sum_{p}B_{p}=-\sum_{v}\prod_{l\ni v}X_{l}-\sum_{p}\prod_{l\in\partial p}Z_{l}.$ (159)
A particle excitation at vertex $v$ corresponds to $A_{v}=-1$. A pair of particle excitations can be created at the endpoints of an open string $s$ by applying $W_{e}\left(s\right)=\prod_{l\in s}Z_{l}$ to the ground state. A loop excitation is created by applying a surface operator $W_{m}\left(A^{*}\right)=\prod_{l\perp A^{*}}X_{l}$, where $A^{*}$ is an open surface on the dual lattice.
Each $X_{l}$ operator affects two plaquette terms sharing the link $l$. Along the boundary $\partial A^{*}$ of the surface, the plaquettes acquire an odd number of $X_{l}$ operators and thus satisfy $B_{p}=-1$. Inside the surface, the plaquettes are flipped twice and remain $B_{p}=+1$. In this way, a loop excitation is created along the boundary of $A^{*}$.
Analogous to the 22d case, the 33d toric code model hosts two higher-form symmetries. One is generated by
$W_{e}\left(L\right)=\prod_{l\in L}Z_{l},$ (160)
where $L$ is a closed curve on the cubic lattice. This operator acts on a codimension-$2$ submanifold and generates a $2$-form symmetry. Its effect can be interpreted as moving a particle excitation along $L$. The other symmetry is generated by
$W_{m}\left(M^{*}\right)=\prod_{l\perp M^{*}}X_{l},$ (161)
where $M^{*}$ is a closed surface on the dual lattice. This operator acts on a codimension-$1$ submanifold and generates a $1$-form symmetry, corresponding to moving a loop excitation along $M^{*}$. The two symmetry operators satisfy
$W_{e}\left(L\right)W_{m}\left(M^{*}\right)=\left(-1\right)^{{\rm Int}\left(L,M^{*}\right)}W_{m}\left(M^{*}\right)W_{e}\left(L\right),$ (162)
where ${\rm Int}\left(L,M^{*}\right)$ denotes the intersection number.
From the effective field theory perspective, the 33d toric code is described by the (3+1)(3+1)D $\mathbb{Z}_{2}$ BFBF theory,
$S=\int_{M_{4}}\frac{2}{2\pi}bda,$ (163)
where $b$ and $a$ are $2$- and $1$-form gauge fields. The equations of motion $\frac{1}{\pi}da=0$ and $\frac{1}{\pi}db=0$ imply two conserved quantities,
$Q_{1}=\int_{\gamma}a,\qquad Q_{2}=\int_{\sigma}b,$ (164)
with closed curve $\gamma$ and closed surface $\sigma$. The corresponding symmetry operators are
$U_{e}\left(\gamma\right)=\exp\left({\rm i}\int_{\gamma}a\right),\qquad U_{m}\left(\sigma\right)=\exp\left({\rm i}\int_{\sigma}b\right),$ (165)
which satisfy
$U_{e}\left(\gamma\right)U_{m}\left(\sigma\right)=\left(-1\right)^{{\rm Lk}\left(\gamma,\sigma\right)}U_{m}\left(\sigma\right)U_{e}\left(\gamma\right).$ (166)
These operators generate a $\mathbb{Z}_{2}$ 22-form and a $\mathbb{Z}_{2}$ 11-form symmetry, respectively.
An equivalent description of the 33d toric code places qubits on the plaquettes of the cubic lattice. The Hamiltonian is
$H_{\text{3dTC}}^{\prime}=-\sum_{c}A_{c}-\sum_{l}B_{l}=-\sum_{c}\prod_{p\in\partial c}X_{p}-\sum_{l}\prod_{p\ni l}Z_{p}.$ (167)
The two descriptions are related by the identity
$S=\int_{M_{4}}\frac{2}{2\pi}bda=\int_{M_{4}}\frac{2}{2\pi}adb+\text{(total derivative)}.$ (168)
In this field-theoretical description, the higher-form symmetries are encoded in the equations $da=0$ and $db=0$, which can be interpreted as continuity equations. The coefficient $\frac{1}{\pi}$ reflects the underlying $\mathbb{Z}_{2}$ structure.
A.2 Non-invertible symmetry in (1+1)(1+1)D transverse-field Ising model
In Appendix A.1, we reviewed generalized symmetries whose operators form a group structure and therefore admit inverses. A complementary generalization of symmetry relaxes the requirement of invertibility. Symmetry operators of this type do not possess inverses and are referred to as non-invertible symmetries.
The (1+1)(1+1)D transverse-field Ising model at its critical point provides a canonical example of a non-invertible symmetry [16, 91, 23]. The Hamiltonian is
$H=-\sum_{i=1}^{N}Z_{i}Z_{i+1}-g\sum_{i=1}^{N}X_{i},$ (169)
with periodic boundary conditions $X_{N+1}=X_{1}$ and $Z_{N+1}=Z_{1}$. This model has a $\mathbb{Z}_{2}$ global symmetry generated by $\mathsf{P}=\prod_{i=1}^{N}X_{i}$, which satisfies $\left[\mathsf{P},H\right]=0$ and $\mathsf{P}^{2}=\mathbb{I}$.
At the critical point $g=1$, the Hamiltonian is invariant under the Kramers–Wannier transformation,
$X_{i}\rightarrow Z_{i}Z_{i+1},\qquad Z_{i}Z_{i+1}\rightarrow X_{i+1},\qquad i=1,\ldots,N.$ (170)
Although one might attempt to realize this transformation using a unitary operator, such an operator does not exist.
If such a unitary operator $U$ existed, satisfying $UHU^{-1}=H$ and $UX_{i}U^{-1}=Z_{i}Z_{i+1}$, then it would imply
$U\mathsf{P}U^{-1}=U\left(\prod_{i=1}^{N}X_{i}\right)U^{-1}=\prod_{i=1}^{N}Z_{i}Z_{i+1}=1,$ (171)
which would force $\mathsf{P}=1$ and lead to a contradiction. Therefore, the Kramers–Wannier transformation cannot be implemented by a unitary operator.
Instead, the Kramers–Wannier transformation is realized by a non-invertible operator $\mathsf{D}$ acting on the Hilbert space,
$\mathsf{D}=U_{\rm KW}\frac{1+\mathsf{P}}{2},$ (172)
where
$U_{\rm KW}=e^{-\frac{{\rm i}2\pi N}{8}}\left(\prod_{i=1}^{N}\frac{1+{\rm i}X_{i}}{\sqrt{2}}\frac{1+{\rm i}Z_{i}Z_{i+1}}{\sqrt{2}}\right)\frac{1+{\rm i}X_{N}}{\sqrt{2}}$ (173)
is a unitary operator, and $(1+\mathsf{P})/2$ is a $\mathbb{Z}_{2}$ projector that projects the Hilbert space onto the $\mathsf{P}=+1$ sector. The operator $\mathsf{D}$ acts on local operators as
$\mathsf{D}X_{i}=Z_{i}Z_{i+1}\mathsf{D},\qquad\mathsf{D}Z_{i}Z_{i+1}=X_{i+1}\mathsf{D},\qquad i=1,\cdots,N,$ (174)
with $X_{N+1}=X_{1}$ and $Z_{N+1}=Z_{1}$, which reproduces the Kramers–Wannier transformation.
For comparison, the action of $U_{\rm KW}$ is $U_{\rm KW}X_{i}U_{\rm KW}^{-1}=Z_{i}Z_{i+1}$ and $U_{\rm KW}Z_{i}Z_{i+1}U_{\rm KW}^{-1}=X_{i+1}$ for $i=1,\cdots,N-1$, while for the boundary terms one finds $U_{\rm KW}X_{N}U_{\rm KW}^{-1}=\mathsf{P}Z_{N}Z_{1}$ and $U_{\rm KW}Z_{N}Z_{1}U_{\rm KW}^{-1}=\mathsf{P}X_{1}$. Hence $U_{\rm KW}$ does not commute with the Hamiltonian and is not a symmetry. By contrast, $\mathsf{D}$ commutes with $H$ at $g=1$,
$\mathsf{D}H=H\mathsf{D},\qquad\text{or}\qquad\left[\mathsf{D},H\right]=0,\qquad\text{when }g=1.$ (175)
On the other hand, because $\mathsf{D}$ contains a projector, its action annihilates all $\mathbb{Z}_{2}$-odd states. Therefore, there is no operator $\mathsf{O}$ such that $\mathsf{O}\mathsf{D}\left|\psi\right\rangle=\left|\psi\right\rangle$. In this sense, $\mathsf{D}$ is non-invertible, and the associated symmetry is referred to as a non-invertible symmetry.
The key ingredient underlying this non-invertible symmetry is the invariance of the critical theory under the Kramers–Wannier transformation, i.e., under gauging. It has also been found that certain (3+1)(3+1)D theories exhibit a related invariance under gauging: one may gauge only part of the system and thereby create a topological defect interface separating the original and gauged regions. Such topological defects are non-invertible [92, 93, 94]. In the present example, the symmetry operator $\mathsf{D}$ is a product of a unitary operator $U_{\rm KW}$ and a projector, where the latter is responsible for the non-invertibility.
Appendix B An example of fusion rule in the setup of $G=\left(\mathbb{Z}_{6}\right)^{3}$ and $p=2$
This example is a supplementary for the discussion of the fusion of symmetry operators in the main text. In Sec. IV.4, we have discussed the examples of $p=1$ and $p=2$ when $G=\left(\mathbb{Z}_{3}\right)^{3}$. Here, we want to provide an example such that $p$ and $N_{123}$ are not coprime. Although we still work with three cyclic factors $\mathbb{Z}_{N_{i}}$ (here, $N_{i}=6$), we choose the level $p=2$ so that $\gcd\!\left(p,N_{123}\right)>1$. (If we instead chose $p=1$, then $\gcd\!\left(p,N_{123}\right)=1$, and the analysis would be analogous to the case $N_{i}=3$ and $p=1$.)
We consider the fusion $L_{123}\times L_{234}$ to illustrate how to compute the fusion rule for two generic symmetry operators. The fusion channels include $L_{351}$ and $L_{351}U_{abc}$ because $\mathcal{U}_{123}\times\mathcal{U}_{234}=\mathcal{U}_{351}$. The possible invertible symmetry factors $U_{abc}$ are determined by analyzing the projectors (delta functions),
$L_{123}$ $=\mathcal{U}_{123}\delta\left(p\int_{\gamma_{p}}a^{2}-2p\int_{\gamma_{p}}a^{1}\right)\delta\left(p\int_{\sigma}b^{3}-3p\int_{\gamma_{p}}a^{1}\right)$
$\times\delta\left(2p\int_{\sigma}b^{3}-3p\int_{\gamma_{p}}a^{2}\right),$ (176)
$L_{234}$ $=\mathcal{U}_{234}\delta\left(2p\int_{\gamma_{p}}a^{2}-3p\int_{\gamma_{p}}a^{1}\right)\delta\left(2p\int_{\sigma}b^{3}-4p\int_{\gamma_{p}}a^{1}\right)$
$\times\delta\left(3p\int_{\sigma}b^{3}-4p\int_{\gamma_{p}}a^{2}\right).$ (177)
We first expand $\mathbb{P}_{123}$ as a linear combination of invertible symmetry operators. Let $\int_{\gamma_{p}}a^{1}=\frac{2\pi k_{1}}{N}$, $\int_{\gamma_{p}}a^{2}=\frac{2\pi k_{2}}{N}$, and $\int_{\sigma}b^{3}=\frac{2\pi k_{3}}{N}$; in this example, $N=6$. Then $p\int_{\gamma_{p}}a^{2}-2p\int_{\gamma_{p}}a^{1}=\frac{2\pi p}{N}\left(k_{2}-2k_{1}\right)$ with $k_{1},k_{2}\in\left\{1,2,\cdots,6\right\}$. Hence we may write
$\delta\left(p\int_{\gamma_{p}}a^{2}-2p\int_{\gamma_{p}}a^{1}\right)=\frac{1}{N}\sum_{l_{1}=0}^{N-1}e^{{\rm i}l_{1}\cdot\left(p\int_{\gamma_{p}}a^{2}-2p\int_{\gamma_{p}}a^{1}\right)}.$
Similarly,
$\delta\left(p\int_{\sigma}b^{3}-3p\int_{\gamma_{p}}a^{1}\right)=\frac{1}{N}\sum_{l_{2}=0}^{N-1}e^{{\rm i}l_{2}\cdot\left(p\int_{\sigma}b^{3}-3p\int_{\gamma_{p}}a^{1}\right)}.$
However, since $p=2$, the factor $e^{{\rm i}3p\int_{\gamma_{p}}a^{1}}$ is in fact the identity. Therefore the above delta function simplifies to
$\delta\left(p\int_{\sigma}b^{3}-3p\int_{\gamma_{p}}a^{1}\right)=\frac{1}{N}\sum_{l_{2}=0}^{N-1}e^{{\rm i}l_{2}\cdot\left(p\int_{\sigma}b^{3}\right)}.$
The third delta function can be expanded as
$\delta\left(2p\int_{\sigma}b^{3}-3p\int_{\gamma_{p}}a^{2}\right)=\frac{1}{N}\sum_{l_{3}=0}^{N-1}e^{{\rm i}l_{3}\cdot\left(2p\int_{\sigma}b^{3}\right)}.$
Note that $\delta\left(p\int_{\sigma}b^{3}\right)$ and $\delta\left(2p\int_{\sigma}b^{3}\right)$ are the same in this setting: both impose $\int_{\sigma}b^{3}\in\frac{2\pi}{6}\cdot 3\mathbb{Z}$, so that $p\int_{\sigma}b^{3}=2\int_{\sigma}b^{3}=0\mod 2\pi$ and also $2p\int_{\sigma}b^{3}=0\mod 2\pi$. In short,
$\mathbb{P}_{123}=$ $\delta\left(p\int_{\gamma_{p}}a^{2}-2p\int_{\gamma_{p}}a^{1}\right)\delta\left(p\int_{\sigma}b^{3}\right)=\frac{1}{N}\sum_{l_{1}=0}^{N-1}e^{{\rm i}l_{1}\cdot\left(p\int_{\gamma_{p}}a^{2}-2p\int_{\gamma_{p}}a^{1}\right)}\times\frac{1}{N}\sum_{l_{2}=0}^{N-1}e^{{\rm i}l_{2}\cdot\left(p\int_{\sigma}b^{3}\right)}.$ (178)
We next expand $\mathbb{P}_{234}$ as a linear combination of invertible symmetry operators. The relevant delta functions are as follows. We have
$\delta\left(2p\int_{\gamma_{p}}a^{2}-3p\int_{\gamma_{p}}a^{1}\right)=\frac{1}{N}\sum_{l_{3}=0}^{N-1}e^{{\rm i}l_{3}\cdot\left(2p\int_{\gamma_{p}}a^{2}\right)},$
because $3p\int_{\gamma_{p}}a^{1}\in 2\pi\mathbb{Z}$. Also,
$\delta\left(2p\int_{\sigma}b^{3}-4p\int_{\gamma_{p}}a^{1}\right)=\frac{1}{N}\sum_{l_{4}=0}^{N-1}e^{{\rm i}l_{4}\cdot\left(2p\int_{\sigma}b^{3}-4p\int_{\gamma_{p}}a^{1}\right)}.$
Finally,
$\delta\left(3p\int_{\sigma}b^{3}-4p\int_{\gamma_{p}}a^{2}\right)=\frac{1}{N}\sum_{l_{5}=0}^{N-1}e^{{\rm i}l_{5}\cdot\left(4p\int_{\gamma_{p}}a^{2}\right)},$
because $3p\int_{\sigma}b^{3}\in 2\pi\mathbb{Z}$. Moreover, $\delta\left(2p\int_{\gamma_{p}}a^{2}\right)$ and $\delta\left(4p\int_{\gamma_{p}}a^{2}\right)$ are the same. In short,
$\mathbb{P}_{234}=$ $\frac{1}{N}\sum_{l_{4}=0}^{N-1}e^{{\rm i}l_{4}\cdot\left(2p\int_{\sigma}b^{3}-4p\int_{\gamma_{p}}a^{1}\right)}\times\frac{1}{N}\sum_{l_{3}=0}^{N-1}e^{{\rm i}l_{3}\cdot\left(2p\int_{\gamma_{p}}a^{2}\right)}.$ (179)
The symmetry operator $L_{351}$ appearing in the fusion channels is
$L_{351}$ $=\mathcal{U}_{351}\delta\left(3p\int_{\gamma_{p}}a^{2}-5p\int_{\gamma_{p}}a^{1}\right)\delta\left(3p\int_{\sigma}b^{3}-p\int_{\gamma_{p}}a^{1}\right)$
$\times\delta\left(5p\int_{\sigma}b^{3}-p\int_{\gamma_{p}}a^{2}\right).$ (180)
The delta functions can be expanded as follows. We have
$\delta\left(3p\int_{\gamma_{p}}a^{2}-5p\int_{\gamma_{p}}a^{1}\right)=\frac{1}{N}\sum_{h_{1}=0}^{N-1}e^{{\rm i}h_{1}\cdot\left(5p\int_{\gamma_{p}}a^{1}\right)},$
since $3p\int_{\gamma_{p}}a^{2}\in 2\pi\mathbb{Z}$. Likewise,
$\delta\left(3p\int_{\sigma}b^{3}-p\int_{\gamma_{p}}a^{1}\right)=\frac{1}{N}\sum_{h_{2}=0}^{N-1}e^{{\rm i}h_{2}\cdot\left(p\int_{\gamma_{p}}a^{1}\right)},$
since $3p\int_{\sigma}b^{3}\in 2\pi\mathbb{Z}$. Moreover, in the present case $p=2$ and $N=6$, we have $\delta\left(5p\int_{\gamma_{p}}a^{1}\right)=\delta\left(p\int_{\gamma_{p}}a^{1}\right)$. Finally,
$\delta\left(5p\int_{\sigma}b^{3}-p\int_{\gamma_{p}}a^{2}\right)=\frac{1}{N}\sum_{h_{3}=0}^{N-1}e^{{\rm i}h_{3}\cdot\left(5p\int_{\sigma}b^{3}-p\int_{\gamma_{p}}a^{2}\right)}.$
In summary, the projector in $L_{351}$ is
$\mathbb{P}_{351}=$ $\frac{1}{N}\sum_{h_{3}=0}^{N-1}e^{{\rm i}h_{3}\cdot\left(5p\int_{\sigma}b^{3}-p\int_{\gamma_{p}}a^{2}\right)}\times\frac{1}{N}\sum_{h_{2}=0}^{N-1}e^{{\rm i}h_{2}\cdot\left(p\int_{\gamma_{p}}a^{1}\right)}.$ (181)
We now compute $\mathbb{P}_{123}\times\mathbb{P}_{234}$. A direct calculation yields
$\mathbb{P}_{123}\times\mathbb{P}_{234}=$ $\left(\frac{1}{6}\right)^{4}\sum_{l_{1},l_{2},l_{3},l_{4}=0}^{5}e^{{\rm i}\left(2l_{1}+4l_{4}\right)\int_{\gamma_{p}}a^{1}+{\rm i}\left(4l_{3}+2l_{1}\right)\int_{\gamma_{p}}a^{2}+{\rm i}\left(2l_{2}+4l_{4}\right)\int_{\sigma}b^{3}},$ (182)
where we have substituted $p=2$. Since $l_{1},l_{2},l_{3},l_{4}\in\left\{0,1,\cdots,5\right\}$, the combinations $\left(2l_{1}+4l_{4}\right)$, $\left(4l_{3}+2l_{1}\right)$, and $\left(2l_{2}+4l_{4}\right)$ each take values in $\left\{0,2,4\right\}$. Therefore only $3^{3}=27$ distinct phase factors appear in the sum, and each such factor occurs $\frac{6^{4}}{27}=48$ times. We can thus rewrite
$\mathbb{P}_{123}\times\mathbb{P}_{234}=$ $\frac{48}{6^{4}}\sum_{\left\{m_{1},m_{2},m_{3}\right\}}e^{{\rm i}m_{1}\int_{\gamma_{p}}a^{1}+{\rm i}m_{2}\int_{\gamma_{p}}a^{2}+{\rm i}m_{3}\int_{\sigma}b^{3}}$ (183)
where $m_{1},m_{2},m_{3}\in\left\{0,2,4\right\}$ and there are $27$ different triples $\left\{m_{1},m_{2},m_{3}\right\}$.
If we instead let $l_{i}$ range from 0 to $5$, we find
$8\times\sum_{\left\{m_{1},m_{2},m_{3}\right\}}e^{{\rm i}m_{1}\int_{\gamma_{p}}a^{1}+{\rm i}m_{2}\int_{\gamma_{p}}a^{2}+{\rm i}m_{3}\int_{\sigma}b^{3}}=\sum_{l_{1},l_{2},l_{3}=0}^{5}e^{{\rm i}2l_{1}\int_{\gamma_{p}}a^{1}+{\rm i}2l_{2}\int_{\gamma_{p}}a^{2}+{\rm i}2l_{3}\int_{\sigma}b^{3}},$ (184)
where we identify $\left\{m_{1},m_{2},m_{3}\right\}$ with $\left\{2l_{1},2l_{2},2l_{3}\right\}$. The left-hand side contains $8\times 27=216$ terms, and the right-hand side contains $6^{3}=216$ terms as well. Therefore,
$\mathbb{P}_{123}\times\mathbb{P}_{234}=\frac{1}{6^{3}}\sum_{l_{1},l_{2},l_{3}=0}^{5}e^{{\rm i}2l_{1}\int_{\gamma_{p}}a^{1}+{\rm i}2l_{2}\int_{\gamma_{p}}a^{2}+{\rm i}2l_{3}\int_{\sigma}b^{3}}.$ (185)
On the other hand, the projector $\mathbb{P}_{351}$ is computed as
$\mathbb{P}_{351}=$ $\left(\frac{1}{6}\right)^{2}\sum_{h_{2},h_{3}=0}^{5}e^{{\rm i}2h_{2}\int_{\gamma_{p}}a^{1}+{\rm i}4h_{3}\cdot\left(\int_{\sigma}b^{3}+\int_{\gamma_{p}}a^{2}\right)}$ (186)
where we have substituted $p=2$ and used $e^{{\rm i}6\int_{\sigma}b^{3}}=e^{{\rm i}6\int_{\gamma_{p}}a^{2}}=1$.
To obtain the fusion rule $L_{123}\times L_{234}$, we compare $\mathbb{P}_{123}\times\mathbb{P}_{234}$ with $\mathbb{P}_{351}$ to determine how they match. The factor $\sum_{l_{1}=0}^{5}e^{{\rm i}2l_{1}\int_{\gamma_{p}}a^{1}}$ appears in both expressions. Next, the contribution $e^{{\rm i}2l_{2}\int_{\gamma_{p}}a^{2}+{\rm i}2l_{3}\int_{\sigma}b^{3}}$ in $\mathbb{P}_{123}\times\mathbb{P}_{234}$ can be simplified as
$\underbrace{\frac{1}{6}\sum_{l_{2}=0}^{5}e^{{\rm i}l_{2}\cdot 2\int_{\gamma_{p}}a^{2}}\times\frac{1}{6}\sum_{l_{3}=0}^{5}e^{{\rm i}l_{3}\cdot 2\int_{\sigma}b^{3}}}_{\text{from }\mathbb{P}_{123}\times\mathbb{P}_{234}}=\frac{1}{3}\sum_{l_{2}=0}^{2}e^{{\rm i}l_{2}\cdot 2\int_{\gamma_{p}}a^{2}}\times\frac{1}{3}\sum_{l_{3}=0}^{2}e^{{\rm i}l_{3}\cdot 2\int_{\sigma}b^{3}}.$ (187)
Moreover, the factor $\sum_{h_{3}=0}^{5}e^{{\rm i}4h_{3}\cdot\left(\int_{\sigma}b^{3}+\int_{\gamma_{p}}a^{2}\right)}$ in $\mathbb{P}_{351}$, when multiplied by $\frac{1}{3}\sum_{m_{2}=0}^{2}e^{{\rm i}m_{2}\cdot 2\int_{\gamma_{p}}a^{2}}$, reproduces the term $\sum_{l_{2},l_{3}=0}^{5}e^{{\rm i}2l_{2}\int_{\gamma_{p}}a^{2}+{\rm i}2l_{3}\int_{\sigma}b^{3}}$ in $\mathbb{P}_{123}\times\mathbb{P}_{234}$:
$\underbrace{\frac{1}{6}\sum_{h_{3}=0}^{5}e^{{\rm i}4h_{3}\cdot\left(\int_{\sigma}b^{3}+\int_{\gamma_{p}}a^{2}\right)}}_{\text{from }\mathbb{P}_{351}}\times\frac{1}{3}\sum_{m_{2}=0}^{2}e^{{\rm i}m_{2}\cdot 2\int_{\gamma_{p}}a^{2}}=$ $\frac{1}{3}\sum_{h_{3}=0}^{2}e^{{\rm i}h_{3}\cdot\left(2\int_{\sigma}b^{3}+2\int_{\gamma_{p}}a^{2}\right)}\times\frac{1}{3}\sum_{m_{2}=0}^{2}e^{{\rm i}m_{2}\cdot 2\int_{\gamma_{p}}a^{2}}$
$=$ $\frac{1}{3}\sum_{h_{3}=0}^{2}e^{{\rm i}h_{3}\cdot\left(2\int_{\sigma}b^{3}\right)}\times\frac{1}{3}\sum_{m_{2}=0}^{2}e^{{\rm i}m_{2}\cdot 2\int_{\gamma_{p}}a^{2}}$
$=$ $\underbrace{\frac{1}{6}\sum_{l_{2}=0}^{5}e^{{\rm i}l_{2}\cdot 2\int_{\gamma_{p}}a^{2}}\times\frac{1}{6}\sum_{l_{3}=0}^{5}e^{{\rm i}l_{3}\cdot 2\int_{\sigma}b^{3}}}_{\text{from }\mathbb{P}_{123}\times\mathbb{P}_{234}}.$ (188)
Therefore,
$\mathbb{P}_{123}\times\mathbb{P}_{234}=$ $\frac{1}{3}\sum_{m_{2}=0}^{2}e^{{\rm i}m_{2}\cdot 2\int_{\gamma_{p}}a^{2}}\mathbb{P}_{351}$ (189)
and the fusion rule is
$L_{123}\times L_{234}=\frac{1}{3}\left(L_{351}+U_{020}L_{351}+U_{040}L_{351}\right).$ (190)
Appendix C Explanation of no higher-dimensional bulk to cancel mixed anomaly of two non-invertible symmetries
At the end of Sec. V.2, we have mentioned that a non-resolvable mixed anomaly exists when we try to gauge two non-invertible symmetries simultaneously, manifesting as that the non-gauge-invariant $S_{\text{gauging}}^{m_{1},m_{2}}$, see Eq. (148), cannot recover gauge invariance by coupling to a higher-dimensional bulk term. Below we explain why we cannot find a bulk term $S_{\text{bulk}}$ such that $S_{\text{gauging}}^{m_{1},m_{2}}+S_{\text{bulk}}$ is gauge invariant under $\begin{cases}\mathcal{B}_{m}^{1}\rightarrow&\mathcal{B}_{m}^{1}+dV_{m}^{1}\\ a^{1}\rightarrow&a^{1}-\frac{1}{N_{1}}V_{m}^{1}\end{cases}$ and $\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases}$.
The gauge transformation of $a^{1}$ is $a^{1}\rightarrow a^{1}-\frac{1}{N_{1}}V_{m}^{1}$. Since $d\phi^{1}=\frac{pN_{1}}{N_{123}}a^{1}$, we suppose $d\phi^{1}\rightarrow d\phi^{1}-\frac{1}{N_{123}}V_{m}^{1}$. Accordingly, we write $\phi^{1}\rightarrow\phi^{1}-\eta^{1}$, where $d\eta^{1}=\frac{1}{N_{123}}V_{m}^{1}$. However, this would imply $d\left(d\eta^{1}\right)=\frac{p}{N_{123}}dV_{m}^{1}=0$, i.e., $dV_{m}^{1}=0$. But $dV_{m}^{1}=0$ would make the gauge transformation of $\mathcal{B}_{m}^{1}$ trivial. This indicates that, when we gauge the symmetry generated by $J^{1}_{m}$, the current $J_{m}^{2}=*\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right)$ does not admit a consistent gauge transformation. This incompatibility of gauge invariance already signals a mixed anomaly.
Below we assume that $\phi^{1}\rightarrow\phi_{new}^{1}$ and $\phi^{2}\rightarrow\phi_{new}^{2}$ such that $d\phi_{new}^{1}=\frac{pN_{1}}{N_{123}}a^{1}-\frac{p}{N_{123}}V_{m}^{1}$ and $d\phi_{new}^{2}=\frac{pN_{2}}{N_{123}}a^{2}-\frac{p}{N_{123}}V_{m}^{2}$. We will show that $S_{\text{gauging}}^{m_{1},m_{2}}+S_{\text{bulk}}$ is not gauge invariant for any choice of $S_{\text{bulk}}$. We denote
$S_{\text{gauging}}^{m_{1},m_{2}}=S_{aab}+S_{\text{coupling}}.$ (191)
Under $\begin{cases}\mathcal{B}_{m}^{1}\rightarrow&\mathcal{B}_{m}^{1}+dV_{m}^{1}\\ a^{1}\rightarrow&a^{1}-\frac{1}{N_{1}}V_{m}^{1}\end{cases}$ and $\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases}$, the variation of $S_{aab}$ is
$\Delta S_{aab}=$ $\int_{M_{4}}-\frac{1}{2\pi}b^{1}dV_{m}^{1}-\frac{1}{2\pi}b^{2}dV_{m}^{2}+q\left(a^{1}-\frac{1}{N_{1}}V_{m}^{1}\right)\left(a^{2}-\frac{1}{N_{2}}V_{m}^{2}\right)b^{3}-qa^{1}a^{2}b^{3}$
$=$ $\int_{M_{4}}-\frac{1}{2\pi}b^{1}dV_{m}^{1}-\frac{1}{2\pi}b^{2}dV_{m}^{2}-\frac{q}{N_{1}}V_{m}^{1}a^{2}b^{3}-\frac{q}{N_{2}}a^{1}V_{m}^{2}b^{3}+\frac{q}{N_{1}N_{2}}V_{m}^{1}V_{m}^{2}b^{3}$ (192)
The coupling term $S_{\text{coupling}}$ becomes
$S_{\text{coupling}}\rightarrow$ $\int_{M_{4}}\frac{1}{2\pi}\left(\mathcal{B}_{m}^{1}+dV_{m}^{1}\right)\wedge\left(b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}\left(a^{2}-\frac{1}{N_{2}}V_{M}^{2}\right)\beta^{3}+\frac{1}{2}\phi_{new}^{2}\frac{N_{3}}{2\pi}b^{3}\right)$
$+\int_{M_{4}}\frac{1}{2\pi}\left(\mathcal{B}_{m}^{2}+dV_{m}^{2}\right)\wedge\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}\left(a^{1}-\frac{1}{N_{1}}V_{m}^{1}\right)\beta^{3}+\frac{1}{2}\phi_{new}^{1}\frac{N_{3}}{2\pi}b^{3}\right)$
$=$ $\int_{M_{4}}\frac{1}{2\pi}\left(\mathcal{B}_{m}^{1}+dV_{m}^{1}\right)\wedge\left(b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}+\frac{1}{2}\frac{1}{2\pi}V_{m}^{2}\beta^{3}+\frac{1}{2}\phi_{new}^{2}\frac{N_{3}}{2\pi}b^{3}\right)$
$+\int_{M_{4}}\frac{1}{2\pi}\left(\mathcal{B}_{m}^{2}+dV_{m}^{2}\right)\wedge\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\frac{1}{2\pi}V_{m}^{1}\beta^{3}+\frac{1}{2}\phi_{new}^{1}\frac{N_{3}}{2\pi}b^{3}\right)$ (193)
and the variation of $S_{\text{coupling}}$ is
$\Delta S_{\text{coupling}}=$ $\int_{M_{4}}\frac{1}{2\pi}dV_{m}^{1}\wedge b^{1}-\frac{1}{2}\frac{N_{2}}{4\pi^{2}}dV_{m}^{1}a^{2}\beta^{3}+\frac{1}{2}\frac{1}{4\pi^{2}}dV_{m}^{1}V_{m}^{2}\beta^{3}+\frac{1}{2}\frac{1}{2\pi}dV_{m}^{1}\phi_{new}^{2}\frac{N_{3}}{2\pi}b^{3}$
$+\frac{1}{2}\frac{1}{4\pi^{2}}\mathcal{B}_{m}^{1}V_{m}^{2}\beta^{3}+\frac{1}{2}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\phi_{new}^{2}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\phi^{2}\frac{N_{3}}{2\pi}b^{3}$
$+\frac{1}{2\pi}dV_{m}^{2}\wedge b^{2}+\frac{1}{2}\frac{1}{2\pi}\frac{N_{1}}{2\pi}dV_{m}^{2}a^{1}\beta^{3}-\frac{1}{2}\frac{1}{4\pi^{2}}dV_{m}^{2}V_{m}^{1}\beta^{3}+\frac{1}{2}\frac{1}{2\pi}dV_{m}^{2}\phi_{new}^{1}\frac{N_{3}}{2\pi}b^{3}$
$-\frac{1}{2}\frac{1}{4\pi^{2}}\mathcal{B}_{m}^{2}V_{m}^{1}\beta^{3}+\frac{1}{2}\frac{1}{2\pi}\mathcal{B}_{m}^{2}\phi_{new}^{1}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\phi^{1}\frac{N_{3}}{2\pi}b^{3}$ (194)
We now simplify $\Delta S_{aab}+\Delta S_{\text{coupling}}$. First, the term $\int_{M_{4}}-\frac{1}{2\pi}b^{1}dV_{m}^{1}-\frac{1}{2\pi}b^{2}dV_{m}^{2}$ from $\Delta S_{aab}$ cancels $\int_{M_{4}}\frac{1}{2\pi}dV_{m}^{1}b^{1}+\frac{1}{2\pi}dV_{m}^{2}b^{2}$ from $\Delta S_{\text{coupling}}$. Assuming $M_{4}$ is closed, we use
$\int_{M_{4}}d\left(\frac{1}{2}\frac{N_{2}}{2\pi}\frac{1}{2\pi}V_{m}^{1}a^{2}\beta^{3}\right)=0=\int_{M_{4}}\frac{1}{2}\frac{N_{2}}{2\pi}\frac{1}{2\pi}dV_{m}^{1}a^{2}\beta^{3}-0+\int_{M_{4}}\frac{1}{2}\frac{1}{2\pi}V_{m}^{1}\frac{N_{2}}{2\pi}a^{2}\frac{pN_{3}}{N_{123}}b^{3},$
which implies
$\int_{M_{4}}\underbrace{-\frac{1}{2}\frac{1}{2\pi}\frac{N_{2}}{2\pi}dV_{m}^{1}a^{2}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}-\underbrace{\frac{1}{2}\frac{1}{2\pi}V_{m}^{1}\frac{N_{2}}{2\pi}a^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{1}}V_{m}^{1}a^{2}b^{3}$ from $\Delta S_{aab}$}}=0.$ (195)
Similarly, from
$\int_{M_{4}}d\left(\frac{1}{2}\frac{1}{2\pi}\frac{N_{1}}{2\pi}V_{m}^{2}a^{1}\beta^{3}\right)=0=\int_{M_{4}}\frac{1}{2}\frac{1}{2\pi}dV_{m}^{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-0+\int_{M_{4}}\frac{1}{2}\frac{1}{2\pi}V_{m}^{2}\frac{N_{1}}{2\pi}a^{1}\frac{pN_{3}}{N_{123}}b^{3},$
we obtain
$\int_{M_{4}}\underbrace{\frac{1}{2}\frac{1}{2\pi}dV_{m}^{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}-\underbrace{\frac{1}{2}\frac{1}{2\pi}\frac{N_{1}}{2\pi}a^{1}V_{m}^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{2}}a^{1}V_{m}^{2}b^{3}$ from $\Delta S_{aab}$}}=0.$ (196)
Finally, using
$\int_{M_{4}}d\left(\frac{1}{2}\frac{1}{4\pi^{2}}V_{m}^{1}V_{m}^{2}\beta^{3}\right)=0=\int_{M_{4}}\frac{1}{2}\frac{1}{4\pi^{2}}dV_{m}^{1}V_{m}^{2}\beta^{3}-\frac{1}{2}\frac{1}{4\pi^{2}}V_{m}^{1}dV_{m}^{2}\beta^{3}+\frac{1}{2}\frac{1}{4\pi^{2}}V_{m}^{1}V_{m}^{2}\frac{pN_{3}}{N_{123}}b^{3},$
we find
$\int_{M_{4}}\underbrace{\frac{1}{2}\frac{1}{4\pi^{2}}dV_{m}^{1}V_{m}^{2}\beta^{3}-\frac{1}{2}\frac{1}{4\pi^{2}}dV_{m}^{2}V_{m}^{1}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}+\underbrace{\frac{1}{2}\frac{1}{4\pi^{2}}V_{m}^{1}V_{m}^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{1}N_{2}}V_{m}^{1}V_{m}^{2}b^{3}$ from $\Delta S_{aab}$ }}=0.$ (197)
Up to this point, we have reduced $\Delta S_{aab}+\Delta S_{\text{coupling}}$ to
$\Delta S_{aab}+\Delta S_{\text{coupling}}=$ $\int_{M_{4}}-\frac{1}{2}\frac{q}{N_{1}}V_{m}^{1}a^{2}b^{3}-\frac{1}{2}\frac{q}{N_{2}}a^{1}V_{m}^{2}b^{3}+\frac{1}{2}\frac{q}{N_{1}N_{2}}V_{m}^{1}V_{m}^{2}b^{3}$
$+\frac{1}{2}\frac{1}{2\pi}dV_{m}^{1}\phi_{new}^{2}\frac{N_{3}}{2\pi}b^{3}+\frac{1}{2}\frac{1}{4\pi^{2}}\mathcal{B}_{m}^{1}V_{m}^{2}\beta^{3}+\frac{1}{2}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\phi_{new}^{2}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\phi^{2}\frac{N_{3}}{2\pi}b^{3}$
$+\frac{1}{2}\frac{1}{2\pi}dV_{m}^{2}\phi_{new}^{1}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{1}{4\pi^{2}}\mathcal{B}_{m}^{2}V_{m}^{1}\beta^{3}+\frac{1}{2}\frac{1}{2\pi}\mathcal{B}_{m}^{2}\phi_{new}^{1}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\phi^{1}\frac{N_{3}}{2\pi}b^{3}$ (198)
We can now explain why we cannot find a bulk term $S_{\text{bulk}}$ such that $\Delta S_{aab}+\Delta S_{\text{coupling}}+\Delta S_{\text{bulk}}=0$ under the gauge transformations $\begin{cases}\mathcal{B}_{m}^{1}\rightarrow&\mathcal{B}_{m}^{1}+dV_{m}^{1}\\ a^{1}\rightarrow&a^{1}-\frac{1}{N_{1}}V_{m}^{1}\end{cases}$ and $\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases}$. The action $S_{\text{bulk}}$ must be a functional of background fields only, since $a^{i}$ and $b^{i}$ live only in (3+1)(3+1)D. However, $\Delta S_{aab}+\Delta S_{\text{coupling}}$ contains terms that depend explicitly on $a^{i}$ and $b^{i}$. Such terms cannot be canceled by $\Delta S_{\text{bulk}}$, because $\Delta S_{\text{bulk}}$ cannot involve $a^{i}$ or $b^{i}$. Therefore, a bulk term $S_{\text{bulk}}$ satisfying $\Delta S_{aab}+\Delta S_{\text{coupling}}+\Delta S_{\text{bulk}}=0$ does not exist.

### v3

@@
-Continuum-Field-Theoretical Construction of Conserved Currents, Non-Invertible Symmetries, and Mixed Anomalies in (3+1)D Non-Abelian Topological Order
+Non-invertible symmetries and mixed anomalies from conserved current construction in (3+1)D twisted $BF$ topological quantum field theories
 Abstract
-A modern perspective views topological order as a symmetry-breaking phase of generalized symmetries, making the identification of such symmetries, especially in non-Abelian topological orders, a fundamental task.
-In this work, we investigate generalized symmetries, with particular emphasis on non-invertible ones, in three-dimensional non-Abelian topological orders hosting both particle- and loop-like excitations.
-We adopt a continuum topological field theory description, focusing on twisted $BF$ theories with gauge group $G=\prod_{i}\mathbb{Z}_{N_{i}}$ and an $a\wedge a\wedge b$ twisted term.
-This field theory supports Borromean-Rings braiding and realizes non-Abelian topological order, which for $G=(\mathbb{Z}_{2})^{3}$ admits a microscopic realization via the $\mathbb{D}_{4}$ Kitaev quantum double lattice model.
-We systematically identify all generalized symmetry operators by extracting conserved currents from the equations of motion.
-Two distinct classes of currents emerge: type-I currents, which generate invertible higher-form symmetries, and type-II currents, which give rise to non-invertible higher-form symmetries.
-The non-invertibility originates from projectors accompanying the symmetry operators, which restrict admissible gauge-field configurations.
-We further analyze the fusion rules of these symmetries, showing that invertible symmetries admit inverses, while non-invertible symmetries fuse through multiple channels.
-Finally, we study mixed anomalies among these generalized symmetries by simultaneously coupling multiple currents to proper types of background gauge fields and examining their gaugeability.
-We identify two types of mixed anomalies: one cancellable by topological field theories in one higher dimension, and another representing an intrinsic gauging obstruction encoded in the (3+1)(3+1)D continuum topological field theory.
-Our results provide a concrete field-theoretical construction of conserved currents and a unified characterization of invertible and non-invertible higher-form symmetries in three-dimensional non-Abelian topological orders, offering a natural framework for studying (2+1)(2+1)D systems via symmetry topological field theory and topological holography.
+We develop a current-based construction of generalized symmetries in (3+1)(3+1)D twisted $BF$ topological quantum field theories (TQFTs), focusing on intrinsically non-invertible higher-form symmetries and their mixed anomalies.
+Starting from the equations of motion, we extract conserved currents and exponentiate the corresponding charges to obtain topological symmetry operators.
+This gives a step-by-step procedure for constructing symmetry operators, fusion, and anomaly diagnostics directly from the continuum action.
+We focus on twisted $BF$ theories with gauge group $G=\prod_{i}\mathbb{Z}_{N_{i}}$ and an $a\wedge a\wedge b$ twist, where $a$’s and $b$ are 1-form and 2-form gauge fields, respectively.
+These theories realize non-Abelian (3+1)(3+1)D TQFTs supporting Borromean-rings braiding and describe three-dimensional non-Abelian topological orders in condensed matter.
+For $G=(\mathbb{Z}_{2})^{3}$, a microscopic realization is given by the $\mathbb{D}_{4}$ Kitaev quantum double model.
+Two distinct classes of conserved currents emerge: Type-I currents generate invertible higher-form symmetries with group-like fusion, while Type-II currents require additional consistency conditions on gauge-field configurations, leading to intrinsically non-invertible symmetries dressed by projectors.
+We compute the fusion algebra: invertible operators admit inverses, while non-invertible ones exhibit multi-channel fusion governed by projector fusion.
+We diagnose mixed anomalies by coupling multiple conserved currents to background gauge fields, revealing two outcomes: anomalies canceled by anomaly inflow from a higher-dimensional theory, and intrinsic gauging obstructions encoded in the (3+1)(3+1)D continuum theory.
+Overall, our results provide a unified and practical approach for constructing and characterizing higher-form symmetries, which can be extended to more general TQFTs and topological orders.
 I Introduction
 Symmetry has long been a central organizing principle in physics.
-Broadly speaking, a symmetry of a physical system is a transformation that leaves the underlying theory invariant when applied to the system as a whole.
+Broadly speaking, a symmetry of a physical system is a transformation that leaves the underlying theory invariant when applied to the system as a whole; equivalently, in the operator language it is implemented by an operator commuting with the Hamiltonian (or with the dynamics in the continuum).
 A familiar example is provided by the one-dimensional [^1] In this paper, “(n+1)(n+1)D” refers to $(n+1)$-dimensional spacetime with $n$-dimensional real space.
 We avoid using “nnD” or “(n)(n)D” unless otherwise specified.
-When referring specifically to spatial dimensions or to the dimensions of geometric objects, we use “nnd”, e.g., 33d ground state, 33d topological order, 22d square lattice, and 11d submanifold.
+When referring specifically to spatial dimensions or to the dimensions of geometric objects, we use “nnd”, e.g., 33d ground state, 33d topological order, and 22d square lattice.
 (11d) transverse-field Ising model, which is invariant under a simultaneous flip of all spins.
@@
 Local operators may transform nontrivially under this operation and are said to carry symmetry charge.
-Since performing the global spin flip twice returns the system to itself, this symmetry has a $\mathbb{Z}_{2}$ group structure.
+Since performing the global spin flip twice returns the system to itself, this symmetry is invertible and has a $\mathbb{Z}_{2}$ group structure.
 In recent years, the notion of symmetry has been substantially generalized from multiple perspectives across condensed matter, high-energy, and mathematical physics [1, 2, 3, 4, 5, 6, 7, 8].
@@
 First, symmetry operations need not act on the entire system; instead, they may be supported on submanifolds [1].
-As a result, charged operators become extended rather than point-like, leading to the concept of higher-form symmetry.
+As a result, symmetry operators can be extended (line/surface) operators, and charged operators become extended rather than point-like, leading to the concept of higher-form symmetry.
 Within this framework, the $\mathbb{Z}_{2}$ symmetry of the 11d transverse-field Ising model discussed above is an example of a 0-form symmetry.
-A canonical illustration of higher-form symmetry is provided by the 22d toric code model [9] on a square lattice, where symmetry operators can be constructed on closed loops rather than on the full lattice.
-These loop operators commute with the Hamiltonian and generate symmetries associated with one-dimensional submanifolds embedded in the two-dimensional system.
+Second, the algebraic structure organizing symmetry operators need not be a group [9, 10, 11, 12, 13, 14, 15, 7, 16, 17].
+More generally, symmetry is encoded by the fusion of topological operators/defects, and is naturally formulated using the language of category theory, giving rise to non-invertible or categorical symmetries.
+As an illustration of the first aspect, the 22d toric code model [18] on a square lattice admits symmetry operators supported on closed loops rather than on the full lattice.
+In condensed-matter language, these loop operators commute with the Hamiltonian and act within the topological ground-state sector; in the continuum/TQFT language, they are topological line operators (Wilson loops) generating higher-form symmetries.
 Operators supported on one-dimensional paths that intersect these loops transform nontrivially under the symmetry and are therefore charged.
 Such symmetries are referred to as $1$-form symmetries, reflecting the co-dimensionality of their support in 22d real space.
-At long wavelengths, the supports of higher-form symmetries are manifold-like and can be continuously deformed without affecting their action.
+At long wavelengths, higher-form symmetry operators are topological in the sense that their supports are manifold-like and can be continuously deformed without affecting their action.
 There exists another important generalization, namely subsystem symmetries, whose supports are rigid rather than deformable and are typically fixed to specific geometric patterns such as lines, fractals, or more exotic structures [19, 20, 21, 22].
 In this work, however, we focus exclusively on higher-form symmetries and leave subsystem symmetries for future investigation.
-Second, the algebraic structure organizing symmetry operators need not be a group [14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
-More generally, it may be described using the language of category theory, giving rise to non-invertible or categorical symmetries.
-A representative example arises in the 11d transverse-field Ising model at criticality, which admits a symmetry operation implementing Kramers-Wannier duality.
+As an illustration of the second aspect, a representative example of a non-invertible symmetry arises in the 11d transverse-field Ising model at criticality, which admits a symmetry operation implementing Kramers–Wannier duality.
 This operation exchanges spin and domain-wall variables and commutes with the Hamiltonian, yet it does not admit an inverse in the conventional sense: acting with it projects out a subset of states.
@@
 These modern generalizations of symmetry provide powerful tools for the study of quantum many-body systems, particularly topologically ordered phases.
-Such phases, traditionally viewed as lying beyond the Landau-Ginzburg paradigm, can be reinterpreted as phases that spontaneously break generalized symmetries.
-For example, the 22d toric code model on a torus exhibits a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 11-form symmetry generated by operators supported on the two noncontractible cycles of the torus.
+Such phases, traditionally viewed as lying beyond the Landau–Ginzburg paradigm, can be reinterpreted as phases characterized by generalized-symmetry structures, and are often described as exhibiting spontaneous breaking of generalized symmetries in the topological ground-state sector.
+For example, the 22d toric code model on a torus exhibits a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 11-form symmetry generated by loop operators supported on the two noncontractible cycles of the torus.
 Since these operators do not commute, no ground state can remain invariant under both simultaneously, leading inevitably to topological ground-state degeneracy.
-This phenomenon can be understood as spontaneous breaking of a higher-form symmetry.
+This phenomenon can be viewed as a manifestation of higher-form symmetry breaking in the ground-state subspace.
 In this way, generalized symmetries offer a unifying framework for understanding both the structure and robustness of topologically ordered phases.
-Building on the concept of generalized symmetry, a framework known as symmetry topological field theory (SymTFT), also referred to as the symmetry-topological-order correspondence, has been rapidly developed.
-The central idea is that a $(d+1)$-dimensional topological quantum field theory (TQFT), or equivalently a (d+1)(d+1)D topological order, encodes the symmetries of a (d1+1)(d-1+1)-dimensional quantum field theory.
-In this correspondence, topological operators in the bulk encode generalized symmetries of the TQFT and can terminate on the boundary in different ways, thereby characterizing distinct lower-dimensional systems.
-This framework has found broad applications, including the classification of gapped and gapless phases, the diagnosis of generalized symmetry anomalies, the analysis of renormalization-group flows, the study of mixed-state phases and phase transitions, and so on [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
+Building on the concept of generalized symmetry, a framework known as symmetry topological field theory (SymTFT) has been rapidly developed.
+The central idea is that a $(d+1)$-dimensional topological quantum field theory (TQFT), or equivalently a (d+1)(d+1)D topological order in condensed-matter language, encodes the generalized symmetries of a (d1+1)(d-1+1)-dimensional quantum field theory.
+In this correspondence, topological operators/defects in the bulk encode generalized symmetries and can terminate on the boundary in different ways, thereby characterizing distinct lower-dimensional systems.
+This framework has found broad applications, including the classification of gapped and gapless phases, the diagnosis of generalized-symmetry anomalies, the analysis of renormalization-group flows, the study of mixed-state phases and phase transitions, and so on [15, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33].
 To date, SymTFT has been particularly successful for (2+1)(2+1)D bulks and (1+1)(1+1)D boundaries.
-Extending this framework to (2+1)(2+1)D systems therefore naturally motivates a detailed understanding of generalized symmetries in (3+1)(3+1)D topologically ordered phases.
+Extending this framework to (2+1)(2+1)D boundary systems therefore naturally motivates a detailed understanding of generalized symmetries realized in (3+1)(3+1)D TQFTs/topologically ordered phases.
 As effective field theories of topological order, TQFTs are expected to encode generalized symmetries in a natural and systematic manner.
-For instance, Wilson operators in Chern-Simons theory [36, 37, 38] act as generators of $1$-form symmetries.
-Analogous to the role of Chern-Simons theory in describing 22d topological orders, $BF$ theory [39, 40], together with its twisted variants [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], provides a unified field-theoretical framework for characterizing 33d topological orders and gauged symmetry-protected topological phases with particle and loop excitations.
-Geometrically, the canonical $BF$ term bdab\,da (with $b$ a $2$-form field and $a$ a $1$-form field) captures the characteristic particle-loop braiding in 33d topological order.
+For instance, Wilson operators in Chern–Simons theory [34, 35, 36] act as generators of $1$-form symmetries.
+Analogous to the role of Chern–Simons theory in describing 22d topological orders, $BF$ theory [37, 38], together with its twisted variants [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49], provides a unified field-theoretical framework for characterizing 33d topological orders and gauged symmetry-protected topological phases with particle and loop excitations.
+Geometrically, the canonical $BF$ term bdab\,da (with $b$ a $2$-form field and $a$ a $1$-form field) captures the characteristic particle–loop braiding in 33d topological order.
 [^2] Throughout this paper, we adopt a condensed notation for differential forms and omit the wedge symbol $\wedge$ when no confusion arises.
 For example, bdab\,da stands for $b\wedge da$, and similarly aadaaada denotes topological terms of $a\wedge a\wedge da$ type.
-More intricate braiding processes, including multi-loop and Borromean rings braiding, are encoded by twisted terms such as aadaaada, aaaaaaaa, and aabaab [40, 44, 48, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65].
+More intricate braiding processes, including multi-loop and Borromean-Rings braiding, are encoded by twisted terms such as aadaaada, aaaaaaaa, and aabaab [38, 42, 46, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63].
 These braiding processes are subject to compatibility conditions imposed by gauge invariance [47].
@@
 [76].
-Given this progress, a natural question arises: how can generalized symmetries of 33d topological orders be systematically identified directly from effective field theory?
-In this work, we address this question by studying generalized symmetries in a class of (3+1)(3+1)D topological quantum field theories describing three-dimensional non-Abelian topological orders with nontrivial Borromean-Rings braiding.
-We focus on twisted $BF$ theories with gauge group $G=\prod_{i}\mathbb{Z}_{N_{i}}$ and an $a\wedge a\wedge b$ twisted term [48, 49, 76].
-For the special case $G=(\mathbb{Z}_{2})^{3}$, the resulting non-Abelian topological order admits a microscopic realization in terms of the $\mathbb{D}_{4}$ Kitaev quantum double lattice model [79].
-These field theories therefore provide a concrete continuum description in which higher-form generalized symmetries, including non-invertible ones, can be systematically analyzed.
-In quantum mechanics, an operator commuting with the Hamiltonian signals a symmetry, while in field theory the corresponding notion is a conserved current.
-Guided by this principle, we identify generalized symmetries directly from the effective field theory by extracting conserved currents from the equations of motion of the gauge fields.
-We find that the equations of motion naturally separate into two distinct classes, leading to two types of conserved currents.
-Type-I currents are conserved identically and generate invertible higher-form symmetries.
-By contrast, type-II currents are conserved only upon imposing additional constraints on the gauge-field configurations.
-As a result, the associated symmetry operators necessarily involve projector structures that enforce these constraints.
-Since such projectors do not admit inverses, the corresponding higher-form symmetries are intrinsically non-invertible.
-With the explicit symmetry operators constructed from these conserved currents, we analyze their fusion rules by examining the composition of symmetry operations.
-We show that invertible higher-form symmetries possess well-defined inverses and obey group-like fusion rules, whereas non-invertible symmetries fuse through multiple channels, reflecting the presence of projector-induced constraints.
-Finally, we investigate mixed anomalies among these generalized symmetries by simultaneously coupling multiple conserved currents to appropriate background gauge fields and attempting to gauge the corresponding symmetries.
-We identify two distinct types of mixed anomalies: one that can be canceled by embedding the theory into a topological field theory in one higher dimension, and another that represents an intrinsic gauging obstruction already encoded in the (3+1)(3+1)D continuum topological field theory.
+Given this progress, a natural question arises: how can generalized symmetries of 33d topological orders—equivalently, of their (3+1)(3+1)D effective TQFT descriptions—be systematically identified directly from continuum field theory, together with their fusion and anomaly data?
+Reader’s guide (condensed-matter vs. TQFT language).
+To minimize translation cost across communities, we will freely use both terminologies.
+In particular, “33d topological order” refers to the gapped topological sector captured by a (3+1)(3+1)D TQFT; point/loop excitations in the many-body system correspond to line/surface operators (topological defects) in the TQFT; and generalized symmetries are implemented by topological operators supported on submanifolds, whose fusion encodes the symmetry algebra.
+When we speak of “gauging” a generalized symmetry, we mean coupling to background gauge fields and testing gaugeability (possibly after adding local counterterms), thereby diagnosing potential ’t Hooft anomalies.
+This work: a constructive and interdisciplinary route.
+In this work we provide a fully constructive, model-resolved derivation of generalized symmetries in a nontrivial family of (3+1)(3+1)D TQFTs/topological orders.
+Our central output is a current-to-operator pipeline: starting from the action and equations of motion, we extract conserved currents and explicitly construct the associated topological symmetry operators/defects, together with their fusion rules and (anomalous) gaugeability.
+The same construction is intended to be useful from both sides of the condensed-matter/HEP interface: it offers a practical way to compute generalized-symmetry data of concrete three-dimensional topological orders described by twisted $BF$-type effective theories, and it organizes that data directly in the defect/anomaly language natural to continuum TQFT.
+Concretely, we study twisted $BF$ theories with gauge group $G=\prod_{i}\mathbb{Z}_{N_{i}}$ and an $a\wedge a\wedge b$ twist [46, 47, 74], which realize non-Abelian (3+1)(3+1)D TQFTs supporting Borromean-Rings braiding and serve as effective continuum descriptions of certain three-dimensional non-Abelian topological orders.
+For the special case $G=(\mathbb{Z}_{2})^{3}$, the corresponding non-Abelian topological order admits a microscopic realization in the $\mathbb{D}_{4}$ Kitaev quantum double lattice model [77].
+This setting provides a concrete arena in which higher-form generalized symmetries—including intrinsically non-invertible ones—can be identified and analyzed systematically.
+Our approach is current-based.
+In quantum mechanics, an operator commuting with the Hamiltonian signals a symmetry; in continuum field theory, the corresponding diagnostic is a conserved current.
+Guided by this principle, we derive the relevant continuity equations from the equations of motion and use the resulting conserved currents to construct symmetry operators/defects.
+In this way, the generalized-symmetry data are obtained directly from conservation laws, and the same consistency conditions that make the currents well-defined determine the resulting fusion structure and (anomalous) gaugeability.
+Main results and roadmap.
+Our results can be summarized as a concrete route from continuum equations to symmetry data of (3+1)(3+1)D TQFTs/topological orders:
+  • Two classes of conserved currents from the equations of motion.
+We show that the equations of motion in the $a\wedge a\wedge b$ twisted $BF$ theories naturally separate into two distinct classes, leading to two qualitatively different types of conserved currents.
+Type-I currents are conserved identically (in the sense of Bianchi/Noether-type identities) and generate invertible higher-form symmetries.
+Type-II currents become conserved only after imposing additional constraints on admissible gauge-field configurations; these constraints are intrinsic to the twisted $BF$ structure and are required for the corresponding conserved quantities to be well-defined.
+  • Explicit symmetry operators, including intrinsically non-invertible higher-form symmetries.
+From the type-I currents we construct standard topological symmetry operators with group-like composition, which in physical terms are the continuum counterparts of Wilson operators for particle/loop excitations.
+From the type-II currents we construct symmetry operators that must be dressed by projectors enforcing the above constraints.
+Because these projectors do not admit inverses, the resulting higher-form symmetries are intrinsically non-invertible.
+In particular, non-invertibility in our setting is not postulated: it is enforced by the consistency conditions behind the conservation laws.
+  • Fusion algebra: group-like fusion vs.
+multi-channel fusion controlled by projector fusion.
+With explicit operators at hand, we compute the fusion rules by composing symmetry operators supported on the same submanifold (equivalently, by bringing the corresponding defects together).
+Invertible higher-form symmetry operators admit inverses and obey group-like fusion rules.
+By contrast, non-invertible higher-form symmetry operators exhibit genuinely non-group-like behavior: the product decomposes into multiple channels, and the channel structure is dictated by the fusion of the projector factors.
+We present representative examples that illustrate how the projector constraints reorganize the fusion algebra in a systematic, computable way.
+  • Mixed anomalies and gauging: inflow-cancelable anomalies vs.
+intrinsic gauging obstructions.
+We diagnose anomalies by coupling multiple conserved currents to appropriate background gauge fields and testing gaugeability.
+This reveals mixed anomalies among the generalized symmetries (including between invertible and non-invertible ones).
+We further identify two qualitatively different outcomes: (i) anomalies that can be canceled by embedding into a topological field theory in one higher dimension (anomaly inflow), and (ii) intrinsic gauging obstructions already encoded in the (3+1)(3+1)D continuum theory.
+We illustrate these diagnostics both in familiar settings (e.g., the 33d $\mathbb{Z}_{2}$ topological order) and in the Borromean-Rings topological order described by the twisted $BF$ theories.
+This structure makes the paper self-contained from a field-theory viewpoint: starting from the action and equations of motion, we extract conserved currents, construct the corresponding topological symmetry operators/defects, compute their fusion algebra, and finally analyze their (anomalous) gaugeability.
 This paper is organized as follows.
@@
 In Sec.
-III, we study generalized symmetries in the effective field theory of a class of 33d non-Abelian topological orders and construct the corresponding symmetry operators from conserved currents.
-Sec.p IV is devoted to the analysis of fusion rules.
+III, we study generalized symmetries in the effective field theory description of a class of 33d non-Abelian topological orders and construct the corresponding symmetry operators from conserved currents.
+Sec.
+IV is devoted to the analysis of fusion rules.
 In Sec.
@@
 II Preliminaries of conserved currents and symmetry operators
-In field theory, a symmetry is indicated by a conserved current, which in turn gives rise to a topological operator, namely the symmetry operator associated with the symmetry.
-From this perspective, a symmetry refers to the mathematical structure formed by these topological operators.
+In relativistic field theory, a continuous symmetry is reflected in a conserved current, and the corresponding conserved charge exponentiates to a topological operator—the symmetry operator.
+From this viewpoint, a “symmetry” refers to the mathematical structure formed by these topological operators (or, equivalently, by their fusion/commutation relations).
 In our discussion of generalized symmetries, we adopt the viewpoint that a higher-form symmetry is generated by topological operators originating from higher-form conserved currents.
-Moreover, this principle naturally extends to the construction of non-invertible symmetries, whose symmetry operators can be traced back to a special class of conserved currents, as will be discussed in Sec.
+Moreover, this principle naturally extends to the construction of non-invertible symmetries: their symmetry operators can be traced back to a special class of conserved currents (together with additional constraints), as will be discussed in Sec.
 III.
@@
 The existence of this global symmetry can thus be traced back to the presence of a conserved current.
-The conservation law ensures that $U_{\alpha}$ is invariant under smooth deformations of $\Sigma_{3}$, hence the terminology topological operator of the 0-form symmetry (i.e., global symmetry) that act on the entire quantum system.
+The conservation law ensures that $U_{\alpha}$ is invariant under smooth deformations of $\Sigma_{3}$, hence the terminology of a topological operator for the 0-form (global) symmetry acting on the entire quantum system.
 The concept of higher-form symmetry, e.g., $p$-form symmetry, generalizes the above structure by allowing the conserved current to be a $(p+1)$-form.
 Correspondingly, the associated symmetry operator is supported on a $\left(D-1-p\right)$-dimensional submanifold in (D1+1)(D-1+1)D spacetime.
-As an illustrative example of higher-form symmetries, we consider the (3+1)(3+1)D Maxwell’s electromagnetic theory with action $S=\frac{1}{4}\int d^{4}xF_{\mu\nu}F^{\mu\nu}$, where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ and $e$ is the gauge coupling constant; also see the similar discussion in, e.g., Refs [8, 1, 3, 6].
+As an illustrative example of higher-form symmetries, we consider the (3+1)(3+1)D Maxwell electromagnetic theory with action $S=\frac{1}{4}\int d^{4}x\,F_{\mu\nu}F^{\mu\nu}$, where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$; for simplicity we set the gauge coupling $e=1$ (equivalently, absorb it into the normalization of fields).
+See also the similar discussion in, e.g., Refs [8, 1, 3, 6].
 First, there exists a conserved $2$-form current satisfying $d\left(*J_{m}\right)=0$, with $J_{m}^{\mu\nu}=\frac{1}{4\pi}\epsilon^{\mu\nu\lambda\sigma}F_{\lambda\sigma}=\frac{1}{2\pi}\left(d\widetilde{A}\right)^{\mu\nu}$.
@@
 $U_{\alpha}^{\left(e\right)}\left(\Sigma_{2}\right)=\exp\left({\rm i}\alpha\int_{\Sigma_{2}}*J_{e}\right)=\exp\left({\rm i}\alpha\int_{\Sigma_{2}}*F\right).$ (3)
-Another illustrative example is provided by the mutual Chern–Simons theory, which serves as the effective field theory of the 22d topological order,
+Another illustrative example is provided by the mutual Chern–Simons theory, which serves as the effective field theory (i.e., a TQFT description) of a 22d topological order,
 $S=\int_{M_{3}}\frac{2}{2\pi}a^{1}da^{2}.$ (4)
-The lattice realization of this field theory is the celebrated toric code model in 2d square lattice.
+The lattice realization of this field theory is the celebrated toric code model in 22d square lattice.
 Here $a^{1}$ and $a^{2}$ are $1$-form gauge fields.
@@
 These quantities measure the number of topological excitations enclosed by the area bounded by a loop $\gamma$.
-They generate two minimal symmetry operators,
+They generate two minimal symmetry operators (topological line operators),
 $U_{e}\left(\gamma\right)=\exp\left({\rm i}\int_{\gamma}a^{2}\right)\,,\,\,\,U_{m}\left(\gamma\right)=\exp\left({\rm i}\int_{\gamma}a^{1}\right)$ (5)
@@
 III Field-theoretical construction of symmetry operators
-In this section, we construct symmetry operators in effective field theories of 33d topological orders.
-We begin with the $\left(3+1\right)$D $\mathbb{Z}_{N}$ BFBF theory as a warm-up, which describes $\mathbb{Z}_{N}$ topological orders.
+In this section, we construct symmetry operators (topological operators/defects) from conserved currents in (3+1)(3+1)D TQFTs that serve as effective field-theoretical descriptions of 33d topological orders.
+We begin with the (3+1)(3+1)D $\mathbb{Z}_{N}$ BFBF theory as a warm-up, which describes $\mathbb{Z}_{N}$ topological orders.
 We then turn to the main goal of this paper: identifying symmetry operators in the effective field theory of a class of 33d non-Abelian topological orders, namely Borromean-Rings (BR) topological orders.
-In particular, we show how conserved currents can be exploited to construct non-invertible symmetry operators.
+In particular, we show how the conserved-current construction can be exploited to obtain intrinsically non-invertible symmetry operators.
 A system with BR topological order [46, 47, 74, 77] exhibits nontrivial BR braiding statistics.
@@
 BR braiding is also distinct from multi-loop braiding because it necessarily involves a particle excitation.
-The corresponding TQFT is a $BF$ theory with an aabaab-type twisted term; see Eq.
+The corresponding TQFT is a twisted $BF$ theory with an aabaab-type term; see Eq.
 (14) below.
-Remarkably, BR topological order supports non-Abelian particle and loop excitations even when all excitations carry gauge charges and fluxes of an Abelian gauge group, making it an ideal playground for exploring generalized symmetries in $\left(3+1\right)$D systems.
+Remarkably, BR topological order supports non-Abelian particle and loop excitations even when all excitations carry gauge charges and fluxes of an Abelian gauge group, making it an ideal playground for exploring generalized symmetries—including non-invertible ones—in (3+1)(3+1)D topological phases/TQFTs.
 In the following, we first derive the equations of motion (EoM) of gauge fields, from which we identify conserved currents.
-These conserved currents imply symmetries and allow us to construct the corresponding symmetry operators.
-We find two classes of conserved currents denoted as type-I and type-II respectively, as shown in Sec.
+These conserved currents imply generalized symmetries and allow us to construct the corresponding symmetry operators.
+We find two classes of conserved currents, denoted as type-I and type-II, respectively, as shown in Sec.
 III.3 and Sec.
 III.4.
-Type-I currents lead to group-like (i.e., invertible) symmetries.
-Type-II currents are conserved only under certain constraints such that corresponding conserved quantities are legitimate.
+Type-I currents lead to group-like (i.e., invertible) higher-form symmetries.
+By contrast, type-II currents are conserved only under additional constraints such that the corresponding conserved quantities are well-defined.
 Accordingly, the symmetry operators generated by these conserved quantities must be accompanied by projectors that enforce the constraints.
@@
 The topological field theories describing the 22d and 33d toric code models discussed above are $\mathbb{Z}_{2}$ BFBF theories.
-It is straightforward to generalize the gauge group from $\mathbb{Z}_{2}$ to a general $\mathbb{Z}_{N}$.
-The $\left(3+1\right)$D $\mathbb{Z}_{N}$ BFBF theory is described by
+It is straightforward to generalize the gauge group from $\mathbb{Z}_{2}$ to $\mathbb{Z}_{N}$.
+The (3+1)(3+1)D $\mathbb{Z}_{N}$ BFBF theory is described by
 $S=\int_{M_{4}}\frac{N}{2\pi}bda.$ (6)
 Here $b$ is a $2$-form gauge field.
-The equations of motion (EoMs) are
+The EoMs are
 $\frac{N}{2\pi}da=0\quad\text{and}\quad\frac{N}{2\pi}db=0.$ (7)
@@
 $d\left(*J_{m}\right)=0\quad\text{with}\quad J_{m}=*b,$ (9)
-where $J_{e}$ and $J_{m}$ are $3$-form and $2$-form conserved currents, respectively. Analogous to ordinary 0-form symmetries, these conserved currents give rise to conserved quantities
+where $J_{e}$ and $J_{m}$ are $3$-form and $2$-form conserved currents, respectively. Analogous to ordinary 0-form symmetries, these conserved currents give rise to conserved quantities,
 $Q_{e}\left(\gamma\right)=\int_{\gamma}*J_{e}=\int_{\gamma}a\,,\,\,\,Q_{m}\left(\sigma\right)=\int_{\sigma}*J_{m}=\int_{\sigma}b,$ (10)
@@
 Acting with $e^{{\rm i}\int_{\gamma}a}$ on a state amounts to transporting the particle along $\gamma$ in spacetime.
-This process probes the gauge flux carried by loop excitations piercing the surface bounded by $\gamma$.
+This process probes the gauge flux carried by loop excitations piercing a surface bounded by $\gamma$.
 From the perspective of conserved currents, since $a=a_{\mu}dx^{\mu}$, the current takes the form $J_{e}=-\frac{1}{3!}\epsilon^{\mu\nu\lambda\rho}a_{\mu}\,dx^{\nu}\wedge dx^{\lambda}\wedge dx^{\rho}$.
-For comparison, in a $\left(3+1\right)$D system with a 0-form $U\left(1\right)$ symmetry, the particle current $j=j_{\mu}dx^{\mu}$ is a $1$-form current, whose Hodge dual is $j=j_{0}dx^{1}\wedge dx^{2}\wedge dx^{3}-j_{1}dx^{0}\wedge dx^{2}\wedge dx^{3}+j_{2}dx^{0}\wedge dx^{1}\wedge dx^{3}-j_{3}dx^{0}\wedge dx^{1}\wedge dx^{2}$.
+For comparison, in a (3+1)(3+1)D system with a 0-form $U\left(1\right)$ symmetry, the particle current $j=j_{\mu}dx^{\mu}$ is a $1$-form current, whose Hodge dual is $j=j_{0}dx^{1}\wedge dx^{2}\wedge dx^{3}-j_{1}dx^{0}\wedge dx^{2}\wedge dx^{3}+j_{2}dx^{0}\wedge dx^{1}\wedge dx^{3}-j_{3}dx^{0}\wedge dx^{1}\wedge dx^{2}$.
 Taking $M_{3}$ to be the entire 33d spatial manifold, $\int_{M_{3}}*j$ integrates the particle-number density $j_{0}$ over $M_{3}$, yielding the total particle number.
 By direct analogy, $\int_{\gamma}*J_{e}$ can be interpreted as a generalized “symmetry charge” associated with the submanifold $\gamma$.
-Recall that $\int_{\gamma}*J_{e}$ counts the gauge fluxes encircled by $\gamma$; in $\left(3+1\right)$D, this quantity is given by the linking number between $\gamma$ and the worldsheets $\sigma$ of loop excitations.
+Recall that $\int_{\gamma}*J_{e}$ counts the gauge flux encircled by $\gamma$; in (3+1)(3+1)D, this quantity is given by the linking number between $\gamma$ and the worldsheets $\sigma$ of loop excitations.
 Equivalently, it can be computed by counting the intersections between $\gamma$ and a three-dimensional volume bounded by $\sigma$ in spacetime.
-In this sense, $J_{e}$ describes the flow of $2$-dimensional objects in $\left(3+1\right)$D spacetime, in close analogy with $j$ describing the flow of 0-dimensional particles.
-These 22d objects are precisely the worldsheets of loop excitations.
+In this sense, $J_{e}$ describes the flow of two-dimensional objects in (3+1)(3+1)D spacetime, in close analogy with $j$ describing the flow of point-like particles.
+These two-dimensional objects are precisely the worldsheets of loop excitations.
 The physical interpretation of $Q_{m}\left(\sigma\right)$ follows in a completely analogous manner.
@@
 The equations of motion imply that $\oint a\in\frac{2\pi}{N}\mathbb{Z}_{N}$, which in turn gives $NQ_{e}\left(\gamma\right)\in 2\pi\mathbb{Z}$, meaning that $q_{e}$ is equivalent to $q_{e}+N$.
-Furthermore, large gauge invariance under $\int_{\gamma}a\rightarrow\int_{\gamma}a+2\pi$ requires $q_{e}$ to be an integer.
+Furthermore, large-gauge invariance under $\int_{\gamma}a\rightarrow\int_{\gamma}a+2\pi$ requires $q_{e}$ to be an integer.
 Consequently, $q_{e}$ takes values in $\mathbb{Z}/N\mathbb{Z}$.
@@
 The BR topological order [46, 47, 74, 77] can be effectively captured by a topological field theory with gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$, whose action takes the form
-$S_{aab}=\int\sum_{i=1}^{3}\frac{N_{i}}{2\pi}b^{i}da^{i}+qa^{1}a^{2}b^{3},$ (14)
+$S_{\text{BR}}=\int\sum_{i=1}^{3}\frac{N_{i}}{2\pi}b^{i}da^{i}+qa^{1}a^{2}b^{3},$ (14)
 where $a^{i}$ and $b^{i}$ are $1$-form and $2$-form gauge fields, respectively.
@@
 The EoMs are obtained by requiring the action to remain invariant (modulo $2\pi$) under infinitesimal variations of the gauge fields.
-For instance, under $b^{1}\rightarrow b^{1}+\delta b^{1}$, the condition $\delta S\left[\delta b^{1}\right]=0\mod{2\pi}$ implies $\frac{N_{1}}{2\pi}da^{1}=0$, which is the equation of motion for $a^{1}$.
-The full set of EoMs, together with the corresponding conserved currents and symmetry operators of $S_{aab}$, is summarized in Table 1.
+For instance, under $b^{1}\rightarrow b^{1}+\delta b^{1}$, the condition $\delta S\left[\delta b^{1}\right]=0\mod{2\pi}$ implies $\frac{N_{1}}{2\pi}da^{1}=0$, which is the equation of motion for $b^{1}$.
+The full set of EoMs, together with the corresponding conserved currents and symmetry operators of $S_{\text{BR}}$, is summarized in Table 1.
 Among these symmetry operators, some are invertible whereas others are non-invertible.
 In particular, due to the presence of the twisted term, two distinct classes of EoMs and conserved currents emerge, which we refer to as type-I and type-II, respectively, and which will be discussed in detail in the following subsections.
-Table 1: Equations of motion (EoMs), conserved currents, and symmetry operators of $S_{aab}=\int\sum_{i=1}^{3}\frac{N_{i}}{2\pi}b^{i}da^{i}+qa^{1}a^{2}b^{3}$, where $q=\frac{pN_{1}N_{2}N_{3}}{\left(2\pi\right)^{2}N_{123}}$ with $p\in\mathbb{Z}_{N_{123}}$.
+Table 1: Equations of motion (EoMs), conserved currents, and symmetry operators of $S_{\text{BR}}=\int\sum_{i=1}^{3}\frac{N_{i}}{2\pi}b^{i}da^{i}+qa^{1}a^{2}b^{3}$, where $q=\frac{pN_{1}N_{2}N_{3}}{\left(2\pi\right)^{2}N_{123}}$ with $p\in\mathbb{Z}_{N_{123}}$.
 The first three symmetry operators are invertible and correspond to Wilson operators of Abelian particle and Abelian loop excitations.
@@
 They take the same functional form as Wilson operators associated with non-Abelian loop and non-Abelian particle excitations, up to overall normalization factors.
-The delta functions appearing in these symmetry operators enforce constraints on gauge-field configurations, under which the corresponding currents are conserved and define genuine symmetries (see Sec.
+The delta functions appearing in these symmetry operators enforce constraints on gauge-field configurations; under these constraints, the corresponding currents are conserved and define genuine symmetries (see Sec.
 III.4).
@@
 IV for a detailed discussion.
-EoM Currents Continuity Eq. Symmetry operators
+EoMs Currents Continuity Eq. Symmetry operators
 $\frac{N_{1}}{2\pi}da^{1}=0$ $J_{e}^{1}=-*a^{1}$ $d\left(*J_{e}^{1}\right)=0$ $U_{e_{1}}\left(\gamma\right)=\exp\left({\rm i}e_{1}\int_{\gamma}a^{1}\right)$
@@
 III.3 Type-I conserved currents and invertible symmetry operators
-We begin by considering a local perturbation of the gauge field $b^{1}$.
-Under $b^{1}\rightarrow b^{1}+\delta b^{1}$, the variation of the action is $\delta S\left[\delta b^{1}\right]=\int_{M^{4}}\frac{N_{1}}{2\pi}\delta b^{1}da^{1}$.
-Requiring $\delta S\left[\delta b^{1}\right]=0$ for an arbitrary infinitesimal perturbation $\delta b^{1}$, we obtain the corresponding equation of motion
+We begin by considering a local variation of the gauge field $b^{1}$. Under $b^{1}\rightarrow b^{1}+\delta b^{1}$, the variation of the action is
+$\delta S\left[\delta b^{1}\right]=\int_{M^{4}}\frac{N_{1}}{2\pi}\,\delta b^{1}\,da^{1}.$
+Requiring $\delta S\left[\delta b^{1}\right]=0$ for an arbitrary infinitesimal perturbation $\delta b^{1}$, we obtain the equation of motion
 $\frac{N_{1}}{2\pi}\,da^{1}=0.$ (18)
-This EoM implies flux quantization, $\oint_{\gamma}a^{1}=\frac{2\pi k_{1}}{N_{1}}$ with $k_{1}\in\mathbb{Z}_{N_{1}}$; namely, the flux of $a^{1}$ piercing the surface enclosed by $\gamma$ is quantized in integer multiples of $\frac{2\pi}{N_{1}}$.
-Since this EoM takes the form of a total derivative set to zero, it naturally resembles a continuity equation.
+This EoM implies flux quantization: $\oint_{\gamma}a^{1}=\frac{2\pi k_{1}}{N_{1}}$ with $k_{1}\in\mathbb{Z}_{N_{1}}$.
+Equivalently, the flux of $a^{1}$ piercing a surface bounded by $\gamma$ is quantized in integer multiples of $\frac{2\pi}{N_{1}}$.
+Since this EoM is a total-derivative condition, it naturally takes the form of a continuity equation.
 Indeed, from Eq.
-(18) we can extract the continuity equation
+(18) we can extract
 $d\left(*J^{1}_{e}\right)=0\quad\text{with}\quad J^{1}_{e}=-*a^{1},$ (19)
@@
 $Q_{1,\gamma}=\int_{\gamma}*J^{1}_{e}=\int_{\gamma}a^{1},$ (20)
-which depends only on the topological class of the closed curve $\gamma$.
-To see that $Q_{1,\gamma}$ is invariant under smooth deformations of $\gamma$, consider $Q_{1,\gamma}-Q_{1,\gamma^{\prime}}=\int_{\gamma-\gamma^{\prime}}*J^{1}_{e}=\int_{\Sigma}d\left(*J^{1}_{e}\right)=0$, where $\Sigma$ is the surface bounded by $\gamma-\gamma^{\prime}$ and we have used Stokes’ theorem.
+which depends only on the homology class of the closed curve $\gamma$.
+To see that $Q_{1,\gamma}$ is invariant under smooth deformations of $\gamma$, consider $Q_{1,\gamma}-Q_{1,\gamma^{\prime}}=\int_{\gamma-\gamma^{\prime}}*J^{1}_{e}=\int_{\Sigma}d\left(*J^{1}_{e}\right)=0$, where $\Sigma$ is a surface with $\partial\Sigma=\gamma-\gamma^{\prime}$ and we have used Stokes’ theorem.
 This invariance confirms the topological nature of $Q_{1,\gamma}$.
@@
 The object charged under this symmetry is the closed worldsheet of the loop excitation described by $b^{1}$.
-Correspondingly, $Q_{1,\gamma}$ counts the number of intersections between $\gamma$ and the three-dimensional volume bounded by the loop worldsheet.
+Correspondingly, $Q_{1,\gamma}$ counts the number of intersections between $\gamma$ and a three-dimensional volume bounded by the loop worldsheet.
 We therefore identify $U_{e_{1}}\left(\gamma\right)$ as a Wilson operator.
-Imposing large gauge invariance together with the quantization of $\int_{\gamma}a^{1}$, we find that $e_{1}$ is an integer defined modulo $N_{1}$.
+Imposing large-gauge invariance together with the quantization of $\int_{\gamma}a^{1}$, we find that $e_{1}$ is an integer defined modulo $N_{1}$.
 As a result, the symmetry operator $U_{e_{1}}\left(\gamma\right)$ with $e_{1}\in\left\{0,1,\cdots,N_{1}-1\right\}$ implements a $\mathbb{Z}_{N_{1}}$ 22-form symmetry.
@@
 $U_{e_{3}}\left(\sigma\right)=\exp\left({\rm i}e_{3}\int_{\sigma}b^{3}\right),$ (25)
-where $e_{2}\in\left\{1,2,\cdots,N_{2}-1\right\}$ and $e_{3}\in\left\{1,2,\cdots,N_{3}-1\right\}$.
+where $e_{2}\in\left\{0,1,\cdots,N_{2}-1\right\}$ and $e_{3}\in\left\{0,1,\cdots,N_{3}-1\right\}$.
 We note that the symmetry operators $U_{e_{1}}\left(\gamma\right)$, $U_{e_{2}}\left(\gamma\right)$, and $U_{e_{3}}\left(\sigma\right)$ coincide with the Wilson operators of Abelian particle and loop excitations.
@@
 As an example, consider a local variation of $a^{1}$, $a^{1}\rightarrow a^{1}+\delta a^{1}$.
-The variation of the action is $\delta S\left[\delta a^{1}\right]=\int_{M^{4}}\frac{N_{1}}{2\pi}\delta a^{1}db^{1}+q\delta a^{1}a^{2}b^{3}$.
+The variation of the action is
+$\delta S\left[\delta a^{1}\right]=\int_{M^{4}}\left(\frac{N_{1}}{2\pi}\,\delta a^{1}\,db^{1}+q\,\delta a^{1}a^{2}b^{3}\right).$
 Requiring $\delta S\left[\delta a^{1}\right]=0\mod{2\pi}$ yields
@@
 The answer is affirmative, but only under additional conditions.
-The key step is to rewrite the EoM into the form of a vanishing total derivative.
+The key step is to rewrite the EoM in the form of a vanishing total derivative.
 To this end, we introduce two auxiliary fields $\phi^{2}$ and $\beta^{3}$ such that $d\phi^{2}=\frac{pN_{2}}{N_{123}}a^{2}$ and $d\beta^{3}=\frac{pN_{3}}{N_{123}}b^{3}$.
-We notice that $d\phi^{2}$ ($d\beta^{3}$) is an integral multiple of $a^{2}$ ($b^{3}$).
-Therefore, the fields $\phi^{2}$ and $\beta^{3}$ satisfy the Dirac quantization condition $\oint d\phi^{2}\in 2\pi\mathbb{Z}$ and $\oint d\beta^{3}\in 2\pi\mathbb{Z}$ when $\oint a^{2}$ and $\oint b^{3}$ take specific values.
+We note that $d\phi^{2}$ ($d\beta^{3}$) is an integer multiple of $a^{2}$ ($b^{3}$).
+Therefore, the fields $\phi^{2}$ and $\beta^{3}$ satisfy the Dirac quantization conditions $\oint d\phi^{2}\in 2\pi\mathbb{Z}$ and $\oint d\beta^{3}\in 2\pi\mathbb{Z}$ when $\oint a^{2}$ and $\oint b^{3}$ take the corresponding allowed values.
 With $\phi^{2}$ and $\beta^{3}$, the EoM (26) can be rewritten as
 $\frac{N_{1}}{2\pi}\left(db^{1}+\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}d\beta^{3}+\frac{1}{2}d\phi^{2}\frac{N_{3}}{2\pi}b^{3}\right)=0.$ (27)
-Here, the two coefficients $\frac{1}{2}$ are not the result of the gauge invariance: the above equation still holds if one chooses two coefficient $c_{1}$ and $c_{2}$ with $c_{1}+c_{2}=1$.
-Instead, we split the term $\frac{2\pi qa^{2}b^{3}}{N_{1}}$ in a symmetric way in order to respect the fact that the twisted term in the action (14) is symmetric under exchanging $a^{1}$ and $a^{2}$, up to an overall anti-symmetry of the wedge product.
-In this way, the currents extracted from the EoMs have the same symmetric formalism, see Table 1.
+Here, the two coefficients $\frac{1}{2}$ are not fixed by gauge invariance: the above equation still holds if one replaces them by two coefficients $c_{1}$ and $c_{2}$ with $c_{1}+c_{2}=1$.
+Rather, we split the term $\frac{2\pi q}{N_{1}}a^{2}b^{3}$ symmetrically in order to reflect that the twisted term in the action (14) is symmetric under exchanging $a^{1}$ and $a^{2}$, up to the overall antisymmetry of the wedge product.
+With this convention, the currents extracted from the EoMs take a symmetric form; see Table 1.
 Using $\frac{N_{2}}{2\pi}da^{2}=0$ and $\frac{N_{3}}{2\pi}db^{3}=0$, we further obtain
@@
 from which we can extract the continuity equation
-$d\left(*J_{m}^{1}\right)=0,\ J_{m}^{1}=*\left(b^{1}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}\right),$ (29)
+$d\left(*J_{m}^{1}\right)=0,\qquad J_{m}^{1}=*\left(b^{1}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}\right),$ (29)
 where $J^{1}_{m}$ is a $2$-form current. Analogous to ordinary symmetries, this conservation law implies a conserved quantity
@@
 Equivalently, these conditions can be expressed as $\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}=0$ and $\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}=0$, because on the closed manifold $\sigma$ one has $\int_{\sigma}d\beta^{3}=0$, and $\int_{\gamma}d\phi^{2}=0$ for any closed curve $\gamma\subset\sigma$.
-In other words, although $\mathcal{U}_{1}$ is a unitary operator originating from $Q_{m}^{1}$, it defines a symmetry operation only when $\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}=0$ and $\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}=0$ are satisfied.
+In other words, although $\mathcal{U}_{1}$ is a unitary operator built from $Q_{m}^{1}$, it defines a symmetry operation only when $\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}=0$ and $\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}=0$ are satisfied.
 Accordingly, we define the symmetry operator as
@@
 This is analogous to the 22d toric code, where moving an $m$ excitation along a closed path measures the number of enclosed $e$ excitations.
-The enclosed number of $\mathbb{Z}_{N_{1}}$ gauge charges, $n_{e_{1}}$, is reflected by the phase factor $\exp\left({\rm i}\frac{2\pi}{N_{1}}\times n_{e_{1}}\right)$.
+The enclosed number of $\mathbb{Z}_{N_{1}}$ gauge charges, $n_{e_{1}}$, is reflected by the phase factor $\exp\left({\rm i}\frac{2\pi}{N_{1}}\,n_{e_{1}}\right)$.
 Since the loop carries one unit of gauge flux, the measurement outcome is defined modulo $N_{1}$.
@@
 with
-$\mathcal{U}_{m_{1}}=e^{{\rm i}Q^{1}_{m_{1}}}=e^{{\rm i}\left(\int_{\sigma}m_{1}b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta_{m_{1}}^{3}+\frac{1}{2}\phi_{m_{1}}^{2}\frac{N_{3}}{2\pi}b^{3}\right)}.$ (36)
+$\mathcal{U}_{m_{1}}$ $=e^{{\rm i}Q^{1}_{m_{1}}}$
+$=\exp\left[{\rm i}\left(\int_{\sigma}m_{1}b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta_{m_{1}}^{3}+\frac{1}{2}\phi_{m_{1}}^{2}\frac{N_{3}}{2\pi}b^{3}\right)\right].$ (36)
 The delta functions in $L^{1}_{m_{1}}\left(\sigma\right)$ project onto a larger subspace than those in $L^{1}\left(\sigma\right)$, in the sense that more gauge-field configurations are allowed.
@@
 Since $L_{m_{1}}^{1}\left(\sigma\right)$ acts on states in the Hilbert space, any overall coefficient can be absorbed into the normalization of the state.
-In the same manner, we can obtain the EoM of gauge field $b^{2}$,
+In the same manner, we can obtain the EoM of the gauge field $b^{2}$,
 $\frac{N_{2}}{2\pi}db^{2}-qa^{1}b^{3}=0.$ (37)
 From it we can extract a type-II conserved current, i.e., a continuity equation,
-$d\left(*J_{m}^{2}\right)=0,\ J_{m}^{2}=*\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right),$ (38)
+$d\left(*J_{m}^{2}\right)=0,\qquad J_{m}^{2}=*\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right),$ (38)
 where the auxiliary fields $\phi^{1}$ and $\beta^{3}$ satisfy $d\phi^{1}=\frac{pN_{1}}{N_{123}}a^{1}$ and $d\beta^{3}=\frac{pN_{3}}{N_{123}}b^{3}$.
@@
 provided that $d\phi_{m_{2}}^{1}=\frac{m_{2}pN_{1}}{N_{123}}a^{1}$ and $d\beta_{m_{2}}^{3}=\frac{m_{2}pN_{3}}{N_{123}}b^{3}$. The corresponding symmetry operator is
-$L_{m_{2}}^{2}\left(\sigma\right)=\mathcal{U}_{m_{2}}\delta\left(\int_{\sigma}\frac{m_{2}pN_{3}}{N_{123}}b^{3}\right)\delta\left(\int_{\gamma\subset\sigma}\frac{m_{2}pN_{2}}{N_{123}}a^{2}\right),$ (41)
+$L_{m_{2}}^{2}\left(\sigma\right)=\mathcal{U}_{m_{2}}\,\delta\left(\int_{\sigma}\frac{m_{2}pN_{3}}{N_{123}}b^{3}\right)\delta\left(\int_{\gamma\subset\sigma}\frac{m_{2}pN_{1}}{N_{123}}a^{1}\right),$ (41)
 with
-$\mathcal{U}_{m_{2}}=$ $e^{{\rm i}\int_{\sigma}*J_{m_{2}}^{2}}=e^{{\rm i}\left(\int_{\sigma}m_{2}b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta_{m_{2}}^{3}-\frac{1}{2}\phi_{m_{2}}^{1}\frac{N_{3}}{2\pi}b^{3}\right)}.$ (42)
+$\mathcal{U}_{m_{2}}$ $=e^{{\rm i}\int_{\sigma}*J_{m_{2}}^{2}}$
+$=\exp\left[{\rm i}\left(\int_{\sigma}m_{2}b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta_{m_{2}}^{3}-\frac{1}{2}\phi_{m_{2}}^{1}\frac{N_{3}}{2\pi}b^{3}\right)\right].$ (42)
 $L_{m_{2}}^{2}\left(\sigma\right)$ is a non-invertible $1$-form symmetry operator, and it corresponds to the non-Abelian loop excitation carrying $m_{2}$ units of $\mathbb{Z}_{N_{2}}$ gauge flux.
-The above discussion also applies to the EoM of gauge field $a^{3}$,
+The above discussion also applies to the EoM of the gauge field $a^{3}$,
 $\frac{N_{3}}{2\pi}da^{3}+qa^{1}a^{2}=0.$ (43)
 We can similarly extract a type-II conserved current given by
-$d\left(*J_{m}^{3}\right)=0,\ J_{m}^{3}=-*\left(a^{3}+\frac{1}{2}\phi^{1}\frac{N_{2}}{2\pi}a^{2}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi^{2}\right),$ (44)
+$d\left(*J_{m}^{3}\right)=0,\qquad J_{m}^{3}=-*\left(a^{3}+\frac{1}{2}\phi^{1}\frac{N_{2}}{2\pi}a^{2}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi^{2}\right),$ (44)
 when $d\phi^{1}=\frac{pN_{1}}{N_{123}}a^{1}$ and $d\phi^{2}=\frac{pN_{2}}{N_{123}}a^{2}$ are satisfied. The corresponding symmetry operator is
@@
 with
-$\mathcal{U}_{m_{3}}=$ $e^{{\rm i}\int_{\gamma}*J_{m_{3}}^{3}}=e^{{\rm i}\left(\int_{\gamma}m_{3}a^{2}+\frac{1}{2}\phi_{m_{3}}^{1}\frac{N_{2}}{2\pi}a^{2}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi_{m_{3}}^{2}\right)}.$ (48)
+$\mathcal{U}_{m_{3}}$ $=e^{{\rm i}\int_{\gamma}*J_{m_{3}}^{3}}$
+$=\exp\left[{\rm i}\left(\int_{\gamma}m_{3}a^{3}+\frac{1}{2}\phi_{m_{3}}^{1}\frac{N_{2}}{2\pi}a^{2}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi_{m_{3}}^{2}\right)\right].$ (48)
 $L_{m_{3}}^{3}\left(\gamma\right)$ is a non-invertible $2$-form symmetry operator and corresponds to the non-Abelian particle excitation carrying $m_{3}$ units of $\mathbb{Z}_{N_{3}}$ gauge charge.
@@
 III.
-Fusion rules characterize the outcome of consecutively applying two symmetry operators supported on the same submanifold, and they encode the underlying algebraic structure of a generalized symmetry.
-From a field-theoretical perspective, fusion is realized by inserting symmetry operators sequentially into correlation functions.
+Fusion rules characterize the result of composing two symmetry operators supported on the same submanifold, and they encode the algebraic structure underlying a generalized symmetry.
+From a field-theoretical perspective, fusion is realized by inserting the corresponding topological operators in close succession in a correlation function.
 For invertible symmetries, fusion rules are group-like: each symmetry operator is labeled by a group element, and their fusion follows the group multiplication law.
-By contrast, for non-invertible symmetries, fusion rules are intrinsically non-group-like.
-In particular, the fusion of two non-invertible symmetry operators can produce a linear combination of several symmetry operators, rather than a single one.
-To analyze fusion rules in a systematic manner, it is crucial to first identify the complete set of symmetry operators in the theory.
+By contrast, for non-invertible symmetries, fusion is intrinsically non-group-like.
+In particular, the product of two non-invertible symmetry operators can decompose into a linear combination of symmetry operators, rather than closing onto a single one.
+To analyze fusion rules systematically, it is crucial to first identify the complete set of symmetry operators in the theory.
 In Sec.
 IV.1, we show that additional symmetry operators can be generated from new conserved currents, which themselves arise from linear combinations of existing conserved currents.
-This construction elucidates the general structure of symmetry operators relevant for fusion.
+This construction clarifies the general structure of symmetry operators that can appear in fusion.
 In Sec.
 IV.2, we specialize to the simplest nontrivial case by choosing the gauge group in the action (14) to be $G=(\mathbb{Z}_{2})^{3}$.
-This choice allows us to explicitly demonstrate representative examples of fusion rules and to highlight characteristic features of non-invertible symmetries in a concrete and transparent setting.
+This choice allows us to present representative fusion processes explicitly and to highlight characteristic features of non-invertible symmetries in a concrete and transparent setting.
 Guided by these examples, we present a general expression for fusion rules in Sec.
@@
 Finally, in Sec.
-IV.4, we provide additional examples of fusion rules for different choices of gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$, illustrating how the fusion structure depends on the underlying gauge data.
+IV.4, we provide additional examples for different choices of gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$, illustrating how the fusion structure depends on the underlying gauge data.
 IV.1 Generating new conserved currents by linear combinations
@@
 $+d\left(\frac{1}{2}\phi_{m_{1},m_{2}}^{2,1}\frac{N_{3}}{2\pi}b^{3}\right)-d\left(\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta_{m_{3},m_{1}}^{3,1}\right)+d\left(\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta_{m_{2},m_{3}}^{3,2}\right),$ (72)
-from which we may identify
+from which we may identify an conserved current
 $*J_{m_{1},m_{2},m_{3}}=$ $m_{1}b^{1}+m_{2}b^{2}+\delta^{\perp}\left(\gamma_{p}\right)m_{3}a^{3}$
@@
 By further introducing projectors that enforce the conditions under which $d\left(*J_{m_{1},m_{2},m_{3}}\right)=0$, we obtain the symmetry operator
-$L_{m_{1}m_{2}m_{3}}\left(\sigma,\gamma_{p}\right)=\mathcal{U}_{m_{1}m_{2}m_{3}}\mathbb{P}_{m_{1}m_{2}m_{3}},$ (75)
+$L_{m_{1}m_{2}m_{3}}\left(\sigma,\gamma_{p}\right)=\mathcal{U}_{m_{1}m_{2}m_{3}}\left(\sigma,\gamma_{p}\right)\mathbb{P}_{m_{1}m_{2}m_{3}},$ (75)
 where the projectors are implemented by delta functions,
@@
 $\left\langle e^{{\rm i}2\int_{\gamma_{p}}a^{1}}\mathcal{O}\right\rangle$
-$=$ $\frac{1}{\mathcal{Z}}\sum_{w\in\left\{0,\pi\right\}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\gamma_{p}}a^{1}=w\right]e^{{\rm i}S_{aab}}e^{{\rm i}2\int_{\gamma_{p}}a^{1}}\mathcal{O},$ (78)
+$=$ $\frac{1}{\mathcal{Z}}\sum_{w\in\left\{0,\pi\right\}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\gamma_{p}}a^{1}=w\right]e^{{\rm i}S_{\text{BR}}}e^{{\rm i}2\int_{\gamma_{p}}a^{1}}\mathcal{O},$ (78)
 where $\mathcal{D}\left[a^{i},b^{i};\int_{\gamma_{p}}a^{1}=w\right]$ denotes configurations with $\int_{\gamma_{p}}a^{1}=w$.
@@
 $\left\langle\delta\left(\int_{\sigma}b^{3}\right)\mathcal{O}\right\rangle$
-$=$ $\frac{1}{\mathcal{Z}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\sigma}b^{3}=0\right]e^{{\rm i}S_{aab}}\mathcal{O}$
-$=$ $\frac{1}{\mathcal{Z}}\sum_{w\in\left\{0,\pi\right\}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\sigma}b^{3}=w\right]e^{{\rm i}S_{aab}}\frac{1}{2}\left(1+e^{{\rm i}\int_{\sigma}b^{3}}\right)\mathcal{O}.$ (84)
+$=$ $\frac{1}{\mathcal{Z}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\sigma}b^{3}=0\right]e^{{\rm i}S_{\text{BR}}}\mathcal{O}$
+$=$ $\frac{1}{\mathcal{Z}}\sum_{w\in\left\{0,\pi\right\}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\sigma}b^{3}=w\right]e^{{\rm i}S_{\text{BR}}}\frac{1}{2}\left(1+e^{{\rm i}\int_{\sigma}b^{3}}\right)\mathcal{O}.$ (84)
 In this way, we obtain $\left\langle\delta\left(\int_{\sigma}b^{3}\right)\mathcal{O}\right\rangle=\left\langle\frac{1}{2}\left(1+e^{{\rm i}\int_{\sigma}b^{3}}\right)\mathcal{O}\right\rangle$.
@@
 $L_{m_{1}=1}\times L_{m_{2}=1}=\frac{1}{2}L_{m_{1}=1,m_{2}=1}+\frac{1}{2}U_{100}L_{m_{1}=1,m_{2}=1}.$ (89)
-Here, $U_{m_{1}=1}L_{m_{1}=1,m_{2}=1}$ denotes the symmetry operator obtained by fusing $U_{m_{1}=1}$ and $L_{m_{1}=1,m_{2}=1}$, whose expression is $e^{{\rm i}\int_{\gamma\subset\sigma}a^{1}+\ln\mathcal{U}_{m_{1}=1,m_{2}=1}}\delta\left(\int_{\gamma\subset\sigma}a^{2}-a^{1}\right)\delta\left(\int_{\sigma}b^{3}\right)$.
+Here, $U_{m_{1}=1}L_{m_{1}=1,m_{2}=1}$ denotes the symmetry operator obtained by fusing $U_{m_{1}=1}$ and $L_{m_{1}=1,m_{2}=1}$, whose expression is $e^{{\rm i}\int_{\gamma\subset\sigma}a^{1}}\mathcal{U}_{m_{1}=1,m_{2}=1}\delta\left(\int_{\gamma\subset\sigma}a^{2}-a^{1}\right)\delta\left(\int_{\sigma}b^{3}\right)$.
 According to the correspondence between symmetry operators and topological excitations, the fusion rules of symmetry operators can be mapped to those of topological excitations [74].
@@
 As discussed earlier, $U_{e_{1}e_{2}e_{3}}$ can also be interpreted as the Wilson operator of an Abelian topological excitation.
-In this interpretation, the integers $e_{i}$ specify the numbers of $\mathbb{Z}_{N_{i}}$ gauge charges and fluxes carried by the excitation.
+In this interpretation, the integers $e_{i}$ specify the $\mathbb{Z}_{N_{i}}$ gauge charges and fluxes carried by the excitation.
 Equivalently, $U_{e_{1}e_{2}e_{3}}$ can be viewed as the fusion product of $U_{e_{1}}\left(\gamma_{p}\right)$, $U_{e_{2}}\left(\gamma_{p}\right)$, and $U_{e_{3}}\left(\sigma\right)$.
@@
 The fusion rules of the non-invertible symmetry operators $L_{m_{1}m_{2}m_{3}}$, for example $L_{ijk}\times L_{rst}$, can be obtained systematically as follows.
-One first treats the fusion of the unitary parts and the projectors separately.
+We first treat the fusion of the unitary parts and the projectors separately.
 The fusion of the unitary operators is straightforward,
@@
 This result reproduces precisely the unitary component of the symmetry operator $L_{\left(i+r\right)\left(j+s\right)\left(k+t\right)}$.
-By contrast, the fusion of the projectors yields the product $\mathbb{P}_{ijk}\mathbb{P}_{rst}$, which can be rewritten, inside correlation functions, as a linear combination of invertible symmetry operators, for example as in Eq.
+By contrast, fusing the projectors yields the product $\mathbb{P}_{ijk}\mathbb{P}_{rst}$, which can be rewritten, inside correlation functions, as a linear combination of invertible symmetry operators, for example as in Eq.
 (84).
@@
 and
-$L_{021}\left(\sigma,\gamma_{p}\right)$ $=\mathcal{U}_{021}\delta\left(2\int_{\gamma_{p}}a^{1}\right)\delta\left(\int_{\gamma_{p}}a^{1}\right)\delta\left(2\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{2}\right).$ (101)
+$L_{021}\left(\sigma,\gamma_{p}\right)$ $=\mathcal{U}_{021}\delta\left(2\int_{\gamma_{p}}a^{1}\right)\delta\left(\int_{\gamma_{p}}a^{1}\right)$
+$\delta\left(2\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{2}\right).$ (101)
 The fusion of the projectors yields $\mathbb{P}_{120}\times\mathbb{P}_{021}=\left(\frac{1}{3}\right)^{3}\sum_{l_{1},l_{2},l_{3}=0}^{2}U_{l_{1}l_{2}l_{3}}$, while the unitary parts fuse as $\mathcal{U}_{120}\times\mathcal{U}_{021}=\mathcal{U}_{111}$.
 Therefore, the fusion channels contain $L_{111}\left(\sigma,\gamma_{p}\right)=\mathcal{U}_{111}\mathbb{P}_{111}$, with projector
-$\mathbb{P}_{111}=$ $\delta\left(\int_{\gamma_{p}}a^{2}-\int_{\gamma_{p}}a^{1}\right)\delta\left(\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{1}\right)\delta\left(\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{2}\right).$ (102)
+$\mathbb{P}_{111}=$ $\delta\left(\int_{\gamma_{p}}a^{2}-\int_{\gamma_{p}}a^{1}\right)\delta\left(\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{1}\right)$
+$\delta\left(\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{2}\right).$ (102)
 Expanding $\mathbb{P}_{111}$ gives
@@
 The independent equations are $k_{3}-k_{1}=0\mod 4$ and $k_{2}-3k_{1}=0\mod 4$.
-Using $4\int_{\gamma_{p}}a^{1}\in 2\pi\mathbb{Z}$, $8\int_{\gamma_{p}}a^{2}\in 2\pi\mathbb{Z}$, and $12\int_{\sigma}b^{3}\in 2\pi\mathbb{Z}$, we can simplified $\mathbb{P}_{357}$ as
+Using $4\int_{\gamma_{p}}a^{1}\in 2\pi\mathbb{Z}$, $8\int_{\gamma_{p}}a^{2}\in 2\pi\mathbb{Z}$, and $12\int_{\sigma}b^{3}\in 2\pi\mathbb{Z}$, we can simplify $\mathbb{P}_{357}$ as
 $\mathbb{P}_{357}=$ $\delta\left(3\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{1}\right)\delta\left(2\int_{\gamma_{p}}a^{2}-3\int_{\gamma_{p}}a^{1}\right)$
@@
 On the other hand, for $L_{200}=\mathcal{U}_{200}\delta\left(\frac{2pN_{2}}{N_{123}}\int_{\gamma_{p}}a^{2}\right)\delta\left(\frac{2pN_{3}}{N_{123}}\int_{\sigma}b^{3}\right)$, the projectors are $\mathbb{P}_{200}=\delta\left(8\int_{\gamma_{p}}a^{2}\right)\delta\left(12\int_{\sigma}b^{3}\right)$, which impose no constraints on the values of $\int_{\gamma_{p}}a^{2}$ and $\int_{\sigma}b^{3}$.
-This is because $8\int_{\gamma_{p}}a^{2}$ and $12\int_{\sigma}b^{3}$ always equal $0\mod 2\pi$, and the delta functions always return $1$.
+This is because $8\int_{\gamma_{p}}a^{2}$ and $12\int_{\sigma}b^{3}$ always equal to $0\mod 2\pi$, and the delta functions always return $1$.
 Therefore,
@@
 The fusion rules of invertible symmetry operators are group-like, while those of non-invertible symmetry operators depend on the fusion of the projectors (delta functions).
-For a general effective field theory $S_{aab}$ with gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$ and a level-$p$ twisted term, the greatest common divisor $N_{123}$ and $p$ determine how a projector can be expressed as a sum of invertible operators in the context of correlation functions, which in turn determines the fusion of the projectors.
+For a general effective field theory $S_{\text{BR}}$ with gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$ and a level-$p$ twisted term, the greatest common divisor $N_{123}$ and $p$ determine how a projector can be expressed as a sum of invertible operators in the context of correlation functions, which in turn determines the fusion of the projectors.
 V Anomalies of (non-)invertible higher-form symmetries
@@
 Below, we first illustrate the mixed anomaly between two higher-form symmetries using the 33d toric code and its effective field theory.
-We then investigate the mixed anomalies among the generalized higher-form symmetries in the BR topological order described by $S_{aab}$ in Sec.
+We then investigate the mixed anomalies among the generalized higher-form symmetries in the BR topological order described by $S_{\text{BR}}$ in Sec.
 III.
-Table 3: Mixed anomalies of generalized symmetries in the effective field theory $S_{aab}$.
+Table 3: Mixed anomalies of generalized symmetries in the effective field theory $S_{\text{BR}}$.
 $J_{e}^{1}=-*a^{1}$, $J_{e}^{2}=-*a^{2}$, and $J_{e}^{3}=*b^{3}$ are conserved currents that generate $2$-, 22-, and $1$-form invertible symmetries, respectively.
@@
 This preserved gauge invariance indicates no anomaly for this $2$-form symmetry at the level of the effective action.
-To gauge the $1$-form symmetry generated by $J_{m}^{1}$, we add the coupling term $\mathcal{B}_{m}^{1}\wedge*J_{m}^{1}$ to $S_{aab}$,
+To gauge the $1$-form symmetry generated by $J_{m}^{1}$, we add the coupling term $\mathcal{B}_{m}^{1}\wedge*J_{m}^{1}$ to $S_{\text{BR}}$,
 $S_{\text{gauged}}^{m_{1}}=$ $\int_{M^{4}}\sum_{i=1}^{3}\frac{N_{i}}{2\pi}b^{i}da^{i}+qa^{1}a^{2}b^{3}$
@@
 For later reference, we list the coupling terms and their associated gauge transformations:
-$S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{2}\wedge*J_{e}^{2},\quad\begin{cases}\mathcal{C}_{e}^{2}\rightarrow&\mathcal{C}_{e}^{2}+dT_{e}^{2}\\ b^{2}\rightarrow&b^{2}+\frac{1}{N_{2}}T_{e}^{2}\end{cases};$ (139)
-$S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}^{2}\wedge*J_{m}^{2},\quad\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases};$ (140)
-$S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{e}^{3}\wedge*J_{e}^{3},\quad\begin{cases}\mathcal{B}_{e}^{3}\rightarrow&\mathcal{B}_{e}^{3}+dV_{e}^{3}\\ a^{3}\rightarrow&a^{3}-\frac{1}{N_{3}}V_{e}^{3}\end{cases};$ (141)
-$S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{m}^{3}\wedge*J_{m}^{3},\quad\begin{cases}\mathcal{C}_{m}^{3}\rightarrow&\mathcal{C}_{m}^{3}+dT_{m}^{3}\\ b^{3}\rightarrow&b^{3}+\frac{1}{N_{3}}T_{m}^{3}\end{cases}.$ (142)
-We now consider gauging two higher-form symmetries simultaneously by adding two coupling terms to $S_{aab}$.
+$S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{2}\wedge*J_{e}^{2},\quad\begin{cases}\mathcal{C}_{e}^{2}\rightarrow&\mathcal{C}_{e}^{2}+dT_{e}^{2}\\ b^{2}\rightarrow&b^{2}+\frac{1}{N_{2}}T_{e}^{2}\end{cases};$ (139)
+$S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}^{2}\wedge*J_{m}^{2},\quad\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases};$ (140)
+$S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{e}^{3}\wedge*J_{e}^{3},\quad\begin{cases}\mathcal{B}_{e}^{3}\rightarrow&\mathcal{B}_{e}^{3}+dV_{e}^{3}\\ a^{3}\rightarrow&a^{3}-\frac{1}{N_{3}}V_{e}^{3}\end{cases};$ (141)
+$S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{m}^{3}\wedge*J_{m}^{3},\quad\begin{cases}\mathcal{C}_{m}^{3}\rightarrow&\mathcal{C}_{m}^{3}+dT_{m}^{3}\\ b^{3}\rightarrow&b^{3}+\frac{1}{N_{3}}T_{m}^{3}\end{cases}.$ (142)
+We now consider gauging two higher-form symmetries simultaneously by adding two coupling terms to $S_{\text{BR}}$.
 All combinations are summarized in Table 3.
@@
 Gauging two invertible symmetries is always free of mixed anomalies. For example, gauging the symmetries generated by $J_{e}^{1}$ and $J_{e}^{2}$ leads to
-$S_{\text{gauged}}^{e_{1},e_{2}}=$ $S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{1}a^{1}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{2}a^{2},$ (143)
+$S_{\text{gauged}}^{e_{1},e_{2}}=$ $S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{1}a^{1}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{2}a^{2},$ (143)
 which is invariant under $\mathcal{C}_{e}^{1}\rightarrow\mathcal{C}_{e}^{1}+dT_{e}^{1}$, $\mathcal{C}_{e}^{2}\rightarrow\mathcal{C}_{e}^{2}+dT_{e}^{2}$ together with $b^{1}\rightarrow b^{1}+\frac{1}{N_{1}}T_{e}^{1}$ and $b^{2}\rightarrow b^{2}+\frac{1}{N_{2}}T_{e}^{2}$.
@@
 Indeed,
-$S_{\text{gauged}}^{e_{1},m_{2}}=$ $S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{1}a^{1}$
+$S_{\text{gauged}}^{e_{1},m_{2}}=$ $S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{1}a^{1}$
 $+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}^{2}\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right),$ (144)
@@
 For example, gauging the invertible symmetry generated by $J_{e}^{3}$ together with the non-invertible symmetry generated by $J_{m}^{3}$ is described by
-$S_{\text{gauged}}^{e_{3},m_{3}}=$ $S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{e}^{3}b^{3}$
+$S_{\text{gauged}}^{e_{3},m_{3}}=$ $S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{e}^{3}b^{3}$
 $+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{m}^{3}\left(a^{3}+\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\phi^{1}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi^{2}\right).$ (145)
@@
 As an illustrative example, consider gauging the non-invertible symmetries generated by $J_{m}^{1}$ and $J_{m}^{2}$ by adding the coupling terms
-$S_{\text{gauging}}^{m_{1},m_{2}}=$ $S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\left(b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}\right)$
+$S_{\text{gauging}}^{m_{1},m_{2}}=$ $S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\left(b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}\right)$
 $+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}^{2}\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right).$ (148)
@@
 VI Summary and Outlook
-VI.1 Summary
-In this paper, we investigate generalized symmetries in the topological field theory of (3+1)(3+1)D BR topological order.
-Our strategy is to construct symmetry operators from conserved currents: we extract continuity equations and conserved currents from the equations of motion of the gauge fields and then write down the associated symmetry operators.
-Geometrically, these generalized symmetries are classified as $1$-form or $2$-form, depending on the dimension of the supporting submanifold.
-Algebraically, the symmetries include both invertible and non-invertible ones.
-The invertible higher-form symmetries exhibit a $\mathbb{Z}_{N_{i}}$-type group structure: their symmetry operators obey group-like fusion rules that match the composition law of group elements, as summarized in Sec.
+In this work we presented a constructive, current-to-defect route to generalized symmetries—including intrinsically non-invertible higher-form symmetries—in a class of (3+1)(3+1)D twisted $BF$ topological field theories with an $a\wedge a\wedge b$ twist and gauge group $G=\prod_{i}\mathbb{Z}_{N_{i}}$.
+These TQFTs serve as effective continuum descriptions of Borromean-Rings (BR) topological order, i.e., three-dimensional non-Abelian topological orders supporting Borromean-Rings braiding.
+The starting point is the field-theory counterpart of “symmetry $\Leftrightarrow$ conserved current”: we derive continuity equations directly from the equations of motion, identify the corresponding conserved charges, and exponentiate them to obtain topological symmetry operators/defects supported on codimension-$p$ submanifolds.
+In this way, the generalized-symmetry data are obtained in a step-by-step and computable manner directly from the continuum action: currents $\rightarrow$ defects $\rightarrow$ fusion/anomaly diagnostics.
+A central structural result is that the equations of motion naturally organize the conserved currents into two qualitatively distinct classes, which in turn lead to two different symmetry structures:
+  • Type-I currents are conserved identically (in the sense of Bianchi/Noether-type identities) and generate invertible higher-form symmetries.
+Their symmetry operators admit inverses and obey group-like fusion, reproducing the expected higher-form composition laws (with minimal generators of $\mathbb{Z}_{N_{i}}$ type); see Sec.
 IV.3.
-Equivalently, each such operator admits an inverse whose fusion yields the identity operator.
-By contrast, the non-invertible symmetry operators (again of either $1$-form or $2$-form type) consist of a unitary operator dressed by a projector.
-The projector enforces the constraints under which the corresponding current is conserved.
-Since a projector has no inverse, the resulting symmetry is intrinsically non-invertible.
-Correspondingly, the fusion of non-invertible symmetry operators can admit multiple fusion channels, as illustrated in Sec.
+  • Type-II currents become conserved only after imposing additional consistency conditions on admissible gauge-field configurations.
+At the operator/defect level, these conditions are implemented by projector dressings of otherwise unitary topological operators.
+This provides an explicit and practical mechanism for intrinsic non-invertibility in (3+1)(3+1)D: the resulting higher-form symmetry operators do not admit inverses, and their fusion is generically multi-channeled, with the channel structure dictated by the fusion of the projector factors; see Sec.
 IV.4.
-We also analyze anomalies of these generalized symmetries by testing whether they can be gauged.
-Gauging a single symmetry amounts to coupling its current to a background field in a gauge-invariant way.
-However, when two symmetries are gauged simultaneously, a mixed anomaly may arise: the action with both coupling terms fails to remain gauge invariant.
-In some cases, such a mixed anomaly can be canceled by anomaly inflow from a higher-dimensional bulk, so that the combined bulk–boundary system is gauge invariant.
-In other cases, the mixed anomaly is intrinsic and persists because no bulk term can restore gauge invariance; see Table 3.
-Overall, our results provide a field-theoretical characterization of invertible and non-invertible higher-form symmetries in 33d topological order.
-We expect that the current-based construction developed here can be generalized to other 33d topological orders, including those described by quantum doubles of finite groups.
-In parallel, recent progress has established diagrammatic representations for 33d and 44d topological orders from continuum field theories, yielding stringent consistency conditions for higher-dimensional topological orders [78].
-It would be interesting to further clarify how the fusion rules of the symmetry operators constructed here are constrained by, and in turn inform, these diagrammatic consistency conditions.
-VI.2 Outlook
+With explicit (non-)invertible symmetry operators in hand, we then computed their fusion algebra by composing defects supported on the same submanifold.
+This yields a unified picture in which group-like fusion for invertible symmetries and multi-channel fusion for non-invertible symmetries arise from the same current-based construction, with the projector constraints providing a computable organizing principle for the non-group-like sector.
+We further diagnosed anomalies by coupling conserved currents to background gauge fields and testing gaugeability.
+While each generalized symmetry can be coupled in a gauge-invariant manner individually, simultaneous coupling of two symmetries can exhibit a mixed anomaly, signaled by an obstruction to maintaining gauge invariance.
+We identify two qualitatively distinct outcomes: (i) mixed anomalies that can be canceled by anomaly inflow from a one-higher-dimensional topological field theory, and (ii) intrinsic gauging obstructions already encoded in the (3+1)(3+1)D continuum theory, for which no bulk counterterm restores gauge invariance; see Table 3.
+Overall, our results provide a systematic field-theoretic characterization of invertible and non-invertible higher-form symmetries and their mixed anomalies in (3+1)(3+1)D TQFTs, organized as a practical pipeline from equations of motion to currents, symmetry defects, fusion, and gaugeability.
+We expect that this construction can be extended to other families of three-dimensional non-Abelian topological orders (including quantum doubles of finite groups) and can serve as a useful interface between condensed-matter realizations and the defect/anomaly language natural in SymTFT.
+In parallel, recent progress has established diagrammatic representations and consistency conditions for 33d and 44d topological orders from continuum field theories [76].
+It would be interesting to clarify how the fusion rules derived here are constrained by, and in turn sharpen, these diagrammatic consistency conditions.
 Several future directions are motivated by the present study.
@@
 Acknowledgements.
-This work was in part supported by National Natural Science Foundation of China (NSFC) under Grants No.
-12474149 and No.
-12274250, Research Center for Magnetoelectric Physics of Guangdong Province under Grant No.
-2024B0303390001, and Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices under Grant No.
-2022B1212010008.
+This work was in part supported by National Natural Science Foundation of China (NSFC) under Grants No. 12474149 and No. 12274250.
 Appendix A Generalized symmetries in lattice models
@@
 $\underbrace{\frac{1}{6}\sum_{l_{2}=0}^{5}e^{{\rm i}l_{2}\cdot 2\int_{\gamma_{p}}a^{2}}\times\frac{1}{6}\sum_{l_{3}=0}^{5}e^{{\rm i}l_{3}\cdot 2\int_{\sigma}b^{3}}}_{\text{from }\mathbb{P}_{123}\times\mathbb{P}_{234}}=\frac{1}{3}\sum_{l_{2}=0}^{2}e^{{\rm i}l_{2}\cdot 2\int_{\gamma_{p}}a^{2}}\times\frac{1}{3}\sum_{l_{3}=0}^{2}e^{{\rm i}l_{3}\cdot 2\int_{\sigma}b^{3}}.$ (187)
+This means that
+$\mathbb{P}_{123}\times\mathbb{P}_{234}=$ $\frac{1}{6}\times\frac{1}{3}\times\frac{1}{3}\sum_{l_{1}=0}^{5}\sum_{l_{2},l_{3}=0}^{2}e^{{\rm i}2l_{1}\int_{\gamma_{p}}a^{1}+{\rm i}l_{2}\cdot 2\int_{\gamma_{p}}a^{2}+{\rm i}l_{3}\cdot 2\int_{\sigma}b^{3}}.$ (188)
 Moreover, the factor $\sum_{h_{3}=0}^{5}e^{{\rm i}4h_{3}\cdot\left(\int_{\sigma}b^{3}+\int_{\gamma_{p}}a^{2}\right)}$ in $\mathbb{P}_{351}$, when multiplied by $\frac{1}{3}\sum_{m_{2}=0}^{2}e^{{\rm i}m_{2}\cdot 2\int_{\gamma_{p}}a^{2}}$, reproduces the term $\sum_{l_{2},l_{3}=0}^{5}e^{{\rm i}2l_{2}\int_{\gamma_{p}}a^{2}+{\rm i}2l_{3}\int_{\sigma}b^{3}}$ in $\mathbb{P}_{123}\times\mathbb{P}_{234}$:
@@
 $=$ $\underbrace{\frac{1}{6}\sum_{l_{2}=0}^{5}e^{{\rm i}l_{2}\cdot 2\int_{\gamma_{p}}a^{2}}\times\frac{1}{6}\sum_{l_{3}=0}^{5}e^{{\rm i}l_{3}\cdot 2\int_{\sigma}b^{3}}}_{\text{from }\mathbb{P}_{123}\times\mathbb{P}_{234}}.$ (189)
+This means that
+$\mathbb{P}_{351}\times\frac{1}{3}\sum_{m_{2}=0}^{2}e^{{\rm i}m_{2}\cdot 2\int_{\gamma_{p}}a^{2}}=$ $\underbrace{\frac{1}{6}\sum_{h_{2}=0}^{5}e^{{\rm i}2h_{2}\int_{\gamma_{p}}a^{1}}\times\frac{1}{6}\sum_{h_{3}=0}^{5}e^{{\rm i}4h_{3}\cdot\left(\int_{\sigma}b^{3}+\int_{\gamma_{p}}a^{2}\right)}}_{\mathbb{P}_{351}}\times\frac{1}{3}\sum_{m_{2}=0}^{2}e^{{\rm i}m_{2}\cdot 2\int_{\gamma_{p}}a^{2}}$
+$=$ $\frac{1}{6}\sum_{h_{2}=0}^{5}e^{{\rm i}2h_{2}\int_{\gamma_{p}}a^{1}}\times\frac{1}{3}\sum_{h_{3}=0}^{2}e^{{\rm i}h_{3}\cdot\left(2\int_{\sigma}b^{3}\right)}\times\frac{1}{3}\sum_{m_{2}=0}^{2}e^{{\rm i}m_{2}\cdot 2\int_{\gamma_{p}}a^{2}}$
+$=$ $\mathbb{P}_{123}\times\mathbb{P}_{234}$ (190)
 Therefore,
@@
 We denote
-$S_{\text{gauging}}^{m_{1},m_{2}}=S_{aab}+S_{\text{coupling}}.$ (191)
-Under $\begin{cases}\mathcal{B}_{m}^{1}\rightarrow&\mathcal{B}_{m}^{1}+dV_{m}^{1}\\ a^{1}\rightarrow&a^{1}-\frac{1}{N_{1}}V_{m}^{1}\end{cases}$ and $\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases}$, the variation of $S_{aab}$ is
-$\Delta S_{aab}=$ $\int_{M_{4}}-\frac{1}{2\pi}b^{1}dV_{m}^{1}-\frac{1}{2\pi}b^{2}dV_{m}^{2}+q\left(a^{1}-\frac{1}{N_{1}}V_{m}^{1}\right)\left(a^{2}-\frac{1}{N_{2}}V_{m}^{2}\right)b^{3}-qa^{1}a^{2}b^{3}$
+$S_{\text{gauging}}^{m_{1},m_{2}}=S_{\text{BR}}+S_{\text{coupling}}.$ (193)
+Under $\begin{cases}\mathcal{B}_{m}^{1}\rightarrow&\mathcal{B}_{m}^{1}+dV_{m}^{1}\\ a^{1}\rightarrow&a^{1}-\frac{1}{N_{1}}V_{m}^{1}\end{cases}$ and $\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases}$, the variation of $S_{\text{BR}}$ is
+$\Delta S_{\text{BR}}=$ $\int_{M_{4}}-\frac{1}{2\pi}b^{1}dV_{m}^{1}-\frac{1}{2\pi}b^{2}dV_{m}^{2}+q\left(a^{1}-\frac{1}{N_{1}}V_{m}^{1}\right)\left(a^{2}-\frac{1}{N_{2}}V_{m}^{2}\right)b^{3}-qa^{1}a^{2}b^{3}$
 $=$ $\int_{M_{4}}-\frac{1}{2\pi}b^{1}dV_{m}^{1}-\frac{1}{2\pi}b^{2}dV_{m}^{2}-\frac{q}{N_{1}}V_{m}^{1}a^{2}b^{3}-\frac{q}{N_{2}}a^{1}V_{m}^{2}b^{3}+\frac{q}{N_{1}N_{2}}V_{m}^{1}V_{m}^{2}b^{3}$ (194)
@@
 $-\frac{1}{2}\frac{1}{4\pi^{2}}\mathcal{B}_{m}^{2}V_{m}^{1}\beta^{3}+\frac{1}{2}\frac{1}{2\pi}\mathcal{B}_{m}^{2}\phi_{new}^{1}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\phi^{1}\frac{N_{3}}{2\pi}b^{3}$ (196)
-We now simplify $\Delta S_{aab}+\Delta S_{\text{coupling}}$.
-First, the term $\int_{M_{4}}-\frac{1}{2\pi}b^{1}dV_{m}^{1}-\frac{1}{2\pi}b^{2}dV_{m}^{2}$ from $\Delta S_{aab}$ cancels $\int_{M_{4}}\frac{1}{2\pi}dV_{m}^{1}b^{1}+\frac{1}{2\pi}dV_{m}^{2}b^{2}$ from $\Delta S_{\text{coupling}}$.
+We now simplify $\Delta S_{\text{BR}}+\Delta S_{\text{coupling}}$.
+First, the term $\int_{M_{4}}-\frac{1}{2\pi}b^{1}dV_{m}^{1}-\frac{1}{2\pi}b^{2}dV_{m}^{2}$ from $\Delta S_{\text{BR}}$ cancels $\int_{M_{4}}\frac{1}{2\pi}dV_{m}^{1}b^{1}+\frac{1}{2\pi}dV_{m}^{2}b^{2}$ from $\Delta S_{\text{coupling}}$.
 Assuming $M_{4}$ is closed, we use
@@
 which implies
-$\int_{M_{4}}\underbrace{-\frac{1}{2}\frac{1}{2\pi}\frac{N_{2}}{2\pi}dV_{m}^{1}a^{2}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}-\underbrace{\frac{1}{2}\frac{1}{2\pi}V_{m}^{1}\frac{N_{2}}{2\pi}a^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{1}}V_{m}^{1}a^{2}b^{3}$ from $\Delta S_{aab}$}}=0.$ (195)
+$\int_{M_{4}}\underbrace{-\frac{1}{2}\frac{1}{2\pi}\frac{N_{2}}{2\pi}dV_{m}^{1}a^{2}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}-\underbrace{\frac{1}{2}\frac{1}{2\pi}V_{m}^{1}\frac{N_{2}}{2\pi}a^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{1}}V_{m}^{1}a^{2}b^{3}$ from $\Delta S_{\text{BR}}$}}=0.$ (197)
 Similarly, from
@@
 we obtain
-$\int_{M_{4}}\underbrace{\frac{1}{2}\frac{1}{2\pi}dV_{m}^{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}-\underbrace{\frac{1}{2}\frac{1}{2\pi}\frac{N_{1}}{2\pi}a^{1}V_{m}^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{2}}a^{1}V_{m}^{2}b^{3}$ from $\Delta S_{aab}$}}=0.$ (196)
+$\int_{M_{4}}\underbrace{\frac{1}{2}\frac{1}{2\pi}dV_{m}^{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}-\underbrace{\frac{1}{2}\frac{1}{2\pi}\frac{N_{1}}{2\pi}a^{1}V_{m}^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{2}}a^{1}V_{m}^{2}b^{3}$ from $\Delta S_{\text{BR}}$}}=0.$ (198)
 Finally, using
@@
 we find
-$\int_{M_{4}}\underbrace{\frac{1}{2}\frac{1}{4\pi^{2}}dV_{m}^{1}V_{m}^{2}\beta^{3}-\frac{1}{2}\frac{1}{4\pi^{2}}dV_{m}^{2}V_{m}^{1}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}+\underbrace{\frac{1}{2}\frac{1}{4\pi^{2}}V_{m}^{1}V_{m}^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{1}N_{2}}V_{m}^{1}V_{m}^{2}b^{3}$ from $\Delta S_{aab}$ }}=0.$ (197)
-Up to this point, we have reduced $\Delta S_{aab}+\Delta S_{\text{coupling}}$ to
-$\Delta S_{aab}+\Delta S_{\text{coupling}}=$ $\int_{M_{4}}-\frac{1}{2}\frac{q}{N_{1}}V_{m}^{1}a^{2}b^{3}-\frac{1}{2}\frac{q}{N_{2}}a^{1}V_{m}^{2}b^{3}+\frac{1}{2}\frac{q}{N_{1}N_{2}}V_{m}^{1}V_{m}^{2}b^{3}$
+$\int_{M_{4}}\underbrace{\frac{1}{2}\frac{1}{4\pi^{2}}dV_{m}^{1}V_{m}^{2}\beta^{3}-\frac{1}{2}\frac{1}{4\pi^{2}}dV_{m}^{2}V_{m}^{1}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}+\underbrace{\frac{1}{2}\frac{1}{4\pi^{2}}V_{m}^{1}V_{m}^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{1}N_{2}}V_{m}^{1}V_{m}^{2}b^{3}$ from $\Delta S_{\text{BR}}$ }}=0.$ (199)
+Up to this point, we have reduced $\Delta S_{\text{BR}}+\Delta S_{\text{coupling}}$ to
+$\Delta S_{\text{BR}}+\Delta S_{\text{coupling}}=$ $\int_{M_{4}}-\frac{1}{2}\frac{q}{N_{1}}V_{m}^{1}a^{2}b^{3}-\frac{1}{2}\frac{q}{N_{2}}a^{1}V_{m}^{2}b^{3}+\frac{1}{2}\frac{q}{N_{1}N_{2}}V_{m}^{1}V_{m}^{2}b^{3}$
 $+\frac{1}{2}\frac{1}{2\pi}dV_{m}^{1}\phi_{new}^{2}\frac{N_{3}}{2\pi}b^{3}+\frac{1}{2}\frac{1}{4\pi^{2}}\mathcal{B}_{m}^{1}V_{m}^{2}\beta^{3}+\frac{1}{2}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\phi_{new}^{2}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\phi^{2}\frac{N_{3}}{2\pi}b^{3}$
 $+\frac{1}{2}\frac{1}{2\pi}dV_{m}^{2}\phi_{new}^{1}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{1}{4\pi^{2}}\mathcal{B}_{m}^{2}V_{m}^{1}\beta^{3}+\frac{1}{2}\frac{1}{2\pi}\mathcal{B}_{m}^{2}\phi_{new}^{1}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\phi^{1}\frac{N_{3}}{2\pi}b^{3}$ (200)
-We can now explain why we cannot find a bulk term $S_{\text{bulk}}$ such that $\Delta S_{aab}+\Delta S_{\text{coupling}}+\Delta S_{\text{bulk}}=0$ under the gauge transformations $\begin{cases}\mathcal{B}_{m}^{1}\rightarrow&\mathcal{B}_{m}^{1}+dV_{m}^{1}\\ a^{1}\rightarrow&a^{1}-\frac{1}{N_{1}}V_{m}^{1}\end{cases}$ and $\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases}$.
+We can now explain why we cannot find a bulk term $S_{\text{bulk}}$ such that $\Delta S_{\text{BR}}+\Delta S_{\text{coupling}}+\Delta S_{\text{bulk}}=0$ under the gauge transformations $\begin{cases}\mathcal{B}_{m}^{1}\rightarrow&\mathcal{B}_{m}^{1}+dV_{m}^{1}\\ a^{1}\rightarrow&a^{1}-\frac{1}{N_{1}}V_{m}^{1}\end{cases}$ and $\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases}$.
 The action $S_{\text{bulk}}$ must be a functional of background fields only, since $a^{i}$ and $b^{i}$ live only in (3+1)(3+1)D.
-However, $\Delta S_{aab}+\Delta S_{\text{coupling}}$ contains terms that depend explicitly on $a^{i}$ and $b^{i}$.
+However, $\Delta S_{\text{BR}}+\Delta S_{\text{coupling}}$ contains terms that depend explicitly on $a^{i}$ and $b^{i}$.
 Such terms cannot be canceled by $\Delta S_{\text{bulk}}$, because $\Delta S_{\text{bulk}}$ cannot involve $a^{i}$ or $b^{i}$.
-Therefore, a bulk term $S_{\text{bulk}}$ satisfying $\Delta S_{aab}+\Delta S_{\text{coupling}}+\Delta S_{\text{bulk}}=0$ does not exist.
+Therefore, a bulk term $S_{\text{bulk}}$ satisfying $\Delta S_{\text{BR}}+\Delta S_{\text{coupling}}+\Delta S_{\text{bulk}}=0$ does not exist.